### v6

# A new framework for nonlinear Kalman filters

Abstract
The Kalman filter (KF) is a state estimation algorithm that optimally combines system knowledge and measurements to minimize the mean squared error of the estimated states. While KF was initially designed for linear systems, numerous extensions of it, such as extended Kalman filter (EKF), unscented Kalman filter (UKF), cubature Kalman filter (CKF), etc., have been proposed for nonlinear systems. Although different types of nonlinear KFs have different pros and cons, they all use the same framework of linear KF. Yet, according to what we found in this paper, the framework often gives overconfident and less accurate state estimations when the measurement functions are nonlinear. Therefore, in this study, we designed a new framework for nonlinear KFs and showed theoretically and empirically that the new framework estimates the states and covariance matrix more accurately than the old one. The new framework was tested on four different nonlinear KFs and five different tasks, showcasing its ability to reduce the estimation errors by several orders of magnitude in low-measurement-noise conditions, with only about a 10–90% increase in computational time. To the best of our knowledge, all existing types of nonlinear KFs can benefit from the new framework, and the benefit will increase as the sensors become more and more accurate in the future. As an example, EKF, the simplest nonlinear KF that was previously believed to work poorly for strongly nonlinear systems, can now perform better than the iterated extended Kalman filter (IEKF) and provide fast and fairly accurate state estimations with the help of the proposed new framework. The codes are available at https://github.com/Shida-Jiang/A-new-framework-for-nonlinear-Kalman-filters.
{IEEEkeywords}
Cubature Kalman filters, extended Kalman filters, Kalman filters, nonlinear filters, state estimation, unscented Kalman filters.
1 Introduction
In control theory, the Kalman filter (KF) is a state estimation algorithm designed for linear systems. In such a system, a set of linear state transition functions describes how the states evolve, and a set of linear measurement functions describes the measured values of certain linear combinations of states. It is assumed that both the state transition and measurement functions contain some additive noise. The KF optimally combines the system knowledge and measurements to minimize the sum of the states’ variances when both process and measurement noise follow a normal (Gaussian) distribution. Since its birth in about 1960, KF has achieved massive success in navigation [1], motion control [2], statistics [3], signal processing [4], etc.
The most significant limitation of KF is that it can only be applied to systems with linear state transition functions and linear measurement functions. Unfortunately, many real-life systems are nonlinear [5]. Therefore, some extensions of KF have been developed to solve the state estimation problem for these nonlinear systems, and the most well-known ones are the extended Kalman filter (EKF) and the unscented Kalman filter (UKF).
EKF linearizes the state transition functions and measurement functions locally at each time step [5]. This technique transforms the nonlinear system into a linear system, where KF can be directly applied. While EKF is relatively simple, the linearization performed in each step neglects the higher-order derivatives, making the state estimations biased and overconfident. Specifically, estimations can be biased since $\operatorname{\mathbb{E}}(f(x))=f(\operatorname{\mathbb{E}}(x))$ is not guaranteed for a nonlinear function, and overconfidence arises when the actual variance is much higher than estimated. These problems are partially solved by the second-order extended Kalman filter (EKF2) [6], which approximates the nonlinear system by its second-order Taylor expansion. However, EKF2 requires the calculation of the Hessian matrix of each state transition function and measurement function, which adds non-trivial computational complexity to the algorithm. Additionally, EKF2 still sometimes gives inaccurate and overconfident estimations, so it is generally used less frequently than EKF.
Both EKF and EKF2 require the calculation of the Jacobian matrices of the state transition functions and measurement functions, which may not always be accessible in real-life systems. In 1997, Simon et al. introduced another extension of KF to nonlinear systems that requires no direct calculation of the Jacobian matrix and avoids the accuracy loss caused by linearization [7]. In later literature, this extension is usually referred to as the UKF [8]. UKF approximates the distribution of the states after nonlinear transformations by sampling several points (called sigma points) around the estimated mean of the states. These points go through the same transformation as the mean, and a more accurate estimation of the states’ distribution after the transformation can be obtained with the help of these extra sample points. In UKF, the sigma points are selected deterministically based on each state’s standard deviation. This technique guarantees that the sigma points are evenly distributed around the central point and can provide a more accurate state estimation. Similar ideas have also been used in several other nonlinear KFs, including cubature Kalman filter (CKF) [9], quadrature Kalman filter [10], square-root quadrature Kalman filter [11], etc. All these filters use deterministic ways to find several additional points around the central point (the estimated states from the previous iteration) and use those points to approximate the first two moments of the states after the nonlinear transformation. All these methods have some links with EKF2. As pointed out by Gustafsson et al. in [12], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the exact estimation as the EKF2. A disadvantage of these nonlinear KFs is that the number of sampling points is at least twice the number of states. Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
While different extensions of the KF use various ways to approximate the first two moments of the states after a nonlinear transformation, they all follow the same framework as in the linear KF. In such a framework, the state estimations in each iteration can be divided into two steps: Predict and Update. In the “predict” step, the state and covariance estimations are predicted using the state transition function and estimates from the previous iteration. Then, in the “update” step, the measurements correct the state predictions with a feedback gain (also known as the Kalman gain), which is designed to minimize the trace of the states’ covariance matrix after the update. After the states are updated, the state covariance matrix is also updated, ending the present iteration. Note that the covariance matrix will have the smallest possible trace after the update since that’s how the Kalman gain is selected.
While such a framework can guarantee the optimality of the state estimation for linear systems, it often leads to overconfident covariance estimation and less accurate state estimation when applied to nonlinear systems. The primary reason for such a phenomenon is the assumption that the Kalman gain can minimize the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions. In fact, as we will show later, the “update” step can sometimes make the state estimation of the nonlinear system less accurate, so the assumption about the optimality of the “update” step can be extremely harmful, as it makes the algorithm become overconfident in an inaccurate estimation.
The problem described above will become increasingly significant as the measurement noise decreases, and the measurements play a more critical role in the state estimation [13]. In some circumstances, it can even lead to a highly undesirable and counterintuitive phenomenon: the actual estimation error increases as the measurement noise decreases. The primary cause of this phenomenon is that the actual estimation errors drop much slower than the estimated standard deviations, and such overconfidence prevents the filter from correcting itself in later iterations. As the development of technology makes all types of sensors more and more accurate, we can reasonably predict that this problem will become increasingly severe if we stick to the conventional framework for nonlinear KFs.
Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence. Although it is possible to mitigate the problem by approximating the measurement function at higher orders (e.g., using EKF2 instead of EKF), these algorithms take a significantly longer computational time, and the problem will still, in general, exist. The lack of theoretical analysis for nonlinear KFs makes finding a general solution even harder. For those few existing theoretical analyses [12, 13, 14], the results are based on the assumption that the higher-ordered terms in the Taylor expansions of the measurement functions can be ignored. This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the ”update” step.
In this study, we proposed a new framework that solves the problem of overconfidence. The key idea of the new framework is to add a step that reevaluates the actual effect of the Kalman gain after the state update. With this additional step, the state covariance estimation can be much more accurate, and any unhelpful update can be withdrawn. The main contributions of this paper are:
  • To the best of our knowledge, we are the first to give theoretical proof that the conventional framework for nonlinear KFs, which has been used for around sixty years, has the problem of underestimating the trace of the state covariance matrix.
  • We proposed a new framework to solve the problem of inaccurate covariance estimation and validated the framework’s effectiveness by using both theoretical and empirical analysis.
  • The proposed new framework can be combined with almost any type of nonlinear KF and reduce the state estimation error by more than an order of magnitude in various applications when the measurement noise is low.
The rest of the paper is organized as follows. Section 2 conducts a rigorous theoretical analysis of the problem of the conventional framework for nonlinear KFs and introduces a new framework that solves the problem. Section 3 presents a comprehensive comparison between the two frameworks. Specifically, both frameworks are combined with four different types of nonlinear KFs and validated on five distinct applications. The conclusions are drawn in Section 4. The detailed simulation setup and the detailed algorithms we used for validation can be found in the appendix.
2 Method
2.1 General formulation of nonlinear Kalman filters
Consider a system with the following dynamics:
$x_{k}=f(x_{k-1},u_{k-1})+w_{k-1}$ (1)
where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow the zero-mean multivariate Gaussian distribution with covariance $Q_{k}$. The estimations of the states made in the previous step are denoted as $\hat{x}_{k-1|k-1}$. Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow the zero-mean multivariate Gaussian distribution with covariance $P_{k-1|k-1}$. Whichever nonlinear KF is used, the first step of each iteration is called “predict”, which predicts the states using (1) and the estimations made in the previous time step. The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k|k-1}$ are assumed to follow the zero-mean multivariate Gaussian distribution with covariance $P_{k|k-1}$.
The second step of each iteration for any nonlinear KF is called “update”, which corrects the predicted states based on the measurements. Suppose that the measurements of the states are as follows:
$z_{k}=h(x_{k})+v_{k}$ (2)
where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow the zero-mean multivariate Gaussian distribution with covariance $R_{k}$. Given (2) and the predicted states, the measurements can also be predicted using a particular method specific to the type of nonlinear KFs. The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow the zero-mean multivariate Gaussian distribution with covariance $P_{y}$. The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy}$. The differences between the predicted and actual measurements are called the measurement residuals $\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$. If we assume that the measurement noises are independent from the predicted states, then $\tilde{y}_{k}$ should have zero mean, and their covariance $S_{k}$ can be calculated by:
$S_{k}=P_{y}+R_{k}$ (3)
KFs use the following equation to update the state estimation:
$\hat{x}_{k|k}=\hat{x}_{k|k-1}+K\tilde{y}_{k}$ (4)
where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (3,4), the covariance of $\hat{x}_{k|k}$ can be calculated as:
$P_{k|k}=P_{k|k-1}+KS_{k}K^{T}-P_{xy,k}K^{T}-KP_{xy,k}^{T}$ (5)
KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of (5), we notice that the trace is minimized when:
$K=K_{op}=P_{xy}S_{k}^{-1}$ (6)
where $K_{op}$ is the theoretical optimal value of the Kalman gain. When (6) is satisfied, (5) becomes:
$P_{k|k}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (7)
Equations (47) illustrate the key idea of KFs, which is to update the states in a way that minimizes the trace of the state covariance matrix after the update. However, it should be noted that these equations cannot be directly applied to systems with nonlinear measurement functions. The primary reason is that the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ are all state-dependent for these nonlinear systems. Since the actual states are unknown, the actual values for these covariance matrices are also unknown. In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that make the conventional framework for nonlinear KFs overconfident and inaccurate.
2.2 The conventional framework and its problem
While calculating $P_{y,k},S_{k}$, and $P_{xy,k}$ at the actual state is impossible, it is possible to approximate their values based on the predicted states. We denote these approximated covariance matrices as $P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$ since they are all estimated based on the prediction. With these approximations, the Kalman gain can be selected as:
$K=P_{xy,k|k-1}S_{k|k-1}^{-1}$ (8)
Although the detailed process of calculating $P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (5) with the estimated values. Namely, the conventional framework estimates the updated state covariance matrix by:
$P_{k|k}$ $=P_{k|k-1}-KS_{k|k-1}K^{T}$ (9)
$=P_{k|k-1}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$
On the other hand, the actual value of the updated state covariance matrix $P_{k|k,ac}$ can be calculated by substituting (8) into (5). Namely:
$P_{k|k,ac}=$ $P_{k|k-1}+P_{xy,k|k-1}S_{k|k-1}^{-1}S_{k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$ (10)
$-P_{xy,k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k}^{T}$
While guaranteeing that $P_{k|k,ac}=P_{k|k}$ holds for all combinations of $P_{xy,k|k-1}$ and $S_{k|k-1}$ is impossible, we would at least hope that $P_{k|k}$ is an unbiased estimation of $P_{k|k,ac}$ (i.e., $\operatorname{\mathbb{E}}(P_{k|k})=P_{k|k,ac}$). Unfortunately, we have the following theorem.
Theorem 1
Define $g(A,B)=\text{tr}(P_{k|k,ac})|_{P_{xy,k}=A,S_{k}=B}$. Suppose that $\operatorname{\mathbb{E}}(P_{xy,k|k-1})=P_{xy,k}$, $\operatorname{\mathbb{E}}(S_{k|k-1})=S_{k}$, and $P_{xy,k|k-1}$ and $S_{k|k-1}$ are independent from each other. When $S_{k|k-1}-S_{k}$ is small, $P_{k|k}$ calculated from (9) has the following characteristic:
$\operatorname{\mathbb{E}}[\text{tr}(P_{k|k})]$ $=\operatorname{\mathbb{E}}[g(P_{xy,k|k-1},S_{k|k-1})]$ (11)
$<g(P_{xy,k},S_{k})=\text{tr}(P_{k|k,ac})$
Proof 2.2.
Denote $P_{xy,k|k-1}=P_{xy,k}+\Delta P$, $S_{k|k}=S_{k}+\Delta S$. According to the definition, $\operatorname{\mathbb{E}}(\Delta P)=O,\operatorname{\mathbb{E}}(\Delta S)=O$. Therefore, we have:
$g(P_{xy,k|k-1},S_{k|k-1})-g(P_{xy,k},S_{k|k-1})$ (12)
$=$ $\text{tr}[P_{xy,k|k-1}(S_{k}+\Delta S)^{-1}\Delta S(S_{k}+\Delta S)^{-1}P_{xy,k|k-1}^{T}]$
$-\text{tr}[\Delta PS_{k|k-1}^{-1}(P_{xy}+\Delta P)^{T}+(P_{xy}+\Delta P)S_{k|k-1}^{-1}\Delta P^{T}]$
Note that $(S_{k}+\Delta S)^{-1}$ can be expanded as the Neumann series. When $\Delta S$ is small, we have the following approximation:
$(S_{k}+\Delta S)^{-1}\Delta S(S_{k}+\Delta S)^{-1}$ (13)
$\approx$ $(S_{k}^{-1}-S_{k}^{-1}\Delta SS_{k}^{-1})\Delta S(S_{k}^{-1}-S_{k}^{-1}\Delta SS_{k}^{-1})$
$\approx$ $S_{k}^{-1}\Delta SS_{k}^{-1}-2S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}$
Since $\Delta P$ and $\Delta S$ are independent from each other and have zero mean, we have:
$\operatorname{\mathbb{E}}[g(P_{xy,k|k-1},S_{k|k-1})-g(P_{xy,k},S_{k})]$ (14)
$\approx$ $\operatorname{\mathbb{E}}[\text{tr}(-2P_{xy,k|k-1}S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}P_{xy,k|k-1}^{T}$
$-2\Delta PS_{k|k-1}^{-1}\Delta P^{T})]<0$
Note that the derivation of (14) utilizes the fact that both $S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}$ and $S_{k|k-1}^{-1}$ are positive semi-definite and symmetrical.
Theorem 1 suggests that using (9) to estimate the state covariance matrix will lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values. Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xy,k}$ are related to the actual state, which is unknown. Note that this problem exists for all types of nonlinear KFs as long as (9) is used to update the state covariance estimation. We can further prove that the magnitude of such overconfidence is related to the accuracy of $S_{k|k-1}$ and $P_{xy,k|k-1}$. Namely, we have the following theorem.
Theorem 2.3.
Suppose that $P_{xy,k|k-1}=P_{xy}+t_{1}P,S_{k|k-1}=S_{k}+t_{2}S$, where $t_{1}\geq 0,t_{2}\geq 0$, $P$ is a random matrix with zero mean, and $S$ is a random symmetrical matrix with zero mean. Suppose that $P$ and $S$ are independent from each other, and $t_{2}S$ is very small. If we define the overconfidence in state covariance estimation as:
$\Delta p_{k|k}(t_{1},t_{2})=\operatorname{\mathbb{E}}[g(P_{xy,k},S_{k|k})-g(P_{xy,k|k-1},S_{k|k-1})]$ (15)
Then, for any $t_{a}\geq t_{b}$ and $t_{c}\geq t_{d}$, we have:
$\Delta p_{k|k}(t_{a},t_{c})\geq\Delta p_{k|k}(t_{b},t_{d})$ (16)
Proof 2.4.
Let $\Delta P=t_{1}P,\Delta S=t_{2}S$. (14) can be rewritten as:
$\Delta p_{k|k}(t_{1},t_{2})=\operatorname{\mathbb{E}}[g(P_{xy,k},S_{k|k})-g(P_{xy,k|k-1},S_{k|k-1})]$ (17)
$=$ $\operatorname{\mathbb{E}}[\text{tr}(2t_{2}^{2}(P_{xy,k}+t_{1}P)S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}(P_{xy,k}+t_{1}P)^{T}$
$+2t_{1}^{2}PS_{k|k-1}^{-1}P^{T})]$
$=$ $2\operatorname{\mathbb{E}}[\text{tr}(t_{2}^{2}P_{xy,k}S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P_{xy,k}^{T}$
$+t_{1}^{2}t_{2}^{2}PS_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P^{T}+t_{1}^{2}PS_{k|k-1}^{-1}P^{T})]$
Since $P_{xy,k}S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P_{xy,k}^{T}$, $PS_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P^{T}$, and $PS_{k|k-1}^{-1}P^{T}$ are all positive semi-definite, their traces should all be greater or equal to zero. Therefore, (16) holds for any $t_{a}\geq t_{b}$ and $t_{c}\geq t_{d}$.
Theorem 2.3 suggests that increasing the noise in $S_{k|k-1}$ and $P_{xy,k|k-1}$ proportionally can make the state covariance estimation less accurate. In other words, the problem of the conventional framework is likely to be more severe when $S_{k}$ and $P_{xy}$ change more dramatically with respect to the states.
In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal. However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the covariance matrix after the update. The reason is that the states will move to a new location after the update, and the previous approximations may be inaccurate. An example of such a problem is shown in Fig. 1, where we used an EKF to estimate a plane’s position. In the example, the state is the plane’s position along the x-axis, and the measurement function is the altitude of the ground beneath the plane, which is a nonlinear function of the state. Due to the inaccurate approximation of the measurement function around the predicted state (the dashed purple line, “Approx. 1”, in Fig. 1), the EKF overestimates the position of the plane, and the absolute state estimation error $|x_{k|k}-x_{k}|$ becomes higher after the “update” step. Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step. Such a contradiction suggests that the effectiveness of the Kalman gain is overestimated. This is visualized by the narrow yellow distribution in the “update” step in Fig. 1), which does not cover the true state.
Figure 1: The basic idea of the new framework for nonlinear KFs. The additional step we propose re-approximates the nonlinear measurement function around the updated states, making the final covariance estimation more accurate.
2.3 The proposed new framework for nonlinear Kalman filters
Once we fundamentally understand the problem of the conventional framework, the solution is simple. We can redo the approximation after the update and use the newly approximated measurement function to update the covariance estimation. In our new framework, this new step is called “recalibrate”. As shown in Fig. 1, with this additional step, the covariance matrix may not have the smallest possible trace after the update; instead, it will honestly reflect the effect of the Kalman gain on the state covariance matrix. This is visualized by the wider yellow distribution in the “recalibrate” step, which now covers the true state.
Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $x_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$. We denote these re-approximated covariance matrices as $P_{y,k|k},S_{k|k}$, and $P_{xy,k|k}$ since they are all estimated based on the updated states. In the new framework, the state covariance matrix is updated as follows:
$P_{k|k}=$ $P_{k|k-1}+KS_{k|k}K^{T}-P_{xy,k|k}K^{T}-KP_{xy,k|k}^{T}$ (18)
$=$ $P_{k|k-1}+P_{xy,k|k-1}S_{k|k-1}^{-1}S_{k|k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$
$-P_{xy,k|k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k|k}^{T}$
Since the new framework selects the same Kalman gain as the conventional framework, the actual value of the updated state covariance matrix $P_{k|k,ac}$ can still be calculated by (10). By comparing (18) and (10), we have the following theorem:
Theorem 2.5.
Assume that $\operatorname{\mathbb{E}}(P_{xy,k|k-1})=\operatorname{\mathbb{E}}(P_{xy,k|k})=P_{xy,k}$ and $\operatorname{\mathbb{E}}(S_{k|k-1})=\operatorname{\mathbb{E}}(S_{k|k})=S_{k}$. If $P_{xy,k|k-1}$, $P_{xy,k|k}$, $S_{k|k-1}$, and $S_{k|k}$ are all independent from each other, $P_{k|k}$ calculated in (18) is an unbiased estimation of $P_{k|k,ac}$. Namely, we have:
$\operatorname{\mathbb{E}}(P_{k|k})=P_{k|k,ac}$ (19)
Proof 2.6.
With the assumptions made in Theorem 2.5, we have:
$\operatorname{\mathbb{E}}(P_{k|k}-P_{k|k,ac})$ (20)
$=$ $\operatorname{\mathbb{E}}[P_{xy,k|k-1}S_{k|k-1}^{-1}(S_{k|k}-S_{k})S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}]$
$-\operatorname{\mathbb{E}}[(P_{xy,k|k}-P_{xy,k})S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}]$
$-\operatorname{\mathbb{E}}[P_{xy,k|k-1}S_{k|k-1}^{-1}(P_{xy,k|k}-P_{xy,k})^{T}]$
$=$ 00 0
Theorem 2.5 suggests that the proposed new framework can prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other. This desirable feature comes from our replacement of (9) with (18), which highlights the fundamental idea of the new framework. That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately. In fact, in the new framework, the trace of the state covariance matrix may increase after the state update and recalibration if the system is strongly nonlinear and the predicted states have high variances, just like the example shown in Fig. 1. This phenomenon means that the update fails to make the estimation more accurate and should be withdrawn. To address this situation, we also add a “back out” step to the framework, which allows the state estimator to back out of the update when the update is unhelpful. The pseudo-code of the new framework is shown in Algorithm 1. Note that the new framework degrades to the conventional framework when the “recalibrate” and “back out” steps are replaced with (9).
Algorithm 1 The proposed framework for nonlinear Kalman filters
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     Estimate $\boldsymbol{\hat{x}}_{k|k-1}$ and $\boldsymbol{P}_{k|k-1}$
8:    (based on $\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{P}_{k-1|k-1},\boldsymbol{f},\boldsymbol{u}_{k},$ and $\boldsymbol{Q}_{k}$)
9:    Update:
10:     Estimate $\boldsymbol{\hat{y}}_{k|k-1},\boldsymbol{P}_{xy,k|k-1}$ and $\boldsymbol{P}_{y,k|k-1}$
11:    (based on $\boldsymbol{\hat{x}}_{k|k-1},\boldsymbol{P}_{k|k-1},$ and $\boldsymbol{h}$)
12:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
13:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
14:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
15:    Recalibrate:
16:     Estimate $\boldsymbol{P}_{xy,k|k}$ and $\boldsymbol{P}_{y,k|k}$
17:    (based on $\boldsymbol{\hat{x}}_{k|k},\boldsymbol{P}_{k|k-1},$ and $\boldsymbol{h}$)
18:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
19:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
20:    Back out:
21:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
22:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
23:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
24:     end if
25:end for
26:where, $\boldsymbol{\hat{x}}$ is the estimated states, $\boldsymbol{P}$ is the states’ covariance matrix, $\boldsymbol{\hat{y}}$ is the estimated measurements, $\boldsymbol{P}_{xy}$ is the covariance between the states and the measurements, $\boldsymbol{P}_{y}$ is the covariance of the estimated measurements, $\boldsymbol{S}$ is the innovation (or residual) covariance, $\boldsymbol{K}$ is the Kalman gain. The subscript $k|k-1$ represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
It is worth mentioning that when the measurement functions are all linear, they would require no approximation. Or, to be more precise, the approximation of the measurement functions would become independent of the states. In this case, the estimations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
To give readers an intuitive understanding of the effect of the “recalibrate” and “back out” steps, in Fig. 2, we showed an example of using different types of nonlinear KFs to do the state update given the measurement $y=x^{3}/3-x^{2}/8-x+1.5383=0$. The measurement function is represented by the thick gray curve in the figure. The predicted state is $x_{1|0}=0$, its variance is $1.5^{2}$, and the standard deviation of the measurement is $\sigma_{y}=0.01$. This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is on the left of the predicted state. The state and covariance estimations made by different algorithms are summarized in different rows in the figure. Generally, the conventional framework always underestimates the actual state estimation error, which can be visualized by the narrow standard deviation estimations in Fig. 2 that cannot capture the true state. By contrast, the state covariance estimation made by the new framework is much more accurate. Additionally, the new framework is aware that the state estimation becomes less accurate after the update because the estimated variance increases. Therefore, after the “recalibrate” step, the “back out” step will be triggered, and the new framework will withdraw the update and use the predicted state and the predicted state variance as the final output.
Figure 2: The comparison of the covariance estimation after the state update and recalibration using different frameworks and different types of nonlinear Kalman filters.
The new framework described above may remind the readers of the iterated Kalman filter (IKF), which also re-approximates the measurement function after calculating the Kalman gain [15, 16, 17]. However, IKF is fundamentally different from the new framework in three ways. Firstly, IKF redoes the approximation and recalculates the Kalman gain iteratively until convergence. However, such a convergence is not guaranteed, so the algorithm is less robust and often requires finetuning the step size of each iteration [18]. Even if the iteration converges, the convergence may take a very long time. By contrast, the new framework requires no iteration, and the re-approximation is only done once. Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 2), which is only possible for EKF and EKF2. Although a few attempts have been made to extend IKF to UKF and CKF [16, 17, 19, 20], a consensus on how to do this robustly has not been reached. Therefore, it is not easy to implement IKF on all types of nonlinear KFs while guaranteeing that the estimator can converge in most cases. Our framework, on the other hand, does not have this problem since there is no convergence issue. Thirdly, and most importantly, the basic idea behind IKF is still to find the optimal Kalman gain and not to estimate the covariance matrix accurately. Namely, IKF always updates the states and covariance matrix simultaneously using (9), resulting in overconfident state estimations. When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence, according to Theorem 2.3. In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.
Algorithm 2 The iterated extended Kalman filter with the old framework
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
8:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})$
9:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}+\boldsymbol{Q}_{k}$
10:    Update:
11:     $\boldsymbol{x}_{0}=\boldsymbol{\hat{x}}_{k|k-1}$
12:     for $i=1$ to 1000 do
13:         $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{x}_{i-1}}$
14:         $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{x}_{i-1})+\boldsymbol{H}_{k|k-1}(\boldsymbol{x}_{i-1}-\boldsymbol{\hat{x}}_{k|k-1})$
15:         $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
16:         $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
17:         $\boldsymbol{S}_{k|k-1,i}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
18:         $\boldsymbol{K}_{i}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1,i}^{-1}$
19:         $\boldsymbol{x}_{i}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}_{i}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
20:         if $i>1$ and $||\boldsymbol{x}_{i}-\boldsymbol{x}_{i-1}||>||\boldsymbol{x}_{i-1}-\boldsymbol{x}_{i-2}||$ then
21:              $\boldsymbol{S}_{k|k-1,i}=\boldsymbol{S}_{k|k-1,i-1}$
22:              $\boldsymbol{K}_{i}=\boldsymbol{K}_{i-1}$
23:              $\boldsymbol{x}_{i}=\boldsymbol{x}_{i-1}$
24:              break
25:         end if
26:         if $\max_{j}|1-\boldsymbol{x}_{i}(j)/\boldsymbol{x}_{i-1}(j)|<0.001$ then
27:              break
28:         end if
29:     end for
30:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{S}_{k|k-1,i}$
31:     $\boldsymbol{K}=\boldsymbol{K}_{i}$
32:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{x}_{i}$
33:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}-\boldsymbol{K}\boldsymbol{S}_{k|k-1}\boldsymbol{K}^{T}$
34:end for
35:Note that rows 16–21 are added to prevent the state estimations from diverging. Namely, when the sign of divergence is detected (row 16), the newest updates of the states, Kalman gain, and measurement residual covariance matrix are withdrawn (rows 17–19). If such a withdrawal is not made, we found in the simulation that the estimation accuracy can be much lower. This characteristic also suggests that the algorithm is not very robust and stable.
Since IKF also uses (9) to update the state covariance matrix, it can also be considered as another framework for nonlinear KFs that is similar to the conventional framework. A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy? Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states ($P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$) are independent of the approximations made at the updated states ($P_{y,k|k},S_{k|k}$, and $P_{xy,k|k}$). However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2.5. For example, the IEKF algorithm shown in Algorithm 2 approximates the measurement function iteratively at $x_{i-1}$ (rows 9–13). When the algorithm converges, in the final iteration, we have $x_{i-1}\approx x_{i}=\hat{x}_{k|k}$, meaning that the approximation of the measurement functions in the final iteration will stay the same after the state update. In fact, to the best of our knowledge, IKF-related algorithms are the only existing types of nonlinear KFs that can not get large benefits from the proposed framework. Nevertheless, for any type of nonlinear KF that can be combined with IKF, its accuracy can be greatly improved when it is combined with our new framework instead. In this sense, we can still say that all existing types of nonlinear KFs can benefit from the proposed framework.
3 Results
To demonstrate the effectiveness of the proposed framework, we validated it in five different applications suitable for nonlinear KFs. A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix 5. For each scenario, we combined the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF. We also implemented the IEKF for comparison. The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 36 in Appendix 6. All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups). To guarantee the fairness of the comparison and reduce the effect of randomness, each KF is simulated 10,000 times for each measurement noise setup, and all the KFs use the same random seed during the simulation. This means that the random initial states’ errors, random process noises, and random measurement noises are the same for the $i^{th}$ simulation of EKF and the $i^{th}$ simulation of UKF. In other words, all these random noises were generated for the 10,000 runs a priori and recalled for simulating each estimation method.
Table 1: A brief description of the five applications of nonlinear Kalman filters investigated in this paper.
Applications
3D target
tracking
Terrain-referenced
navigation
Synchronous generator
state estimation
Pendulum state
estimation
Battery state
estimation
Number of states 6 2 4 2 3
Linear state transition function? Yes Yes No No No
Number of measurements 2 1 1 1 1
Linear measurement functions? No No No No No
Number of inputs 3 2 3 1 1
Number of iterations 30 100 100 100 180
The root mean squared error (RMSE) of the state estimations given by different nonlinear KFs are shown in Figs. 37. In all figures, different types of nonlinear KFs are represented by different colors and markers, and the old and new frameworks are represented by dotted lines and solid lines, respectively. The x-axis represents the standard deviations of the measurements. The root mean squared error is calculated from the estimation errors in 10,000 simulations. The estimation error is calculated as the difference between the estimated state and the true state at the end of the final iteration of each simulation. Note that we only showed the RMSE of two representative states for systems with more than two states. For example, in 3D target tracking, the six states are the object’s speed and position along the three axes, and Fig. 3 examines the object’s speed and position estimation accuracy along the x-axis. As can be seen from the five figures, the simulation results are very similar to each other. Therefore, only the results of the first task, 3D target tracking, are analyzed in depth in this section.
Figure 3: The root mean squared error of the state estimations under different measurement noise setups (3D target tracking).
Figure 4: The root mean squared error of the state estimations under different measurement noise setups (terrain-referenced navigation).
Figure 5: . The root mean squared error of the state estimations under different measurement noise setups (synchronous generator state estimation).
Figure 6: The root mean squared error of the state estimations under different measurement noise setups (pendulum state estimation).
Figure 7: The root mean squared error of the state estimations under different measurement noise setups (battery state estimation).
As shown in Fig. 37, the proposed framework (solid lines) significantly improves the accuracy of all types of nonlinear KFs. The improvement becomes stronger as the measurement noise becomes smaller. This result aligns with our expectations because the proposed framework only changes how the measurements are used in the nonlinear KF, and the estimations will rely more on the measurements instead of system knowledge as the measurement noise decreases. For all these applications, when the standard deviations of the measurements are small, the RMSE of all four types of nonlinear KFs can usually be reduced by more than an order of magnitude, which is a significant improvement.
Another interesting observation of Figs. 37 is that the RMSE of EKF2, UKF, and CKF becomes quite similar after the new framework is used. This phenomenon can be explained by the fact that all three methods consider the first and second-order terms in the Taylor expansion of the nonlinear state transition and measurement functions in certain ways [12]. On the other hand, EKF is generally slightly worse than the other three types of nonlinear KFs because it only considers the first-order terms in the Taylor expansion.
By comparing the performance of IEKF and EKF (with and without the new framework) in Figs. 37, we can see that IEKF does perform better than EKF with the old framework, yet it is almost always worse than EKF with the new framework despite requiring longer computational time. As we mentioned in the Introduction, the main reason for such a difference is that IEKF still gives overconfident covariance estimation, especially when it fails to converge, which happens quite often in practice.
As we mentioned in the Introduction, the fundamental idea of the proposed framework is to make a more accurate covariance estimation by re-approximating the measurement functions after the state update. To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the system) in the old and new framework in Fig. 8. In the figure, for each type of nonlinear KFs (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation. Apparently, when the standard deviations of the two measurements are lower than 0.1 meters, the conventional framework always significantly underestimates the covariance, sometimes by several orders of magnitude, regardless of the type of nonlinear KFs used. On the other hand, for the new framework, only the RMSE of EKF is underestimated, and the covariance estimation of other types of nonlinear KFs matches the actual value very well. This, again, can be explained by the fact that EKF neglects the second and higher-order information of the measurement functions, so its covariance estimation is less accurate than other methods. Nevertheless, the covariance estimation of EKF is still much more precise when combined with the new framework.
Figure 8: The actual and estimated root mean squared error of the x-axis position estimation.
Another interesting observation of Fig. 8 is that the estimated RMSE of the x-axis position (the dashed lines) is very similar for different combinations of frameworks and KFs. This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction – to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
Another important metric for evaluating the performance of different filters is the convergence characteristics. We visualized this metric by fixing the standard deviations of the two measurements to 0.01 m and plotted the RMSE of the state estimations after different numbers of iterations. The results are shown in Fig. 9. From the figure, we can see that all types of nonlinear KFs can converge faster after being combined with the new framework. This result suggests that the proposed framework is very suitable for applications that require high convergence speed.
Figure 9: The root mean squared error of the state estimations after different numbers of iterations.
Since the extra steps introduced in the new framework take time, it is also important to compare the computational time of the two frameworks. The results are summarized in Fig. 10. Note that the average runtime shown in the figure is calculated from over 100,000 simulations run on the same computer. Specifically, for each combination of nonlinear KF and framework, we recorded the method’s runtime whenever it was used and calculated the average. After the average runtime is calculated, the value is normalized with respect to the average runtime of the conventional EKF for the same application. In other words, the normalized average runtime of “EKF (old)” always equals one in Fig. 10.
Figure 10: Comparison of the computational time of different methods and frameworks.
From the result, we can see that the new framework generally increases the computational time by 10–90%, varying for different nonlinear systems. Namely, when the system has linear state transition functions and nonlinear measurement functions (which is true for the first two applications in the table), the “update” step should be accountable for most of the computational time in the conventional framework and re-approximating the measurement functions can almost double the runtime in the worst case. The normalized runtime of the new framework is related to the type of nonlinear KFs as well. In UKF and CKF, one of the most time-consuming steps is to calculate the square root of the state covariance matrix. This step is respectively done one, two, two, and three times in each iteration of UKF (old framework), UKF (new framework), CKF (old framework), and CKF (new framework). Therefore, the computational time increase (in percentage) brought by the new framework is longer for UKF than for CKF. Noticing that the new framework can reduce the estimation error by more than an order of magnitude when the measurement noise is small, we can say that the extra computational complexity is generally worth it. Another thing to note is that the combination of EKF and the new framework requires significantly less computational time than the combination of any other nonlinear KFs and the old framework. At the same time, the former’s accuracy is generally higher than the latter in the applications we examined in this paper. Therefore, even in situations where the computational time is of major concern, the proposed framework can still be helpful because the new EKF can outperform the old EKF2, UKF, and CKF, in addition to saving a lot of computational time.
4 Conclusion
In this paper, we proposed a new framework that can be applied to all types of nonlinear KFs. The new framework re-approximates the measurement functions after the state update, solving the problem of overconfident state covariance estimation, which is an inherent problem in the conventional framework for nonlinear KFs. The new framework was tested in five different applications, reducing the estimation errors by more than an order of magnitude when the measurement noise is low and significantly reducing the convergence time of the algorithm. Considering that all types of sensors will become increasingly more accurate in the future, the proposed framework is expected to bring greater benefits for all kinds of nonlinear KFs over time.
The main limitation of this work is that the proposed framework can only be used in discrete-time nonlinear KFs. For continuous-time nonlinear KFs, the state estimations change continuously over time, so the problem of overconfident covariance estimation does not exist. However, all digital sensors in real life give discrete-time measurements. Even though it is possible to make the signals continuous through interpolation, the continuous-time state estimation problem sometimes includes several partial differential equations, so it is hard and takes a long time to solve [21]. Therefore, most applications of nonlinear KFs in real life use the discrete-time form, and they can all benefit from the new framework proposed in this paper.
{appendices}
5 The detailed simulation setup
This paper uses five different applications of nonlinear KFs to test the effectiveness of the new framework: 3D target tracking, terrain-referenced navigation, synchronous generator state estimation, pendulum state estimation, and battery state estimation.
Figure 11: Examples of nonlinear systems used in this paper for validation. (a) 3D object tracking. (b) Terrain-referenced navigation. (c) Pendulum state estimation. (d) Battery state estimation.
5.1 3D target tracking
This example is adapted from Example 3.3.1 in [22]. As shown in Fig. 11(a), a target moves in 3D space, and two sensors are used to track its location. The first sensor is at the origin, and the second sensor moves in a circle. The second sensor’s location $(x_{s1},x_{s2},x_{s3})$ at step $k=1,2,\ldots,30$ is represented by:
$\begin{cases}x_{s1,k}=20+20\cos{\big{(}\frac{(k-1)\pi}{15}\big{)}}\\ x_{s2,k}=20+20\sin{\big{(}\frac{(k-1)\pi}{15}\big{)}}\\ x_{s3,k}=0\end{cases}$ (21)
The states are the target’s positions $(x_{1},x_{2},x_{3})$ and speeds $(v_{1},v_{2},v_{3})$ along the three axes. The measurements $y_{1},y_{2}$ are the distances between the target and the two sensors. The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (22,23). Note that the noises are omitted in the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ v_{1,k}\\ v_{2,k}\\ v_{3,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+v_{1,k-1}\\ x_{2,k-1}+v_{2,k-1}\\ x_{3,k-1}+v_{3,k-1}\\ v_{1,k-1}\\ v_{2,k-1}\\ v_{3,k-1}\end{bmatrix}$ (22)
$\begin{bmatrix}y_{1,k}\\ y_{2,k}\end{bmatrix}=\begin{bmatrix}\sqrt{\sum_{i=1}^{3}x_{i,k}^{2}}\\ \sqrt{\sum_{i=1}^{3}(x_{i,k}-x_{si,k})^{2}}\end{bmatrix}$ (23)
The true initial states of the target are $x_{1,0}=10,x_{2,0}=-10,x_{3,0}=50,v_{1,0}=1,v_{2,0}=2,v_{3,0}=0$; the initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are 100, 100, 100, 0.01, 0.01, and 0.01. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are 0, 0, 0, $10^{–6}$, $10^{–6}$, and $10^{–6}$.
5.2 Terrain-referenced navigation
This example is adapted from Example 3.3.2 in [22]. As shown in Fig. 11(b), a plane flies above an uneven landscape. The states are the planes’ positions along the two axes, denoted as $x_{1}$ and $x_{2}$. The measurement $y_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements. The plane knows the contour map in advance, so the terrain’s elevation can tell the aircraft its possible locations. However, this task is difficult because all the points on a contour line have the same elevation, and system knowledge must be combined for accurate state estimations. The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (24, 25). Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+0.5\\ x_{2,k-1}\end{bmatrix}$ (24)
$y_{k}=\sin{\sqrt{\big{(}\frac{x_{1,k}}{40}\big{)}^{2}+\big{(}\frac{x_{2,k}}{40}\big{)}^{2}}}$ (25)
The true initial states of the target are $x_{1,0}=10$ km, $x_{2,0}=10$ km; the initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are both 1 km2. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are both 0.25 m2.
5.3 Synchronous generator state estimation
This example is adapted from a research paper [23]. The system model is precisely the same as in the paper. The states of the system are the rotor angle $\delta$, rotor speed $\Delta\omega$, the q-axis component of the voltage $e^{\prime}_{q}$, and the d-axis component of the voltage $e^{\prime}_{d}$. For simplicity, these states are denoted as $x_{1}$ to $x_{4}$. The measurement $y_{k}$ is electrical output power. The system’s inputs include the mechanical input torque $T_{m}$, the steady-state internal voltage of the armature $E_{fd}$, and the terminal bus voltage $V_{t}$. For simplicity, these inputs are denoted as $u_{1}$ to $u_{3}$. The time interval is set to $\Delta t=0.1$ ms. The input profiles are the same as in the paper, which are:
$\begin{cases}u_{1,k}=0.8\\ u_{2,k}=2.11+0.0002k\\ u_{3,k}=1.002\end{cases}$ (26)
All the model parameters are selected to be the same as in [23]. The system’s state transition and measurement functions can be written as (27, 28). Note that the noises are omitted from the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ x_{4,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}\\ x_{2,k-1}\\ x_{3,k-1}\\ x_{4,k-1}\end{bmatrix}$ (27)
$+\begin{bmatrix}377x_{2,k-1}\Delta t\\ \frac{\Delta t}{13}\big{[}u_{1,k-1}-\frac{u_{3,k-1}x_{3,k-1}\sin{x_{1,k-1}}}{0.375}\big{]}\\ \frac{\Delta t}{0.131}[u_{2,k-1}-x_{3,k-1}]\\ \frac{-x_{4,k-1}\Delta t}{0.0131}\end{bmatrix}$
$+\begin{bmatrix}0\\ \frac{\Delta t}{13}\big{[}0.9215u_{3,k-1}^{2}\sin(2x_{1,k-1})-0.05x_{2,k-1}\big{]}\\ \frac{-4.4933\Delta t}{0.131}(x_{3,k-1}-u_{3,k-1}\cos{x_{1,k-1}})\\ \frac{0.6911u_{3,k-1}\sin{x_{1,k-1}}\Delta t}{0.0131}\end{bmatrix}$
$y_{k}=\frac{u_{3,k-1}x_{3,k}\sin{x_{1,k}}}{0.375}+0.9215u_{3,k-1}^{2}\sin(2x_{1,k})$ (28)
The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$. The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{–4}$, $10^{–10}$, $10^{–4}$, and $10^{–4}$. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are $10^{–10}$, $10^{–16}$, $10^{–10}$, and $10^{–10}$.
5.4 Pendulum state estimation
As shown in Fig. 11(c), one end of the pendulum is fixed to the ceiling with a thin rope. The system’s states are the pendulum’s angular position $\theta$ and angular speed $\omega$. The measurement $y_{k}$ is the horizontal component of tension in the rope. The system input is the external torque exerted on the pendulum, which is zero in this example. The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is $g=9.8$m/s2. The time interval is set to $\Delta t=0.01$s. According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (29,30). Note that the noises are omitted in the equations.
$\begin{bmatrix}\omega_{k}\\ \theta_{k}\end{bmatrix}=\begin{bmatrix}\omega_{k-1}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \theta_{k-1}+\omega_{k-1}\Delta t\end{bmatrix}$ (29)
$y_{k}=mg\cos{\theta_{k}}\sin{\theta_{k}}+ml\omega_{k}^{2}\sin{\theta_{k}}$ (30)
The true initial states of the system are $\omega_{0}=0,\theta_{0}=\pi/4$. The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are both $(\pi/18)^{2}$. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal; its diagonal values are $10^{-10}$ and 0.
5.5 Battery state estimation
This example is adapted from a research paper [24]. As shown in Fig. 11(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$. A battery’s two most important states are its state of charge (SOC) and state of health (SOH). The SOC is the ratio of the battery’s remaining capacity to its present maximum capacity, and the SOH is the ratio of the battery’s present maximum capacity to its initial maximum capacity. In general, SOC describes the battery’s energy level, while SOH describes the battery’s aging level. The states of this system are the SOC, SOH, and capacitor voltage $U_{c}$. The measurement $y_{k}$ is the battery’s terminal voltage. The input is the charging current, which, in this example, is a 3-level square wave shown below:
$I_{k}=\begin{cases}-2,75\geq k>15\\ 2,165\geq k>105\\ 0,\text{else}\end{cases}$ (31)
The battery’s initial maximum capacity is $Q_{0}=1$ Ah. The parameters in the equivalent circuit model are chosen to be $R_{1}=0.01,R_{2}=0.05,1/(R_{2}C_{1})=0.008$. The OCV is a function of SOC and SOH, and it can be written as:
$OCV=\sum_{i=0}^{9}\big{(}\frac{SOH-0.8}{0.2}a_{100,i+1}+\frac{1-SOH}{0.2}a_{80,i+1}\big{)}$ (32)
where $a_{100}=$ [1390.38, -6961.31, 14760.31, -17230.92, 12055.71, -5162.75, 1330.60, -196.37, 15.60, 2.96], $a_{80}=$ [813.94, -4229.96, 9345.49, -11415.38, 8396.15, -3801.07, 1043.09, -165.29, 14.28, 2.96]. The time interval is set to $\Delta t=1$ s, and the total simulation time is 180 seconds. The system’s state transition and measurement functions can be written as (33,34). Note that the noises are omitted in the equations.
$\begin{bmatrix}SOC_{k}\\ U_{c,k}\\ SOH_{k}\end{bmatrix}=$ $\begin{bmatrix}SOC_{k-1}\\ U_{c,k-1}\exp{\big{(}\frac{-\Delta t}{R_{2}C_{1}}\big{)}}\\ SOH_{k-1}\end{bmatrix}$ (33)
$+\begin{bmatrix}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \big{[}1-\exp{\big{(}\frac{-\Delta t}{R_{2}C_{1}}\big{)}}\big{]}R_{2}I_{k-1}\\ 0\end{bmatrix}$
$y_{k}=OCV(SOC_{k},SOH_{k})+U_{c,k}+R_{1}I_{k}$ (34)
The true initial states of the system are $SOC_{0}=60\%,U_{c,0}=0,SOH_{0}=90\%$. The initial state covariance matrix $P_{0}$ is diagonal; its diagonal elements are 0.04, $10^{-10}$, and 0.01. The initial state estimations are 80%, 0, and 100%, respectively. The input current is noisy, and its standard deviation is 1 mA. This input noise is the only source of the process noises. Since the input current also appears in the measurement function, the actual measurement noise is the combination of the measurement noise of the voltmeter and the noise caused by the term $R_{1}I_{k}$ in the measurement function.
6 The detailed algorithm of various nonlinear Kalman filters used in the paper
The detailed algorithms of EKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 36. Note that all these algorithms degrade to the conventional framework when the ”recalibrate” and ”update” steps are replaced by (9).
Algorithm 3 The extended Kalman filter with the new framework
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
8:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})$
9:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}+\boldsymbol{Q}_{k}$
10:    Update:
11:     $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k-1}}$
12:     $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{\hat{x}}_{k|k-1})$
13:     $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
14:     $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
15:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
16:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
17:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
18:    Recalibrate:
19:     $\boldsymbol{H}_{k|k}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k}}$
20:     $\boldsymbol{P}_{xy,k|k}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
21:     $\boldsymbol{P}_{y,k|k}=\boldsymbol{H}_{k|k}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
22:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
23:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
24:    Back out:
25:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
26:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
27:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
28:     end if
29:end for
Algorithm 4 The second-order extended Kalman filter with the new framework
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, measurements number $n_{m}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
8:     $\boldsymbol{F}_{i,k}^{*}=\big{[}\frac{\partial^{2}f_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}\big{]}_{mn},i=1,\ldots,n_{x}$
9:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{F}_{i,k}^{*}\boldsymbol{P}_{k-1|k-1})\big{]}_{i}$
10:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}$
11:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{F}_{i,k}^{*}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{j,k}^{*}\boldsymbol{P}_{k-1|k-1})\big{]}_{ij}+\boldsymbol{Q}_{k}$
12:    Update:
13:     $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k-1}}$
14:     $\boldsymbol{H}_{i,k|k-1}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k|k-1}}\big{]}_{mn},i=1,\ldots,n_{m}$
15:     $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k-1}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{i}$
16:     $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
17:     $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
18:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k-1}^{*}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{j,k|k-1}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{ij}$
19:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
20:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
21:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
22:    Recalibrate:
23:     $\boldsymbol{H}_{k|k}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k}}$
24:     $\boldsymbol{H}_{i,k|k}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k|k}}\big{]}_{mn},i=1,\ldots,n_{m}$
25:     $\boldsymbol{P}_{xy,k|k}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
26:     $\boldsymbol{P}_{y,k|k}=\boldsymbol{H}_{k|k}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
27:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k}^{*}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{j,k|k}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{ij}$
28:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
29:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
30:    Back out:
31:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
32:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
33:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
34:     end if
35:end for
Algorithm 5 The unscented Kalman filter with the new framework
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$, hyperparameters $\alpha=10^{-3},\beta=2,\kappa=0$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:$\lambda=\alpha^{2}(n_{x}+\kappa)-n_{x}$
6:$W_{i}^{(c)}=W_{i}^{(m)}=1/(2n_{x}+2\lambda),i=1,\ldots,n_{x}$
7:$W_{0}^{(c)}=\lambda/(n_{x}+\lambda)+1-\alpha^{2}+\beta$
8:$W_{0}^{(m)}=\lambda/(n_{x}+\lambda)$
9:for every time step $k$ do
10:    Predict:
11:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k-1|k-1}}}]_{i},i=1,...,n_{x}$
12:     $\boldsymbol{\mathcal{X}}_{0,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}$
13:     $\boldsymbol{\mathcal{X}}_{i,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}+\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
14:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}-\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
15:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}=\boldsymbol{f}(\boldsymbol{\mathcal{X}}_{i,k-1},\boldsymbol{u}_{k-1}),i=0,1,\ldots,2n_{x}$
16:     $\boldsymbol{\hat{x}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{X}}_{i,k|k-1}$
17:     $\boldsymbol{P}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{X}}_{i,k|k-1}$
18:      $-\boldsymbol{\hat{x}}_{k|k-1}]^{T}+\boldsymbol{Q}_{k}$
19:    Update:
20:     $\boldsymbol{\mathcal{Y}}_{i,k|k-1}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
21:     $\boldsymbol{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
22:     $\boldsymbol{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
23:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
24:     $\boldsymbol{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{Y}}_{i,k|k-1}-\boldsymbol{\hat{y}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
25:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
26:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
27:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
28:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
29:    Recalibrate:
30:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k|k-1}}}]_{i},i=1,...,n_{x}$
31:     $\boldsymbol{\mathcal{X}}_{0,k|k}=\boldsymbol{\hat{x}}_{k|k}$
32:     $\boldsymbol{\mathcal{X}}_{i,k|k}=\boldsymbol{\hat{x}}_{k|k}+\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
33:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k}=\boldsymbol{\hat{x}}_{k|k}-\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
34:     $\boldsymbol{\mathcal{Y}}_{i,k|k}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
35:     $\boldsymbol{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{Y}}_{i,k|k}$
36:     $\boldsymbol{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k}-\boldsymbol{\hat{x}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
37:     $\boldsymbol{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
38:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
39:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
40:    Back out:
41:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
42:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
43:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
44:     end if
45:end for
Algorithm 6 The cubature Kalman filter with the new framework
1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
2:Initialization:
3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
5:for every time step $k$ do
6:    Predict:
7:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k-1|k-1}}}]_{i},i=1,...,n_{x}$
8:     $\boldsymbol{\mathcal{X}}_{i,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
9:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
10:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}=\boldsymbol{f}(\boldsymbol{\mathcal{X}}_{i,k-1},\boldsymbol{u}_{k-1}),i=1,\ldots,2n_{x}$
11:     $\boldsymbol{\hat{x}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}$
12:     $\boldsymbol{P}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}$
13:      $-\boldsymbol{\hat{x}}_{k|k-1}]^{T}+\boldsymbol{Q}_{k}$
14:    Update:
15:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k|k-1}}}]_{i},i=1,...,n_{x}$
16:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}=\boldsymbol{\hat{x}}_{k|k-1}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
17:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k-1}=\boldsymbol{\hat{x}}_{k|k-1}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
18:     $\boldsymbol{\mathcal{Y}}_{i,k|k-1}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
19:     $\boldsymbol{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
20:     $\boldsymbol{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
21:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
22:     $\boldsymbol{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{Y}}_{i,k|k-1}-\boldsymbol{\hat{y}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
23:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
24:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
25:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
26:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
27:    Recalibrate:
28:     $\boldsymbol{\mathcal{X}}_{i,k|k}=\boldsymbol{\hat{x}}_{k|k}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
29:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k}=\boldsymbol{\hat{x}}_{k|k}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
30:     $\boldsymbol{\mathcal{Y}}_{i,k|k}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
31:     $\boldsymbol{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{Y}}_{i,k|k}$
32:     $\boldsymbol{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k}-\boldsymbol{\hat{x}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
33:     $\boldsymbol{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
34:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
35:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
36:    Back out:
37:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
38:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
39:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
40:     end if
41:end for
Acknowledgment
This material is based upon work supported, in part, by the National Science Foundation under Grant No. 1847177.

### v10

@@
 The Kalman filter (KF) is a state estimation algorithm that optimally combines system knowledge and measurements to minimize the mean squared error of the estimated states.
-While KF was initially designed for linear systems, numerous extensions of it, such as extended Kalman filter (EKF), unscented Kalman filter (UKF), cubature Kalman filter (CKF), etc., have been proposed for nonlinear systems.
+While KF was initially designed for linear systems, numerous extensions of it, such as extended Kalman filter (EKF), unscented Kalman filter (UKF), cubature Kalman filter (CKF), etc., have been proposed for nonlinear systems over the last sixty years.
 Although different types of nonlinear KFs have different pros and cons, they all use the same framework of linear KF.
-Yet, according to what we found in this paper, the framework often gives overconfident and less accurate state estimations when the measurement functions are nonlinear.
-Therefore, in this study, we designed a new framework for nonlinear KFs and showed theoretically and empirically that the new framework estimates the states and covariance matrix more accurately than the old one.
-The new framework was tested on four different nonlinear KFs and five different tasks, showcasing its ability to reduce the estimation errors by several orders of magnitude in low-measurement-noise conditions, with only about a 10–90% increase in computational time.
-To the best of our knowledge, all existing types of nonlinear KFs can benefit from the new framework, and the benefit will increase as the sensors become more and more accurate in the future.
-As an example, EKF, the simplest nonlinear KF that was previously believed to work poorly for strongly nonlinear systems, can now perform better than the iterated extended Kalman filter (IEKF) and provide fast and fairly accurate state estimations with the help of the proposed new framework.
-The codes are available at https://github.com/Shida-Jiang/A-new-framework-for-nonlinear-Kalman-filters.
-{IEEEkeywords}
-Cubature Kalman filters, extended Kalman filters, Kalman filters, nonlinear filters, state estimation, unscented Kalman filters.
+Yet, according to our theoretical and empirical analysis, the framework tends to give overconfident and less accurate state estimations when the measurement functions are nonlinear.
+Therefore, in this study, we designed a new framework that can be combined with any existing type of nonlinear KFs and showed theoretically and empirically that the new framework estimates the states and covariance more accurately than the old one.
+The new framework was tested on four different nonlinear KFs and five different tasks, showcasing its ability to reduce estimation errors by several orders of magnitude in low-measurement-noise conditions.
+The codes are available at https://github.com/Shida-Jiang/A-new-framework-for-nonlinear-Kalman-filters
+keywords:
+Kalman filtering; nonlinear observer and filter design; extended Kalman filter; unscented Kalman filter; estimation theory; parameter and state estimation
+thanks: This paper was not presented at any IFAC meeting. Corresponding author Shida Jiang Tel. +1 510-944-9339.
 1 Introduction
@@
 It is assumed that both the state transition and measurement functions contain some additive noise.
-The KF optimally combines the system knowledge and measurements to minimize the sum of the states’ variances when both process and measurement noise follow a normal (Gaussian) distribution.
+The KF optimally combines the system knowledge and measurements to minimize the expected value of the mean-squared error of the state estimation.
+Additionally, when the process noise, measurement noise, and state estimation errors all follow a multivariate normal (Gaussian) distribution, the state estimation given by the Kalman filter is also the maximum likelihood estimation [1].
+Due to these desirable features, KF is often referred to as the optimal linear state estimator.
 Since its birth in about 1960, KF has achieved massive success in navigation [2], motion control [3], statistics [4], signal processing [5], etc.
@@
 As pointed out by Gustafsson et al.
-in [12], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the exact estimation as the EKF2.
-A disadvantage of these nonlinear KFs is that the number of sampling points is at least twice the number of states.
+in [13], when the sigma points are infinitely close to the center, and the hyperparameters in the UKF are set to specific values, UKF will give the same state estimation as the EKF2, although the covariance estimation can still be a bit different.
+A disadvantage of these nonlinear KFs is that the number of sampling points is usually twice the number of states.
 Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
 While different extensions of the KF use various ways to approximate the first two moments of the states after a nonlinear transformation, they all follow the same framework as in the linear KF.
+In this paper, this framework is referred to as “the conventional framework” or “the old framework”.
 In such a framework, the state estimations in each iteration can be divided into two steps: Predict and Update.
@@
 While such a framework can guarantee the optimality of the state estimation for linear systems, it often leads to overconfident covariance estimation and less accurate state estimation when applied to nonlinear systems.
-The primary reason for such a phenomenon is the assumption that the Kalman gain can minimize the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions.
+The primary reason for such a phenomenon is the assumption that the Kalman gain minimizes the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions.
 In fact, as we will show later, the “update” step can sometimes make the state estimation of the nonlinear system less accurate, so the assumption about the optimality of the “update” step can be extremely harmful, as it makes the algorithm become overconfident in an inaccurate estimation.
@@
 For those few existing theoretical analyses [13, 14, 15], the results are based on the assumption that the higher-ordered terms in the Taylor expansions of the measurement functions can be ignored.
-This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the ”update” step.
+This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the “update” step.
 In this study, we proposed a new framework that solves the problem of overconfidence.
@@
 The main contributions of this paper are:
-  • To the best of our knowledge, we are the first to give theoretical proof that the conventional framework for nonlinear KFs, which has been used for around sixty years, has the problem of underestimating the trace of the state covariance matrix.
+  • To the best of our knowledge, we are the first to give a mathematical proof that the conventional framework for nonlinear KFs, which has been used for around sixty years, generally has the problem of underestimating the trace of the state covariance matrix.
   • We proposed a new framework to solve the problem of inaccurate covariance estimation and validated the framework’s effectiveness by using both theoretical and empirical analysis.
@@
 $x_{k}=f(x_{k-1},u_{k-1})+w_{k-1}$ (1)
-where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow the zero-mean multivariate Gaussian distribution with covariance $Q_{k}$.
+where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow a zero-mean multivariate distribution with covariance $Q_{k}$.
 The estimations of the states made in the previous step are denoted as $\hat{x}_{k-1|k-1}$.
-Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow the zero-mean multivariate Gaussian distribution with covariance $P_{k-1|k-1}$.
+Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow a zero-mean multivariate distribution with covariance $P_{k-1|k-1}$.
 Whichever nonlinear KF is used, the first step of each iteration is called “predict”, which predicts the states using (1) and the estimations made in the previous time step.
-The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k|k-1}$ are assumed to follow the zero-mean multivariate Gaussian distribution with covariance $P_{k|k-1}$.
+The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k}$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{k|k-1}$.
 The second step of each iteration for any nonlinear KF is called “update”, which corrects the predicted states based on the measurements. Suppose that the measurements of the states are as follows:
 $z_{k}=h(x_{k})+v_{k}$ (2)
-where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow the zero-mean multivariate Gaussian distribution with covariance $R_{k}$.
+where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow a zero-mean multivariate distribution with covariance $R_{k}$.
 Given (2) and the predicted states, the measurements can also be predicted using a particular method specific to the type of nonlinear KFs.
-The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow the zero-mean multivariate Gaussian distribution with covariance $P_{y}$.
-The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy}$.
-The differences between the predicted and actual measurements are called the measurement residuals $\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$.
+The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{y,k}$.
+The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy,k}$.
+The differences between the predicted and actual measurements are called the measurement residuals, denoted as $\tilde{y}_{k}$.
+Namely:
+$\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$ (3)
 If we assume that the measurement noises are independent from the predicted states, then $\tilde{y}_{k}$ should have zero mean, and their covariance $S_{k}$ can be calculated by:
-$S_{k}=P_{y}+R_{k}$ (3)
+$S_{k}=P_{y,k}+R_{k}$ (4)
 KFs use the following equation to update the state estimation:
 $\hat{x}_{k|k}=\hat{x}_{k|k-1}+K\tilde{y}_{k}$ (5)
-where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (3,4), the covariance of $\hat{x}_{k|k}$ can be calculated as:
+where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (35), the covariance of $\hat{x}_{k|k}$ can be calculated as [16]:
 $P_{k|k}=P_{k|k-1}+KS_{k}K^{T}-P_{xy,k}K^{T}-KP_{xy,k}^{T}$ (6)
-KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of (5), we notice that the trace is minimized when:
-$K=K_{op}=P_{xy}S_{k}^{-1}$ (6)
+KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of tr$(P_{k|k})$ in (6), we find that the trace is minimized when:
+$K_{op}=P_{xy,k}S_{k}^{-1}$ (7)
 where $K_{op}$ is the theoretical optimal value of the Kalman gain. When (7) is satisfied, (6) becomes:
-$P_{k|k}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (7)
+$P_{k|k,op}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (8)
 Equations (58) illustrate the key idea of KFs, which is to update the states in a way that minimizes the trace of the state covariance matrix after the update.
 However, it should be noted that these equations cannot be directly applied to systems with nonlinear measurement functions.
-The primary reason is that the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ are all state-dependent for these nonlinear systems.
-Since the actual states are unknown, the actual values for these covariance matrices are also unknown.
-In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that make the conventional framework for nonlinear KFs overconfident and inaccurate.
+The primary reason is that $P_{y,k},S_{k}$, and $P_{xy,k}$ can only be calculated if the actual state and the distribution of the states are known.
+Since neither of these holds in practice, the actual values for these covariance matrices are also unknown.
+In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that can make the conventional framework for nonlinear KFs overconfident and inaccurate.
 2.2 The conventional framework and its problem
 While calculating $P_{y,k},S_{k}$, and $P_{xy,k}$ at the actual state is impossible, it is possible to approximate their values based on the predicted states.
-We denote these approximated covariance matrices as $P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$ since they are all estimated based on the prediction.
-With these approximations, the Kalman gain can be selected as:
-$K=P_{xy,k|k-1}S_{k|k-1}^{-1}$ (8)
-Although the detailed process of calculating $P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (5) with the estimated values.
-Namely, the conventional framework estimates the updated state covariance matrix by:
-$P_{k|k}$ $=P_{k|k-1}-KS_{k|k-1}K^{T}$ (9)
-$=P_{k|k-1}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$
-On the other hand, the actual value of the updated state covariance matrix $P_{k|k,ac}$ can be calculated by substituting (8) into (5). Namely:
-$P_{k|k,ac}=$ $P_{k|k-1}+P_{xy,k|k-1}S_{k|k-1}^{-1}S_{k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$ (10)
-$-P_{xy,k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k}^{T}$
-While guaranteeing that $P_{k|k,ac}=P_{k|k}$ holds for all combinations of $P_{xy,k|k-1}$ and $S_{k|k-1}$ is impossible, we would at least hope that $P_{k|k}$ is an unbiased estimation of $P_{k|k,ac}$ (i.e., $\operatorname{\mathbb{E}}(P_{k|k})=P_{k|k,ac}$).
-Unfortunately, we have the following theorem.
-Theorem 1
-Define $g(A,B)=\text{tr}(P_{k|k,ac})|_{P_{xy,k}=A,S_{k}=B}$.
-Suppose that $\operatorname{\mathbb{E}}(P_{xy,k|k-1})=P_{xy,k}$, $\operatorname{\mathbb{E}}(S_{k|k-1})=S_{k}$, and $P_{xy,k|k-1}$ and $S_{k|k-1}$ are independent from each other.
-When $S_{k|k-1}-S_{k}$ is small, $P_{k|k}$ calculated from (9) has the following characteristic:
-$\operatorname{\mathbb{E}}[\text{tr}(P_{k|k})]$ $=\operatorname{\mathbb{E}}[g(P_{xy,k|k-1},S_{k|k-1})]$ (11)
-$<g(P_{xy,k},S_{k})=\text{tr}(P_{k|k,ac})$
-Proof 2.2.
-Denote $P_{xy,k|k-1}=P_{xy,k}+\Delta P$, $S_{k|k}=S_{k}+\Delta S$. According to the definition, $\operatorname{\mathbb{E}}(\Delta P)=O,\operatorname{\mathbb{E}}(\Delta S)=O$. Therefore, we have:
-$g(P_{xy,k|k-1},S_{k|k-1})-g(P_{xy,k},S_{k|k-1})$ (12)
-$=$ $\text{tr}[P_{xy,k|k-1}(S_{k}+\Delta S)^{-1}\Delta S(S_{k}+\Delta S)^{-1}P_{xy,k|k-1}^{T}]$
-$-\text{tr}[\Delta PS_{k|k-1}^{-1}(P_{xy}+\Delta P)^{T}+(P_{xy}+\Delta P)S_{k|k-1}^{-1}\Delta P^{T}]$
-Note that $(S_{k}+\Delta S)^{-1}$ can be expanded as the Neumann series. When $\Delta S$ is small, we have the following approximation:
-$(S_{k}+\Delta S)^{-1}\Delta S(S_{k}+\Delta S)^{-1}$ (13)
-$\approx$ $(S_{k}^{-1}-S_{k}^{-1}\Delta SS_{k}^{-1})\Delta S(S_{k}^{-1}-S_{k}^{-1}\Delta SS_{k}^{-1})$
-$\approx$ $S_{k}^{-1}\Delta SS_{k}^{-1}-2S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}$
-Since $\Delta P$ and $\Delta S$ are independent from each other and have zero mean, we have:
-$\operatorname{\mathbb{E}}[g(P_{xy,k|k-1},S_{k|k-1})-g(P_{xy,k},S_{k})]$ (14)
-$\approx$ $\operatorname{\mathbb{E}}[\text{tr}(-2P_{xy,k|k-1}S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}P_{xy,k|k-1}^{T}$
-$-2\Delta PS_{k|k-1}^{-1}\Delta P^{T})]<0$
-Note that the derivation of (14) utilizes the fact that both $S_{k}^{-1}\Delta SS_{k}^{-1}\Delta SS_{k}^{-1}$ and $S_{k|k-1}^{-1}$ are positive semi-definite and symmetrical.
-Theorem 1 suggests that using (9) to estimate the state covariance matrix will lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values.
+Different nonlinear filters (e.g.
+UKF, EKF2, CKF) take different approaches to approximating these covariance matrices.
+In the subsequent analysis, we denote these approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ since they are all estimated based on the prediction.
+Although the detailed process of calculating ${\color[rgb]{1,0,0}P_{y,k|k-1}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (7) and (8) with the approximated values.
+Namely, the Kalman gain is selected as:
+${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ (9)
+and the state covariance matrix is updated by:
+${\color[rgb]{1,0,0}P_{k|k,est}}$ $=P_{k|k-1}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k-1}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (10)
+$=P_{k|k-1}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
+On the other hand, the actual value of the updated state covariance matrix ${\color[rgb]{1,0,0}P_{k|k,ac}}$ given the selected Kalman gain in (9) can be calculated by substituting (9) into (6).
+Namely:
+${\color[rgb]{1,0,0}P_{k|k,ac}}=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}S_{k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$ (11)
+$\quad-P_{xy,k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}P_{xy,k}^{T}$
+In other words, ${\color[rgb]{1,0,0}P_{k|k,ac}}$ plays the role of $P_{k|k}$ in (6), and ${\color[rgb]{1,0,0}P_{k|k,est}}$ plays the role of $P_{k|k,op}$ in (8).
+Intuitively, tr$({\color[rgb]{1,0,0}P_{k|k,ac}})$ tends to be larger than tr$({\color[rgb]{1,0,0}P_{k|k,est}})$ since tr$(P_{k|k,op})=\min$tr$(P_{k|k})\leq$ tr$(P_{k|k})$.
+However, this relationship of size is not always preserved after the approximations.
+For example, according to (10) and (11), when $P_{xy,k}=S_{k}=1$ and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}={\color[rgb]{1,0,0}S_{k|k-1}}=0.9$, we have ${\color[rgb]{1,0,0}P_{k|k,est}}-{\color[rgb]{1,0,0}P_{k|k,ac}}=0.1>0$.
+Nevertheless, in most cases in practice [17, 18], we do find that tr$({\color[rgb]{1,0,0}P_{k|k,ac}})>$tr$({\color[rgb]{1,0,0}P_{k|k,est}})$, meaning that using (10) will often (although not always) underestimate the actual covariance.
+In order to make this statement more rigorous in math, we consider the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$.
+Specifically, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k-1}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k-1}}$ is a random positive definite matrix.
+Note that we mark all the random covariance matrices in red in this section for clarity.
+Additionally, we assume that the “noises” that arise in the approximations are unbiased, so these random matrices should satisfy:
+$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k-1}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k-1}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}]=P_{xy,k}\end{cases}$ (12)
+These additional definitions and assumptions are made so that we can consider the “average” case and give a more mathematically rigorous definition of the overconfidence caused by (10).
+Namely, we have the following theorem.
+Theorem 1.
+Suppose that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ and ${\color[rgb]{1,0,0}S_{k|k-1}}$ are random matrices that satisfy (12).
+${\color[rgb]{1,0,0}S_{k|k-1}}$ is symmetric and positive definite.
+Then, $P_{k|k,est}$ calculated from (10) satisfies:
+$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}]\succeq 0$ (13)
+where $\succeq$” represents positive semidefiniteness, and the equality only holds when ${\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ is constant.
+{pf}
+Let ${\color[rgb]{1,0,0}P}={\color[rgb]{1,0,0}P_{xy,k|k-1}}S_{k}^{-\frac{1}{2}}$, ${\color[rgb]{1,0,0}S}=S_{k}^{-\frac{1}{2}}{\color[rgb]{1,0,0}S_{k|k-1}}S_{k}^{-\frac{1}{2}}$. According to (12),
+$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P}]=\bar{P}=P_{xy,k}S_{k}^{-\frac{1}{2}}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S}]=I\end{cases}$ (14)
+According to (10, 11),
+${\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}$ (15)
+$=$ ${\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-2}P^{T}}-\bar{P}{\color[rgb]{1,0,0}S^{-1}P^{T}}-{\color[rgb]{1,0,0}PS^{-1}}\bar{P}^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}$
+$=$ $({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}$
+For any fixed unit vector $\bm{v}$, ${\color[rgb]{1,0,0}P^{T}}\bm{v}$ is a random vector with a mean of $\bar{P}^{T}\bm{v}$. According to Lemma 4 in Appendix A,
+$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-1}}{\color[rgb]{1,0,0}P^{T}}\bm{v}]\geq\bm{v}^{T}\bar{P}\bar{P}^{T}\bm{v}.$ (16)
+Moving the right hand side to the left, the following must hold true:
+$\mathbb{E}[{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}]\succeq 0$ (17)
+Therefore, the expectation of (15) is always p.s.d.
+due to (17) and the fact that the first term in (15) is p.s.d.
+The equality only holds when ${\color[rgb]{1,0,0}PS^{-1}}$ is constant, which is equivalent to ${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ being constant (i.e., free from approximation errors).
+Theorem 1 suggests that using (10) to estimate the state covariance matrix tends to lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values.
+Admittedly, in practice, the errors of the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ are not as simple as zero-mean random matrices.
+However, Theorem 1 does give us a structured explanation as to why nonlinear KFs often underestimate the state covariance matrix.
+Note that an overconfident state covariance estimation can be very detrimental since it prevents the filter from correcting itself in subsequent iterations, thereby affecting the state estimation accuracy.
 Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xy,k}$ are related to the actual state, which is unknown.
-Note that this problem exists for all types of nonlinear KFs as long as (9) is used to update the state covariance estimation.
-We can further prove that the magnitude of such overconfidence is related to the accuracy of $S_{k|k-1}$ and $P_{xy,k|k-1}$.
-Namely, we have the following theorem.
-Theorem 2.3.
-Suppose that $P_{xy,k|k-1}=P_{xy}+t_{1}P,S_{k|k-1}=S_{k}+t_{2}S$, where $t_{1}\geq 0,t_{2}\geq 0$, $P$ is a random matrix with zero mean, and $S$ is a random symmetrical matrix with zero mean.
-Suppose that $P$ and $S$ are independent from each other, and $t_{2}S$ is very small.
-If we define the overconfidence in state covariance estimation as:
-$\Delta p_{k|k}(t_{1},t_{2})=\operatorname{\mathbb{E}}[g(P_{xy,k},S_{k|k})-g(P_{xy,k|k-1},S_{k|k-1})]$ (15)
-Then, for any $t_{a}\geq t_{b}$ and $t_{c}\geq t_{d}$, we have:
-$\Delta p_{k|k}(t_{a},t_{c})\geq\Delta p_{k|k}(t_{b},t_{d})$ (16)
-Proof 2.4.
-Let $\Delta P=t_{1}P,\Delta S=t_{2}S$. (14) can be rewritten as:
-$\Delta p_{k|k}(t_{1},t_{2})=\operatorname{\mathbb{E}}[g(P_{xy,k},S_{k|k})-g(P_{xy,k|k-1},S_{k|k-1})]$ (17)
-$=$ $\operatorname{\mathbb{E}}[\text{tr}(2t_{2}^{2}(P_{xy,k}+t_{1}P)S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}(P_{xy,k}+t_{1}P)^{T}$
-$+2t_{1}^{2}PS_{k|k-1}^{-1}P^{T})]$
-$=$ $2\operatorname{\mathbb{E}}[\text{tr}(t_{2}^{2}P_{xy,k}S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P_{xy,k}^{T}$
-$+t_{1}^{2}t_{2}^{2}PS_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P^{T}+t_{1}^{2}PS_{k|k-1}^{-1}P^{T})]$
-Since $P_{xy,k}S_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P_{xy,k}^{T}$, $PS_{k}^{-1}SS_{k}^{-1}SS_{k}^{-1}P^{T}$, and $PS_{k|k-1}^{-1}P^{T}$ are all positive semi-definite, their traces should all be greater or equal to zero.
-Therefore, (16) holds for any $t_{a}\geq t_{b}$ and $t_{c}\geq t_{d}$.
-Theorem 2.3 suggests that increasing the noise in $S_{k|k-1}$ and $P_{xy,k|k-1}$ proportionally can make the state covariance estimation less accurate.
-In other words, the problem of the conventional framework is likely to be more severe when $S_{k}$ and $P_{xy}$ change more dramatically with respect to the states.
+Therefore, this problem exists for all types of nonlinear KFs as long as (10) is used to update the state covariance estimation.
 In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal.
@@
 1”, in Fig.
-1), the EKF overestimates the position of the plane, and the absolute state estimation error $|x_{k|k}-x_{k}|$ becomes higher after the “update” step.
+1), the EKF overestimates the position of the plane, and the absolute value of the state estimation error $|\hat{x}_{k|k}-x_{k}|$ becomes higher after the “update” step.
 Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step.
@@
 This is visualized by the narrow yellow distribution in the “update” step in Fig.
-1), which does not cover the true state.
+1, which does not cover the true state.
 Figure 1: The basic idea of the new framework for nonlinear KFs.
@@
 This is visualized by the wider yellow distribution in the “recalibrate” step, which now covers the true state.
-Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $x_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$.
-We denote these re-approximated covariance matrices as $P_{y,k|k},S_{k|k}$, and $P_{xy,k|k}$ since they are all estimated based on the updated states.
-In the new framework, the state covariance matrix is updated as follows:
-$P_{k|k}=$ $P_{k|k-1}+KS_{k|k}K^{T}-P_{xy,k|k}K^{T}-KP_{xy,k|k}^{T}$ (18)
-$=$ $P_{k|k-1}+P_{xy,k|k-1}S_{k|k-1}^{-1}S_{k|k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}$
-$-P_{xy,k|k}S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}-P_{xy,k|k-1}S_{k|k-1}^{-1}P_{xy,k|k}^{T}$
-Since the new framework selects the same Kalman gain as the conventional framework, the actual value of the updated state covariance matrix $P_{k|k,ac}$ can still be calculated by (10).
-By comparing (18) and (10), we have the following theorem:
-Theorem 2.5.
-Assume that $\operatorname{\mathbb{E}}(P_{xy,k|k-1})=\operatorname{\mathbb{E}}(P_{xy,k|k})=P_{xy,k}$ and $\operatorname{\mathbb{E}}(S_{k|k-1})=\operatorname{\mathbb{E}}(S_{k|k})=S_{k}$.
-If $P_{xy,k|k-1}$, $P_{xy,k|k}$, $S_{k|k-1}$, and $S_{k|k}$ are all independent from each other, $P_{k|k}$ calculated in (18) is an unbiased estimation of $P_{k|k,ac}$.
+Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $\hat{x}_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$.
+We denote these re-approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ since they are all estimated based on the updated states.
+In the new framework, the state covariance matrix is updated by replacing $S_{k}$ with ${\color[rgb]{1,0,0}S_{k|k}}$ and $P_{xy,k}$ with ${\color[rgb]{1,0,0}P_{xy,k|k}}$ in (11):
+${\color[rgb]{1,0,0}P_{k|k,new}}=P_{k|k-1}+{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (18)
+$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
+$=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
+$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
+Note that the proposed new framework selects the same Kalman gain and does the same state update as the conventional one; the only difference is that the state covariance update equation is now a direct approximation of (11), which is the actual state covariance regardless of the selection of the Kalman gain $K$.
+Intuitively, this can prevent the KF from overconfidence because the equation no longer relies on the problematic assumption that the Kalman gain is optimal.
+To explain this advantage more rigorously, we consider the newly approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$.
+Similar to our previous assumptions about ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k}}$ is a random positive definite matrix.
+Additionally, we assume that the “noises” that arise in these approximations are unbiased, so these new random matrices should also satisfy:
+$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k}}]=P_{xy,k}\end{cases}$ (19)
+With these additional definitions and assumptions, we are able to give the following theorem about the advantage of the new framework:
+Theorem 2.
+Assume that $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k}})=P_{xy,k}$ and $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$.
+If ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, ${\color[rgb]{1,0,0}P_{xy,k|k}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}S_{k|k}}$ are all independent from each other, ${\color[rgb]{1,0,0}P_{k|k,new}}$ calculated in (18) is an unbiased estimation of ${\color[rgb]{1,0,0}P_{k|k,ac}}$.
 Namely, we have:
-$\operatorname{\mathbb{E}}(P_{k|k})=P_{k|k,ac}$ (19)
-Proof 2.6.
-With the assumptions made in Theorem 2.5, we have:
-$\operatorname{\mathbb{E}}(P_{k|k}-P_{k|k,ac})$ (20)
-$=$ $\operatorname{\mathbb{E}}[P_{xy,k|k-1}S_{k|k-1}^{-1}(S_{k|k}-S_{k})S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}]$
-$-\operatorname{\mathbb{E}}[(P_{xy,k|k}-P_{xy,k})S_{k|k-1}^{-1}P_{xy,k|k-1}^{T}]$
-$-\operatorname{\mathbb{E}}[P_{xy,k|k-1}S_{k|k-1}^{-1}(P_{xy,k|k}-P_{xy,k})^{T}]$
+$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})=0$ (20)
+{pf}
+With the assumptions made in Theorem 2, we have:
+$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})$ (21)
+$=$ $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}S_{k|k}}-S_{k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
+$-\operatorname{\mathbb{E}}[({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
+$-\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k})^{T}]$
 $=$ 00 0
-Theorem 2.5 suggests that the proposed new framework can prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other.
-This desirable feature comes from our replacement of (9) with (18), which highlights the fundamental idea of the new framework.
+The last equality equals zero because the terms in parentheses are zero, due to the unbiased (e.g. $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$) and independence assumptions.
+Theorem 2 suggests that the proposed new framework can generally prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other.
+Admittedly, this assumption about the independence of the approximation errors is very hard to verify and may not hold in practice.
+Nonetheless, the theorem does explain why recalibrating the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ after the state update is advantageous for accurate state covariance estimation.
+Specifically, the advantage comes from our replacement of (10) with (18), which highlights the fundamental idea of the new framework.
 That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately.
@@
 Algorithm 1 The proposed framework for nonlinear Kalman filters
-1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
 2:Initialization:
-3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
-4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
+3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
+4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
 5:for every time step $k$ do
 6:    Predict:
-7:     Estimate $\boldsymbol{\hat{x}}_{k|k-1}$ and $\boldsymbol{P}_{k|k-1}$
-8:    (based on $\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{P}_{k-1|k-1},\boldsymbol{f},\boldsymbol{u}_{k},$ and $\boldsymbol{Q}_{k}$)
+7:     Estimate $\bm{\hat{x}}_{k|k-1}$ and $\bm{P}_{k|k-1}$
+8:    (based on $\bm{\hat{x}}_{k-1|k-1},\bm{P}_{k-1|k-1},\bm{f},\bm{u}_{k},$ and $\bm{Q}_{k}$)
 9:    Update:
-10:     Estimate $\boldsymbol{\hat{y}}_{k|k-1},\boldsymbol{P}_{xy,k|k-1}$ and $\boldsymbol{P}_{y,k|k-1}$
-11:    (based on $\boldsymbol{\hat{x}}_{k|k-1},\boldsymbol{P}_{k|k-1},$ and $\boldsymbol{h}$)
-12:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
-13:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
-14:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
+10:     Estimate $\bm{\hat{y}}_{k|k-1},\bm{P}_{xy,k|k-1}$ and $\bm{P}_{y,k|k-1}$
+11:    (based on $\bm{\hat{x}}_{k|k-1},\bm{P}_{k|k-1},$ and $\bm{h}$)
+12:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
+13:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
+14:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
 15:    Recalibrate:
-16:     Estimate $\boldsymbol{P}_{xy,k|k}$ and $\boldsymbol{P}_{y,k|k}$
-17:    (based on $\boldsymbol{\hat{x}}_{k|k},\boldsymbol{P}_{k|k-1},$ and $\boldsymbol{h}$)
-18:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
-19:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
+16:     Estimate $\bm{P}_{xy,k|k}$ and $\bm{P}_{y,k|k}$
+17:    (based on $\bm{\hat{x}}_{k|k},\bm{P}_{k|k-1},$ and $\bm{h}$)
+18:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
+19:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
 20:    Back out:
-21:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
-22:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
-23:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
+21:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
+22:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
+23:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
 24:     end if
 25:end for
-26:where, $\boldsymbol{\hat{x}}$ is the estimated states, $\boldsymbol{P}$ is the states’ covariance matrix, $\boldsymbol{\hat{y}}$ is the estimated measurements, $\boldsymbol{P}_{xy}$ is the covariance between the states and the measurements, $\boldsymbol{P}_{y}$ is the covariance of the estimated measurements, $\boldsymbol{S}$ is the innovation (or residual) covariance, $\boldsymbol{K}$ is the Kalman gain.
+26:where, $\bm{\hat{x}}$ is the estimated states, $\bm{P}$ is the states’ covariance matrix, $\bm{\hat{y}}$ is the estimated measurements, $\bm{P}_{xy}$ is the covariance between the states and the measurements, $\bm{P}_{y}$ is the covariance of the estimated measurements, $\bm{S}$ is the innovation (or residual) covariance, $\bm{K}$ is the Kalman gain.
 The subscript $k|k-1$ represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
@@
 Or, to be more precise, the approximation of the measurement functions would become independent of the states.
-In this case, the estimations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
+In this case, the approximations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
 To give readers an intuitive understanding of the effect of the “recalibrate” and “back out” steps, in Fig.
@@
 The predicted state is $x_{1|0}=0$, its variance is $1.5^{2}$, and the standard deviation of the measurement is $\sigma_{y}=0.01$.
-This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is on the left of the predicted state.
-The state and covariance estimations made by different algorithms are summarized in different rows in the figure.
+This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is to the left of the predicted state.
+The state and covariance estimations made by different algorithms shown in different rows in the figure.
 Generally, the conventional framework always underestimates the actual state estimation error, which can be visualized by the narrow standard deviation estimations in Fig.
@@
 Figure 2: The comparison of the covariance estimation after the state update and recalibration using different frameworks and different types of nonlinear Kalman filters.
+2.4 Relationship to Iterated Kalman Filter
 The new framework described above may remind the readers of the iterated Kalman filter (IKF), which also re-approximates the measurement function after calculating the Kalman gain [19, 20, 21].
@@
 Namely, IKF always updates the states and covariance matrix simultaneously using (10), resulting in overconfident state estimations.
-When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence, according to Theorem 2.3.
+When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence.
 In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.
 Algorithm 2 The iterated extended Kalman filter with the old framework
-1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
 2:Initialization:
-3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
-4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
+3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
+4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
 5:for every time step $k$ do
 6:    Predict:
-7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
-8:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})$
-9:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}+\boldsymbol{Q}_{k}$
+7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
+8:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
+9:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
 10:    Update:
-11:     $\boldsymbol{x}_{0}=\boldsymbol{\hat{x}}_{k|k-1}$
+11:     $\bm{x}_{0}=\bm{\hat{x}}_{k|k-1}$
 12:     for $i=1$ to 1000 do
-13:         $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{x}_{i-1}}$
-14:         $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{x}_{i-1})+\boldsymbol{H}_{k|k-1}(\boldsymbol{x}_{i-1}-\boldsymbol{\hat{x}}_{k|k-1})$
-15:         $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
-16:         $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
-17:         $\boldsymbol{S}_{k|k-1,i}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
-18:         $\boldsymbol{K}_{i}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1,i}^{-1}$
-19:         $\boldsymbol{x}_{i}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}_{i}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
-20:         if $i>1$ and $||\boldsymbol{x}_{i}-\boldsymbol{x}_{i-1}||>||\boldsymbol{x}_{i-1}-\boldsymbol{x}_{i-2}||$ then
-21:              $\boldsymbol{S}_{k|k-1,i}=\boldsymbol{S}_{k|k-1,i-1}$
-22:              $\boldsymbol{K}_{i}=\boldsymbol{K}_{i-1}$
-23:              $\boldsymbol{x}_{i}=\boldsymbol{x}_{i-1}$
+13:         $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{x}_{i-1}}$
+14:         $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{x}_{i-1})+\bm{H}_{k|k-1}(\bm{x}_{i-1}-\bm{\hat{x}}_{k|k-1})$
+15:         $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+16:         $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+17:         $\bm{S}_{k|k-1,i}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
+18:         $\bm{K}_{i}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1,i}^{-1}$
+19:         $\bm{x}_{i}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{i}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+20:         if $i>1$ and $||\bm{x}_{i}-\bm{x}_{i-1}||>||\bm{x}_{i-1}-\bm{x}_{i-2}||$ then
+21:             $\bm{S}_{k|k-1,i}=\bm{S}_{k|k-1,i-1}$
+22:             $\bm{K}_{i}=\bm{K}_{i-1}$
+23:             $\bm{x}_{i}=\bm{x}_{i-1}$
 24:             break
 25:         end if
-26:         if $\max_{j}|1-\boldsymbol{x}_{i}(j)/\boldsymbol{x}_{i-1}(j)|<0.001$ then
+26:         if $\max_{j}|1-\bm{x}_{i}(j)/\bm{x}_{i-1}(j)|<0.001$ then
 27:              break
@@
 29:     end for
-30:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{S}_{k|k-1,i}$
-31:     $\boldsymbol{K}=\boldsymbol{K}_{i}$
-32:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{x}_{i}$
-33:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}-\boldsymbol{K}\boldsymbol{S}_{k|k-1}\boldsymbol{K}^{T}$
+30:     $\bm{S}_{k|k-1}=\bm{S}_{k|k-1,i}$
+31:     $\bm{K}=\bm{K}_{i}$
+32:     $\bm{\hat{x}}_{k|k}=\bm{x}_{i}$
+33:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}-\bm{K}\bm{S}_{k|k-1}\bm{K}^{T}$
 34:end for
@@
 A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy?
-Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states ($P_{y,k|k-1},S_{k|k-1}$, and $P_{xy,k|k-1}$) are independent of the approximations made at the updated states ($P_{y,k|k},S_{k|k}$, and $P_{xy,k|k}$).
-However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2.5.
+Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states (${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$) are independent of the approximations made at the updated states (${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$).
+However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2.
 For example, the IEKF algorithm shown in Algorithm 2 approximates the measurement function iteratively at $x_{i-1}$ (rows 9–13).
@@
 To demonstrate the effectiveness of the proposed framework, we validated it in five different applications suitable for nonlinear KFs.
-A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix 5.
+A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix B.
 For each scenario, we combined the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF.
 We also implemented the IEKF for comparison.
-The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 36 in Appendix 6.
+The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 35 in Appendix C.
 All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups).
@@
 As we mentioned in the Introduction, the fundamental idea of the proposed framework is to make a more accurate covariance estimation by re-approximating the measurement functions after the state update.
-To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the system) in the old and new framework in Fig.
+To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the first system) in the old and new framework in Fig.
 8.
-In the figure, for each type of nonlinear KFs (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation.
+In the figure, for each type of nonlinear KF (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation.
 Apparently, when the standard deviations of the two measurements are lower than 0.1 meters, the conventional framework always significantly underestimates the covariance, sometimes by several orders of magnitude, regardless of the type of nonlinear KFs used.
@@
 8 is that the estimated RMSE of the x-axis position (the dashed lines) is very similar for different combinations of frameworks and KFs.
-This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction – to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
+This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction — to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
 Another important metric for evaluating the performance of different filters is the convergence characteristics.
@@
 Namely, when the system has linear state transition functions and nonlinear measurement functions (which is true for the first two applications in the table), the “update” step should be accountable for most of the computational time in the conventional framework and re-approximating the measurement functions can almost double the runtime in the worst case.
-The normalized runtime of the new framework is related to the type of nonlinear KFs as well.
-In UKF and CKF, one of the most time-consuming steps is to calculate the square root of the state covariance matrix.
-This step is respectively done one, two, two, and three times in each iteration of UKF (old framework), UKF (new framework), CKF (old framework), and CKF (new framework).
-Therefore, the computational time increase (in percentage) brought by the new framework is longer for UKF than for CKF.
-Noticing that the new framework can reduce the estimation error by more than an order of magnitude when the measurement noise is small, we can say that the extra computational complexity is generally worth it.
+However, noticing that the new framework can reduce the estimation error by more than an order of magnitude when the measurement noise is small, we can say that the extra computational complexity is generally worth it.
 Another thing to note is that the combination of EKF and the new framework requires significantly less computational time than the combination of any other nonlinear KFs and the old framework.
@@
 Therefore, most applications of nonlinear KFs in real life use the discrete-time form, and they can all benefit from the new framework proposed in this paper.
-{appendices}
-5 The detailed simulation setup
+{ack}
+This material is based upon work supported, in part, by the National Science Foundation under Grant No. 1847177.
+Appendix A Supplementary proof for Theorem 1
+Lemma 3.
+Let $(\Omega,\mathcal{F},\mathbb{P})$ be an underlying probability space. Assume the random column vectors
+$x:\Omega\to\mathbb{R}^{n},\;y:\Omega\to\mathbb{R}^{n}$
+satisfy $\mathbb{E}[||x||^{2}]<\infty$ and $\mathbb{E}[||y||^{2}]<\infty$ (i.e., $x,y\in L^{2}(\Omega;\mathbb{R}^{n})$). Then, the operation on $L^{2}(\Omega;\mathbb{R}^{n})$ by
+$f(u,v)=\mathbb{E}[u^{T}v]$
+is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$.
+{pf}
+We start by proving that $f$ is well-defined on $L^{2}$.
+For any $u,v\in L^{2}$, both $\mathbb{E}[u^{T}u]$ and $\mathbb{E}[v^{T}v]$ are finite.
+Therefore, $\mathbb{E}[(u+v)^{T}(u+v)]$ should also be finite.
+Notice that
+$\mathbb{E}[(u+v)^{T}(u+v)]=\mathbb{E}[u^{T}u]+\mathbb{E}[v^{T}v]+2\mathbb{E}[u^{T}v],$
+we can conclude that $\mathbb{E}[u^{T}v]$ should also be finite.
+Therefore, $f$ is well-defined on $L^{2}$.
+Next, to prove that $f$ is an inner product, we need to show that it satisfies symmetry, linearity in the first argument, and positive-definiteness.
+Symmetry is satisfied because $\mathbb{E}[u^{T}v]=\mathbb{E}[v^{T}u]$.
+Linearity is satisfied because for any scalars $a,b$ and $x,y,z\in L^{2}$,
+$\mathbb{E}[(ax+by)^{T}z]=a\mathbb{E}[x^{T}z]+b\mathbb{E}[y^{T}z].$
+Positive-definiteness is satisfied because for any $x\in L^{2}$, $\mathbb{E}[x^{T}x]\geq 0$, and the equality sign only holds when $x\equiv 0$.
+Lemma 4.
+Let ${\color[rgb]{1,0,0}S}$ be a random symmetric and positive definite matrix whose mean is the identity matrix $I$.
+Let $\bm{v}$ be a random vector whose mean is $\bm{v}_{0}$.
+The following inequality holds:
+$\mathbb{E}(\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v})\geq\bm{v}_{0}^{T}\bm{v}_{0}$ (22)
+Note that the equality sign only holds when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
+{pf}
+Let $\bm{x}={\color[rgb]{1,0,0}S^{-\frac{1}{2}}}\bm{v},\bm{y}={\color[rgb]{1,0,0}S^{\frac{1}{2}}}\bm{v}_{0}$. From the definitions,
+$\begin{cases}\mathbb{E}[\bm{x}^{T}\bm{x}]=\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\\ \mathbb{E}[\bm{y}^{T}\bm{y}]=\mathbb{E}[\bm{v}_{0}^{T}{\color[rgb]{1,0,0}S}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\\ \mathbb{E}[\bm{x}^{T}\bm{y}]=\mathbb{E}[\bm{v}^{T}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\end{cases}$ (23)
+According to Lemma 3, $\mathbb{E}[\bm{x}^{T}\bm{y}]$ is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$. Applying Cauchy–Schwarz inequality, we have:
+$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\geq\frac{(\bm{v}_{0}^{T}\bm{v}_{0})^{2}}{\bm{v}_{0}^{T}\bm{v}_{0}}$ (24)
+which completes the proof. Note that the equality holds only when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
+Appendix B The detailed simulation setup
 This paper uses five different applications of nonlinear KFs to test the effectiveness of the new framework: 3D target tracking, terrain-referenced navigation, synchronous generator state estimation, pendulum state estimation, and battery state estimation.
 Figure 11: Examples of nonlinear systems used in this paper for validation. (a) 3D object tracking. (b) Terrain-referenced navigation. (c) Pendulum state estimation. (d) Battery state estimation.
-5.1 3D target tracking
+B.1 3D target tracking
 This example is adapted from Example 3.3.1 in [26].
@@
 The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example.
-When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (22,23).
+When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (26,27).
 Note that the noises are omitted in the equations.
@@
 The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states.
-The process noise covariance matrix is also diagonal, and its diagonal values are 0, 0, 0, $10^{–6}$, $10^{–6}$, and $10^{–6}$.
-5.2 Terrain-referenced navigation
+The process noise covariance matrix is also diagonal, and its diagonal values are 0, 0, 0, $10^{-6}$, $10^{-6}$, and $10^{-6}$.
+B.2 Terrain-referenced navigation
 This example is adapted from Example 3.3.2 in [26].
@@
 The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively.
-When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (24, 25).
+When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (28, 29).
 Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.
@@
 The process noise covariance matrix is also diagonal, and its diagonal values are both 0.25 m2.
-5.3 Synchronous generator state estimation
+B.3 Synchronous generator state estimation
 This example is adapted from a research paper [27].
@@
 All the model parameters are selected to be the same as in [27].
-The system’s state transition and measurement functions can be written as (27, 28).
+The system’s state transition and measurement functions can be written as (31, 32).
 Note that the noises are omitted from the equations.
@@
 The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$.
-The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{–4}$, $10^{–10}$, $10^{–4}$, and $10^{–4}$.
+The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{-4}$, $10^{-10}$, $10^{-4}$, and $10^{-4}$.
 The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states.
-The process noise covariance matrix is also diagonal, and its diagonal values are $10^{–10}$, $10^{–16}$, $10^{–10}$, and $10^{–10}$.
-5.4 Pendulum state estimation
+The process noise covariance matrix is also diagonal, and its diagonal values are $10^{-10}$, $10^{-16}$, $10^{-10}$, and $10^{-10}$.
+B.4 Pendulum state estimation
 As shown in Fig.
@@
 The time interval is set to $\Delta t=0.01$s.
-According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (29,30).
+According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (33,34).
 Note that the noises are omitted in the equations.
@@
 The process noise covariance matrix is also diagonal; its diagonal values are $10^{-10}$ and 0.
-5.5 Battery state estimation
+B.5 Battery state estimation
 This example is adapted from a research paper [28].
@@
 The measurement $y_{k}$ is the battery’s terminal voltage.
-The input is the charging current, which, in this example, is a 3-level square wave shown below:
+The input is the charging current, which, in this example, is a three-level square wave shown below:
 $I_{k}=\begin{cases}-2,75\geq k>15\\ 2,165\geq k>105\\ 0,\text{else}\end{cases}$ (35)
@@
 $OCV=\sum_{i=0}^{9}\big{(}\frac{SOH-0.8}{0.2}a_{100,i+1}+\frac{1-SOH}{0.2}a_{80,i+1}\big{)}$ (36)
-where $a_{100}=$ [1390.38, -6961.31, 14760.31, -17230.92, 12055.71, -5162.75, 1330.60, -196.37, 15.60, 2.96], $a_{80}=$ [813.94, -4229.96, 9345.49, -11415.38, 8396.15, -3801.07, 1043.09, -165.29, 14.28, 2.96].
+where $a_{100}=[1390.38,-6961.31,14760.31,-17230.92,$ 12055.71,5162.75,1330.60,196.37,15.60,2.96]12055.71,-5162.75,1330.60,-196.37,15.60,2.96]12055.71 , - 5162.75 , 1330.60 , - 196.37 , 15.60 , 2.96 ], $a_{80}=$ [813.94,4229.96,9345.49,11415.38,8396.15,3801.07,[813.94,-4229.96,9345.49,-11415.38,8396.15,-3801.07,[ 813.94 , - 4229.96 , 9345.49 , - 11415.38 , 8396.15 , - 3801.07 , 1043.09,165.29,14.28,2.96]1043.09,-165.29,14.28,2.96]1043.09 , - 165.29 , 14.28 , 2.96 ].
 The time interval is set to $\Delta t=1$ s, and the total simulation time is 180 seconds.
-The system’s state transition and measurement functions can be written as (33,34).
+The system’s state transition and measurement functions can be written as (37,38).
 Note that the noises are omitted in the equations.
@@
 Since the input current also appears in the measurement function, the actual measurement noise is the combination of the measurement noise of the voltmeter and the noise caused by the term $R_{1}I_{k}$ in the measurement function.
-6 The detailed algorithm of various nonlinear Kalman filters used in the paper
+Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
 The detailed algorithms of EKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 35.
-Note that all these algorithms degrade to the conventional framework when the ”recalibrate” and ”update” steps are replaced by (9).
+Note that all these algorithms degrade to the conventional framework when the “Recalibrate” and “Back out” steps are replaced by (10).
 Algorithm 3 The extended Kalman filter with the new framework
-1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
 2:Initialization:
-3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
-4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
+3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
+4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
 5:for every time step $k$ do
 6:    Predict:
-7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
-8:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})$
-9:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}+\boldsymbol{Q}_{k}$
+7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
+8:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
+9:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
 10:    Update:
-11:     $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k-1}}$
-12:     $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{\hat{x}}_{k|k-1})$
-13:     $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
-14:     $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
-15:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
-16:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
-17:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
+11:     $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k-1}}$
+12:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})$
+13:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+14:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+15:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
+16:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
+17:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
 18:    Recalibrate:
-19:     $\boldsymbol{H}_{k|k}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k}}$
-20:     $\boldsymbol{P}_{xy,k|k}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
-21:     $\boldsymbol{P}_{y,k|k}=\boldsymbol{H}_{k|k}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
-22:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
-23:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
+19:     $\bm{H}_{k|k}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k}}$
+20:     $\bm{P}_{xy,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+21:     $\bm{P}_{y,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+22:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
+23:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
 24:    Back out:
-25:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
-26:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
-27:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
+25:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
+26:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
+27:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
 28:     end if
@@
 Algorithm 4 The second-order extended Kalman filter with the new framework
-1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, measurements number $n_{m}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, states number $n_{x}$, measurements number $n_{m}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
 2:Initialization:
-3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
-4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
+3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
+4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
 5:for every time step $k$ do
 6:    Predict:
-7:     $\boldsymbol{F}_{k}=\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}$
-8:     $\boldsymbol{F}_{i,k}^{*}=\big{[}\frac{\partial^{2}f_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1}}\big{]}_{mn},i=1,\ldots,n_{x}$
-9:     $\boldsymbol{\hat{x}}_{k|k-1}=\boldsymbol{f}(\boldsymbol{\hat{x}}_{k-1|k-1},\boldsymbol{u}_{k-1})+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{F}_{i,k}^{*}\boldsymbol{P}_{k-1|k-1})\big{]}_{i}$
-10:     $\boldsymbol{P}_{k|k-1}=\boldsymbol{F}_{k}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{k}^{T}$
-11:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{F}_{i,k}^{*}\boldsymbol{P}_{k-1|k-1}\boldsymbol{F}_{j,k}^{*}\boldsymbol{P}_{k-1|k-1})\big{]}_{ij}+\boldsymbol{Q}_{k}$
+7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
+8:     $\bm{F}_{i,k}^{*}=\big{[}\frac{\partial^{2}f_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}\big{]}_{mn},i=1,\ldots,n_{x}$
+9:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})+\frac{1}{2}\big{[}\text{tr}(\bm{F}_{i,k}^{*}\bm{P}_{k-1|k-1})\big{]}_{i}$
+10:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}$
+11:      $+\frac{1}{2}\big{[}\text{tr}(\bm{F}_{i,k}^{*}\bm{P}_{k-1|k-1}\bm{F}_{j,k}^{*}\bm{P}_{k-1|k-1})\big{]}_{ij}+\bm{Q}_{k}$
 12:    Update:
-13:     $\boldsymbol{H}_{k|k-1}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k-1}}$
-14:     $\boldsymbol{H}_{i,k|k-1}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k|k-1}}\big{]}_{mn},i=1,\ldots,n_{m}$
-15:     $\boldsymbol{\hat{y}}_{k|k-1}=\boldsymbol{h}(\boldsymbol{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k-1}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{i}$
-16:     $\boldsymbol{P}_{xy,k|k-1}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
-17:     $\boldsymbol{P}_{y,k|k-1}=\boldsymbol{H}_{k|k-1}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k-1}^{T}$
-18:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k-1}^{*}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{j,k|k-1}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{ij}$
-19:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
-20:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
-21:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
+13:     $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k-1}}$
+14:     $\bm{H}_{i,k|k-1}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k|k-1}}\big{]}_{mn},i=1,\ldots,n_{m}$
+15:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{i}$
+16:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+17:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+18:      $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
+19:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
+20:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
+21:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
 22:    Recalibrate:
-23:     $\boldsymbol{H}_{k|k}=\frac{\partial\boldsymbol{h}}{\partial\boldsymbol{x}}|_{\boldsymbol{\hat{x}}_{k|k}}$
-24:     $\boldsymbol{H}_{i,k|k}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\boldsymbol{\hat{x}}_{k|k}}\big{]}_{mn},i=1,\ldots,n_{m}$
-25:     $\boldsymbol{P}_{xy,k|k}=\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
-26:     $\boldsymbol{P}_{y,k|k}=\boldsymbol{H}_{k|k}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{k|k}^{T}$
-27:      $+\frac{1}{2}\big{[}\text{tr}(\boldsymbol{H}_{i,k|k}^{*}\boldsymbol{P}_{k|k-1}\boldsymbol{H}_{j,k|k}^{*}\boldsymbol{P}_{k|k-1})\big{]}_{ij}$
-28:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
-29:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
+23:     $\bm{H}_{k|k}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k}}$
+24:     $\bm{H}_{i,k|k}^{*}=\big{[}\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k|k}}\big{]}_{mn},i=1,\ldots,n_{m}$
+25:     $\bm{P}_{xy,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+26:     $\bm{P}_{y,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+27:       $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
+28:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
+29:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
 30:    Back out:
-31:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
-32:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
-33:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
+31:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
+32:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
+33:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
 34:     end if
 35:end for
+Algorithm 5 The cubature Kalman filter with the new framework
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, states number $n_{x}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
+2:Initialization:
+3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
+4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
+5:for every time step $k$ do
+6:    Predict:
+7:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$,
+8:        where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
+9:     $\bm{\mathcal{X}}_{i,k-1}=\bm{\hat{x}}_{k-1|k-1}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+10:     $\bm{\mathcal{X}}_{i+n_{x},k-1}=\bm{\hat{x}}_{k-1|k-1}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+11:     $\bm{\mathcal{X}}_{i,k|k-1}^{*}=\bm{f}(\bm{\mathcal{X}}_{i,k-1},\bm{u}_{k-1}),i=1,\ldots,2n_{x}$
+12:     $\bm{\hat{x}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{X}}_{i,k|k-1}^{*}$
+13:     $\bm{P}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}^{*}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{X}}_{i,k|k-1}^{*}$
+14:        $-\bm{\hat{x}}_{k|k-1}]^{T}+\bm{Q}_{k}$
+15:    Update:
+16:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k|k-1})$
+17:     $\bm{\mathcal{X}}_{i,k|k-1}=\bm{\hat{x}}_{k|k-1}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+18:     $\bm{\mathcal{X}}_{i+n_{x},k|k-1}=\bm{\hat{x}}_{k|k-1}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+19:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
+20:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k-1}$
+21:     $\bm{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
+22:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
+23:     $\bm{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
+24:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
+25:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
+26:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
+27:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+28:    Recalibrate:
+29:     $\bm{\mathcal{X}}_{i,k|k}=\bm{\hat{x}}_{k|k}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+30:     $\bm{\mathcal{X}}_{i+n_{x},k|k}=\bm{\hat{x}}_{k|k}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+31:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
+32:     $\bm{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k}$
+33:     $\bm{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
+34:     $\bm{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
+35:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
+36:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+37:    Back out:
+38:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
+39:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
+40:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
+41:     end if
+42:end for
 Algorithm 6 The unscented Kalman filter with the new framework
-1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$, hyperparameters $\alpha=10^{-3},\beta=2,\kappa=0$
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, states number $n_{x}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$, hyperparameters $\alpha=10^{-3},\beta=2,\kappa=0$
 2:Initialization:
-3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
-4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
+3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
+4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
 5:$\lambda=\alpha^{2}(n_{x}+\kappa)-n_{x}$
@@
 10:    Predict:
-11:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k-1|k-1}}}]_{i},i=1,...,n_{x}$
-12:     $\boldsymbol{\mathcal{X}}_{0,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}$
-13:     $\boldsymbol{\mathcal{X}}_{i,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}+\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-14:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}-\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-15:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}=\boldsymbol{f}(\boldsymbol{\mathcal{X}}_{i,k-1},\boldsymbol{u}_{k-1}),i=0,1,\ldots,2n_{x}$
-16:     $\boldsymbol{\hat{x}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{X}}_{i,k|k-1}$
-17:     $\boldsymbol{P}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{X}}_{i,k|k-1}$
-18:      $-\boldsymbol{\hat{x}}_{k|k-1}]^{T}+\boldsymbol{Q}_{k}$
+11:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$, where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
+12:     $[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=i^{th}\text{ column of }\bm{L},i=1,...,n_{x}$
+13:     $\bm{\mathcal{X}}_{0,k-1}=\bm{\hat{x}}_{k-1|k-1}$
+14:     $\bm{\mathcal{X}}_{i,k-1}=\bm{\hat{x}}_{k-1|k-1}+\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+15:     $\bm{\mathcal{X}}_{i+n_{x},k-1}=\bm{\hat{x}}_{k-1|k-1}-\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+16:     $\bm{\mathcal{X}}_{i,k|k-1}=\bm{f}(\bm{\mathcal{X}}_{i,k-1},\bm{u}_{k-1}),i=0,1,\ldots,2n_{x}$
+17:     $\bm{\hat{x}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{X}}_{i,k|k-1}$
+18:     $\bm{P}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}]^{T}+\bm{Q}_{k}$
 19:    Update:
-20:     $\boldsymbol{\mathcal{Y}}_{i,k|k-1}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
-21:     $\boldsymbol{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
-22:     $\boldsymbol{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
-23:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
-24:     $\boldsymbol{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{Y}}_{i,k|k-1}-\boldsymbol{\hat{y}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
-25:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
-26:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
-27:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
-28:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
-29:    Recalibrate:
-30:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k|k-1}}}]_{i},i=1,...,n_{x}$
-31:     $\boldsymbol{\mathcal{X}}_{0,k|k}=\boldsymbol{\hat{x}}_{k|k}$
-32:     $\boldsymbol{\mathcal{X}}_{i,k|k}=\boldsymbol{\hat{x}}_{k|k}+\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-33:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k}=\boldsymbol{\hat{x}}_{k|k}-\sqrt{\lambda+n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-34:     $\boldsymbol{\mathcal{Y}}_{i,k|k}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
-35:     $\boldsymbol{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\boldsymbol{\mathcal{Y}}_{i,k|k}$
-36:     $\boldsymbol{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{X}}_{i,k|k}-\boldsymbol{\hat{x}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
-37:     $\boldsymbol{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
-38:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
-39:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
-40:    Back out:
-41:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
-42:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
-43:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
-44:     end if
-45:end for
-Algorithm 6 The cubature Kalman filter with the new framework
-1:Input: Process noise covariance matrix $\boldsymbol{Q}_{k}$, Measurement noise covariance matrix $\boldsymbol{R}_{k}$, states number $n_{x}$, state transition function $\boldsymbol{f}(\boldsymbol{x},\boldsymbol{u})$, measurement function $\boldsymbol{h}{(\boldsymbol{x})}$, system inputs $\boldsymbol{u}_{k}$, measurements $\boldsymbol{z}_{k}$
-2:Initialization:
-3:$\boldsymbol{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\boldsymbol{x}_{0}]$
-4:$\boldsymbol{P}_{0|0}=\operatorname{\mathbb{E}}[(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})(\boldsymbol{\hat{x}}_{0|0}-\boldsymbol{x}_{0})^{T}]$
-5:for every time step $k$ do
-6:    Predict:
-7:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k-1|k-1}}}]_{i},i=1,...,n_{x}$
-8:     $\boldsymbol{\mathcal{X}}_{i,k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-9:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k-1}=\boldsymbol{\hat{x}}_{k-1|k-1}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-10:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}=\boldsymbol{f}(\boldsymbol{\mathcal{X}}_{i,k-1},\boldsymbol{u}_{k-1}),i=1,\ldots,2n_{x}$
-11:     $\boldsymbol{\hat{x}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}$
-12:     $\boldsymbol{P}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{X}}_{i,k|k-1}^{*}$
-13:      $-\boldsymbol{\hat{x}}_{k|k-1}]^{T}+\boldsymbol{Q}_{k}$
-14:    Update:
-15:     $\boldsymbol{L}_{i}=[\sqrt{\boldsymbol{\boldsymbol{P}_{k|k-1}}}]_{i},i=1,...,n_{x}$
-16:     $\boldsymbol{\mathcal{X}}_{i,k|k-1}=\boldsymbol{\hat{x}}_{k|k-1}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-17:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k-1}=\boldsymbol{\hat{x}}_{k|k-1}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-18:     $\boldsymbol{\mathcal{Y}}_{i,k|k-1}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
-19:     $\boldsymbol{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
-20:     $\boldsymbol{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k-1}-\boldsymbol{\hat{x}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
-21:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
-22:     $\boldsymbol{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{Y}}_{i,k|k-1}-\boldsymbol{\hat{y}}_{k|k-1}][\boldsymbol{\mathcal{Y}}_{i,k|k-1}$
-23:      $-\boldsymbol{\hat{y}}_{k|k-1}]^{T}$
-24:     $\boldsymbol{S}_{k|k-1}=\boldsymbol{P}_{y,k|k-1}+\boldsymbol{R}_{k}$
-25:     $\boldsymbol{K}=\boldsymbol{P}_{xy,k|k-1}\boldsymbol{S}_{k|k-1}^{-1}$
-26:     $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}+\boldsymbol{K}(\boldsymbol{\boldsymbol{z}_{k}-\hat{y}}_{k|k-1})$
-27:    Recalibrate:
-28:     $\boldsymbol{\mathcal{X}}_{i,k|k}=\boldsymbol{\hat{x}}_{k|k}+\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-29:     $\boldsymbol{\mathcal{X}}_{i+n_{x},k|k}=\boldsymbol{\hat{x}}_{k|k}-\sqrt{n_{x}}\boldsymbol{L}_{i},i=1,\ldots,n_{x}$
-30:     $\boldsymbol{\mathcal{Y}}_{i,k|k}=\boldsymbol{h}(\boldsymbol{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
-31:     $\boldsymbol{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\boldsymbol{\mathcal{Y}}_{i,k|k}$
-32:     $\boldsymbol{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{X}}_{i,k|k}-\boldsymbol{\hat{x}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
-33:     $\boldsymbol{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}][\boldsymbol{\mathcal{Y}}_{i,k|k}-\boldsymbol{\hat{y}}_{k|k}]^{T}$
-34:     $\boldsymbol{S}_{k|k}=\boldsymbol{P}_{y,k|k}+\boldsymbol{R}_{k}$
-35:     $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}+\boldsymbol{K}\boldsymbol{S}_{k|k}\boldsymbol{K}^{T}-\boldsymbol{P}_{xy,k|k}\boldsymbol{K}^{T}-\boldsymbol{K}\boldsymbol{P}_{xy,k|k}^{T}$
-36:    Back out:
-37:     if $\text{tr}(\boldsymbol{P}_{k|k})>\text{tr}(\boldsymbol{P}_{k|k-1})$ then
-38:         $\boldsymbol{\hat{x}}_{k|k}=\boldsymbol{\hat{x}}_{k|k-1}$
-39:         $\boldsymbol{P}_{k|k}=\boldsymbol{P}_{k|k-1}$
-40:     end if
-41:end for
-Acknowledgment
-This material is based upon work supported, in part, by the National Science Foundation under Grant No. 1847177.
+20:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k})$
+21:     $\bm{\mathcal{X}}_{0,k|k-1}=\bm{\hat{x}}_{k|k-1}$
+22:     $\bm{\mathcal{X}}_{i,k|k-1}=\bm{\hat{x}}_{k|k-1}+\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+23:     $\bm{\mathcal{X}}_{i+n_{x},k|k-1}=\bm{\hat{x}}_{k|k-1}-\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
+24:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
+25:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k-1}$
+26:     $\bm{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
+27:     $\bm{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
+28:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
+29:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
+30:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+31:    Recalibrate:
+32:     $\bm{\mathcal{X}}_{i,k|k}=\bm{\mathcal{X}}_{i,k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1}),i=0,\ldots,2n_{x}$
+33:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
+34:     $\bm{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k}$
+35:     $\bm{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
+36:     $\bm{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
+37:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
+38:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+39:    Back out:
+40:     if $\text{tr}(\bm{P}_{k|k})>\text{tr}(\bm{P}_{k|k-1})$ then
+41:         $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}$
+42:         $\bm{P}_{k|k}=\bm{P}_{k|k-1}$
+43:     end if
+44:end for

### v11

@@
 It is assumed that both the state transition and measurement functions contain some additive noise.
-The KF optimally combines the system knowledge and measurements to minimize the expected value of the mean-squared error of the state estimation.
+Among all linear estimators, the KF optimally combines the system knowledge and measurements to minimize the expected value of the mean-squared error of the state estimation.
 Additionally, when the process noise, measurement noise, and state estimation errors all follow a multivariate normal (Gaussian) distribution, the state estimation given by the Kalman filter is also the maximum likelihood estimation [1].
@@
 Additionally, EKF2 still sometimes gives inaccurate and overconfident estimations, so it is generally used less frequently than EKF.
-Both EKF and EKF2 require the calculation of the Jacobian matrices of the state transition functions and measurement functions, which may not always be accessible in real-life systems.
-In 1997, Simon et al.
-introduced another extension of KF to nonlinear systems that requires no direct calculation of the Jacobian matrix and avoids the accuracy loss caused by linearization [8].
+Classical EKF and EKF2 formulations are usually presented in a way that requires explicit Jacobian matrices of the state-transition and measurement functions, which may not always be accessible or convenient to compute in real-world systems.
+Although derivative-free variants of the EKF and EKF2 have been proposed [8, 9], a particularly influential derivative-free extension of the KF to nonlinear systems is the UKF.
+In 1997, Julier and Uhlmann introduced a derivative-free extension of the KF to nonlinear systems that avoids explicit Jacobian computations and mitigates the accuracy loss caused by first-order linearization [10].
 In later literature, this extension is usually referred to as the UKF [11].
-UKF approximates the distribution of the states after nonlinear transformations by sampling several points (called sigma points) around the estimated mean of the states.
-These points go through the same transformation as the mean, and a more accurate estimation of the states’ distribution after the transformation can be obtained with the help of these extra sample points.
-In UKF, the sigma points are selected deterministically based on each state’s standard deviation.
-This technique guarantees that the sigma points are evenly distributed around the central point and can provide a more accurate state estimation.
+The UKF approximates the mean and covariance of the states after nonlinear transformations by propagating a set of deterministically chosen sigma points around the estimated state mean.
+These sigma points are constructed from the square root of the state covariance matrix so that their weighted mean and covariance match the prior mean and covariance.
+After applying the nonlinear transformation to each sigma point, the transformed mean and covariance are recovered as weighted averages of the transformed points, which typically yields a more accurate approximation of the first two moments than a first-order linearization.
 Similar ideas have also been used in several other nonlinear KFs, including cubature Kalman filter (CKF) [12], quadrature Kalman filter [13], square-root quadrature Kalman filter [14], etc.
@@
 Therefore, the method can be increasingly more complicated than EKF as the number of states increases.
-While different extensions of the KF use various ways to approximate the first two moments of the states after a nonlinear transformation, they all follow the same framework as in the linear KF.
-In this paper, this framework is referred to as “the conventional framework” or “the old framework”.
-In such a framework, the state estimations in each iteration can be divided into two steps: Predict and Update.
-In the “predict” step, the state and covariance estimations are predicted using the state transition function and estimates from the previous iteration.
-Then, in the “update” step, the measurements correct the state predictions with a feedback gain (also known as the Kalman gain), which is designed to minimize the trace of the states’ covariance matrix after the update.
-After the states are updated, the state covariance matrix is also updated, ending the present iteration.
-Note that the covariance matrix will have the smallest possible trace after the update since that’s how the Kalman gain is selected.
-While such a framework can guarantee the optimality of the state estimation for linear systems, it often leads to overconfident covariance estimation and less accurate state estimation when applied to nonlinear systems.
-The primary reason for such a phenomenon is the assumption that the Kalman gain minimizes the trace of the state covariance matrix, which is generally not guaranteed for systems with nonlinear measurement functions.
-In fact, as we will show later, the “update” step can sometimes make the state estimation of the nonlinear system less accurate, so the assumption about the optimality of the “update” step can be extremely harmful, as it makes the algorithm become overconfident in an inaccurate estimation.
-The problem described above will become increasingly significant as the measurement noise decreases, and the measurements play a more critical role in the state estimation [14].
-In some circumstances, it can even lead to a highly undesirable and counterintuitive phenomenon: the actual estimation error increases as the measurement noise decreases.
-The primary cause of this phenomenon is that the actual estimation errors drop much slower than the estimated standard deviations, and such overconfidence prevents the filter from correcting itself in later iterations.
-As the development of technology makes all types of sensors more and more accurate, we can reasonably predict that this problem will become increasingly severe if we stick to the conventional framework for nonlinear KFs.
-Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence.
+While different extensions of the Kalman filter (KF) approximate the first- and second-order moments of the state after a nonlinear transformation in different ways, they all retain the same algorithmic structure as the linear KF.
+In this paper, we refer to this structure as the “conventional framework” or “old framework”.
+In each iteration, the estimation process consists of two stages: Predict and Update.
+In the “predict” step, the state mean and covariance are propagated through the process model using the estimates from the previous iteration.
+In the “update” step, the incoming measurement is used to correct the state prediction through a feedback term known as the Kalman gain.
+In the linear KF, this gain is analytically derived to minimize the trace of the a posteriori covariance matrix, and the corresponding covariance is then updated consistently with this optimal gain.
+While this framework is the optimal linear state estimator for linear systems, applying it to nonlinear measurement functions often results in overconfident covariance estimates and degraded state accuracy.
+The fundamental issue is that the Kalman gain is computed as though it still minimizes the posterior covariance trace, even though this optimality no longer holds when the required moments are only approximated.
+As we will show later, the “update” step may even increase the estimation error in nonlinear systems.
+Consequently, the implicit assumption of optimality in the “update” step can be very harmful, because the filter becomes increasingly confident in estimates that are systematically biased or inaccurate.
+The overconfidence phenomenon described above is well known in the literature.
+Graphical illustrations of EKF overconfidence have been provided in [15], and in the context of simultaneous localization and mapping (SLAM), researchers have repeatedly observed that nonlinear KFs tend to produce overconfident, rather than conservative, state estimates, especially after a long time [16, 17, 18].
+Moreover, this overconfidence issue becomes more pronounced when the measurement noise variance is small [19, 20], because the linearization (or other approximation) error then constitutes a larger fraction of the total error budget, making the filter’s covariance underestimate more severe.
+In some cases, this leads to a highly undesirable and counterintuitive behavior: the actual estimation error increases as the measurement noise decreases.
+The underlying reason is that the actual estimation errors decrease much more slowly than the estimated standard deviations, so the filter becomes overconfident and is less able to correct itself in later iterations.
+As sensor technology continues to improve and measurement noise is further reduced across many applications, we can reasonably expect this problem to become even more severe if we continue to rely on the conventional framework for nonlinear KFs.
+Unfortunately, so far, there is no general solution that can effectively prevent the nonlinear KF from overconfidence and improve its estimation accuracy.
 Although it is possible to mitigate the problem by approximating the measurement function at higher orders (e.g., using EKF2 instead of EKF), these algorithms take a significantly longer computational time, and the problem will still, in general, exist.
@@
 This assumption may not be reasonable because the predicted states can be very far away from the updated states, and ignoring higher-ordered terms can cause significant errors in the “update” step.
-In this study, we proposed a new framework that solves the problem of overconfidence.
+Apart from modifying the structure of the filter itself, another widely used strategy for mitigating filter overconfidence is covariance inflation [22, 23].
+In this approach, the estimated state covariance is deliberately enlarged, and the associated inflation factor is adjusted to compensate for under-dispersion and errors caused by nonlinearity.
+While covariance inflation can alleviate overconfidence, its performance depends heavily on the choice of the inflation factor, whose value typically requires extensive fine-tuning.
+This tuning can be computationally costly and problem-dependent, thereby undermining the universality of the approach [24, 25].
+In this study, we propose a new framework that solves the problem of overconfidence.
 The key idea of the new framework is to add a step that reevaluates the actual effect of the Kalman gain after the state update.
@@
 The conclusions are drawn in Section 4.
-The detailed simulation setup and the detailed algorithms we used for validation can be found in the appendix.
 2 Method
 2.1 General formulation of nonlinear Kalman filters
-Consider a system with the following dynamics:
-$x_{k}=f(x_{k-1},u_{k-1})+w_{k-1}$ (1)
-where $x_{k}$ are the actual states at the time step $k$, $u_{k}$ are the inputs at the time step $k$, and $w_{k}$ are the process noises that follow a zero-mean multivariate distribution with covariance $Q_{k}$.
-The estimations of the states made in the previous step are denoted as $\hat{x}_{k-1|k-1}$.
-Since the actual states are unknown, the KF assumes that the estimation errors $\hat{x}_{k-1|k-1}-x_{k-1}$ follow a zero-mean multivariate distribution with covariance $P_{k-1|k-1}$.
-Whichever nonlinear KF is used, the first step of each iteration is called “predict”, which predicts the states using (1) and the estimations made in the previous time step.
-The predicted states are denoted as $\hat{x}_{k|k-1}$, and the errors $\hat{x}_{k|k-1}-x_{k}$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{k|k-1}$.
-The second step of each iteration for any nonlinear KF is called “update”, which corrects the predicted states based on the measurements. Suppose that the measurements of the states are as follows:
-$z_{k}=h(x_{k})+v_{k}$ (2)
-where $z_{k}$ are the measurements at time step $k$, and $v_{k}$ are the measurement noises that follow a zero-mean multivariate distribution with covariance $R_{k}$.
-Given (2) and the predicted states, the measurements can also be predicted using a particular method specific to the type of nonlinear KFs.
-The predicted measurements are denoted as $\hat{y}_{k|k-1}$, whose errors $\hat{y}_{k|k-1}-h(x_{k})$ are assumed to follow a zero-mean multivariate distribution with covariance $P_{y,k}$.
-The actual cross-covariance matrix between the predicted states and the predicted measurements is denoted as $P_{xy,k}$.
-The differences between the predicted and actual measurements are called the measurement residuals, denoted as $\tilde{y}_{k}$.
-Namely:
-$\tilde{y}_{k}=z_{k}-\hat{y}_{k|k-1}$ (3)
-If we assume that the measurement noises are independent from the predicted states, then $\tilde{y}_{k}$ should have zero mean, and their covariance $S_{k}$ can be calculated by:
-$S_{k}=P_{y,k}+R_{k}$ (4)
-KFs use the following equation to update the state estimation:
-$\hat{x}_{k|k}=\hat{x}_{k|k-1}+K\tilde{y}_{k}$ (5)
-where $\hat{x}_{k|k}$ are the updated states, and $K$ is the Kalman gain. With (35), the covariance of $\hat{x}_{k|k}$ can be calculated as [16]:
-$P_{k|k}=P_{k|k-1}+KS_{k}K^{T}-P_{xy,k}K^{T}-KP_{xy,k}^{T}$ (6)
-KFs select the Kalman gain to minimize the trace of $P_{k|k}$. By taking the derivative of tr$(P_{k|k})$ in (6), we find that the trace is minimized when:
-$K_{op}=P_{xy,k}S_{k}^{-1}$ (7)
-where $K_{op}$ is the theoretical optimal value of the Kalman gain. When (7) is satisfied, (6) becomes:
-$P_{k|k,op}=P_{k|k-1}-K_{op}S_{k}K_{op}^{T}$ (8)
-Equations (58) illustrate the key idea of KFs, which is to update the states in a way that minimizes the trace of the state covariance matrix after the update.
-However, it should be noted that these equations cannot be directly applied to systems with nonlinear measurement functions.
-The primary reason is that $P_{y,k},S_{k}$, and $P_{xy,k}$ can only be calculated if the actual state and the distribution of the states are known.
-Since neither of these holds in practice, the actual values for these covariance matrices are also unknown.
-In fact, as we will show next, some over-simplifications regarding these matrices are the key factors that can make the conventional framework for nonlinear KFs overconfident and inaccurate.
+Consider a discrete-time system with nonlinear dynamics
+$x_{k}=f(x_{k-1},u_{k-1})+w_{k-1},$ (1)
+where $x_{k}$ denotes the (unknown) true states at time step $k$, $u_{k-1}$ is the known input applied between time steps k1k-1 and $k$, and $w_{k-1}$ is the process noise, assumed zero-mean with covariance $Q_{k-1}$, and independent of $x_{k-1}$ and past measurements.
+At time step k1k-1, the filter maintains an a posteriori state estimate $\hat{x}_{k-1|k-1}$ with associated error covariance $P_{k-1|k-1}$.
+For any nonlinear KF variant, the first step of each iteration is the “predict” step, which propagates $\hat{x}_{k-1|k-1}$ and $P_{k-1|k-1}$ through the process model (1) to obtain the a priori quantities $\hat{x}_{k|k-1}$ and $P_{k|k-1}$.
+The predicted state estimate at time step $k$ is denoted by $\hat{x}_{k|k-1}$, and the corresponding estimation error $x_{k}-\hat{x}_{k|k-1}$ is assumed to follow a zero-mean multivariate distribution with covariance $P_{k|k-1}$.
+Formally,
+$\begin{cases}\mathbb{E}[x_{k}-\hat{x}_{k|k-1}|z_{1:k-1}]=0,\\ P_{k|k-1}=\mathrm{Var}(x_{k}-\hat{x}_{k|k-1}|z_{1:k-1}),\end{cases}$ (2)
+where $z_{1:k-1}$ denotes the past measurements.
+In the subsequent analysis, we work conditionally on the observation history.
+In particular, for any fixed realization of $z_{1:k-1}$, $P_{k|k-1}$ is a deterministic function of that realization; likewise, $P_{k-1|k-1}$ is deterministic given $z_{1:k-1}$.
+Unless stated otherwise, we suppress the conditioning in the notation for all means, covariances, and cross-covariances (e.g., $P_{k|k-1}$, $P_{z,k}$, $P_{xz,k}$, and $S_{k}$, which will be formally defined later).
+The second step of each iteration for a nonlinear KF is called the “update”, which corrects the predicted state based on the new measurement. The measurement model is
+$z_{k}=h(x_{k})+v_{k},$ (3)
+where $z_{k}$ is the measurement at time step $k$, and $v_{k}$ is the measurement noise, assumed zero-mean with covariance $R_{k}$.
+Given the predicted state, the measurement can be predicted (using a specific nonlinear KF algorithm), and we denote this predicted measurement by $\hat{z}_{k|k-1}$.
+The corresponding (noise-free) prediction error $h(x_{k})-\hat{z}_{k|k-1}$ is assumed zero-mean with covariance $P_{z,k}$.
+The cross-covariance between the state prediction error and the measurement prediction error is denoted by $P_{xz,k}$.
+The difference between the actual and predicted measurements is called the measurement residual (or innovation), denoted by
+$\tilde{z}_{k}:=z_{k}-\hat{z}_{k|k-1}.$ (4)
+Since both $v_{k}$ and $h(x_{k})-\hat{z}_{k|k-1}$ are zero-mean, the innovation $\tilde{z}_{k}$ is also zero-mean.
+If, in addition, the measurement noise $v_{k}$ is assumed to be independent of the (noise-free) prediction error $h(x_{k})-\hat{z}_{k|k-1}$, the covariance of the innovation is
+$S_{k}=P_{z,k}+R_{k}.$ (5)
+The KF update for the state estimate is
+$\hat{x}_{k|k}=\hat{x}_{k|k-1}+K_{k}\tilde{z}_{k},$ (6)
+where $\hat{x}_{k|k}$ is the updated state estimate and $K_{k}$ is the Kalman gain at time step $k$. Using (4)–(6), the a posteriori error covariance can be written as [26]
+$P_{k|k}=P_{k|k-1}+K_{k}S_{k}K_{k}^{T}-P_{xz,k}K_{k}^{T}-K_{k}P_{xz,k}^{T}.$ (7)
+The Kalman gain $K_{k}$ is chosen to minimize $\mathrm{tr}(P_{k|k})$. By differentiating $\mathrm{tr}(P_{k|k})$ in (7) with respect to $K_{k}$, we obtain the optimal gain
+$K_{k,\text{op}}=P_{xz,k}S_{k}^{-1},$ (8)
+where $K_{k,\mathrm{op}}$ denotes the optimal Kalman gain. Substituting (8) into (7) yields the well-known expression
+$P_{k|k,\mathrm{op}}=P_{k|k-1}-K_{k,\mathrm{op}}S_{k}K_{k,\mathrm{op}}^{T}.$ (9)
+Equations (6)–(9) summarize the key principle of KF: the state estimate is updated in such a way that the trace of the a posteriori error covariance is minimized.
+However, for systems with nonlinear measurement functions $h(\cdot)$, these expressions involve covariance terms $P_{k|k-1}$, $P_{z,k}$, $S_{k}$, and $P_{xz,k}$ that, in general, are defined as expectations with respect to the full distribution of the state and cannot be uniquely determined from the first two moments alone.
+Since nonlinear KFs propagate only the mean and covariance of the state, this information is generally insufficient to recover the true values of these matrices, so they must be approximated.
+In fact, as we will show next, inaccurate approximations regarding these matrices are the key factors that can make the conventional framework for nonlinear KFs overconfident and inaccurate.
+Before concluding this subsection, we note that the formulation above extends directly to non-additive noise.
+Specifically, if the process noise enters the dynamics as $x_{k}=f(x_{k-1},u_{k-1},w_{k-1})$, we can define the prediction-time augmented variable $\chi_{k-1}:=[x_{k-1}^{T}\;w_{k-1}^{T}]^{T}$.
+Likewise, if the measurement noise is non-additive, $z_{k}=h(x_{k},v_{k})$, we can define the update-time augmented variable $\eta_{k}:=[x_{k}^{T}\;v_{k}^{T}]^{T}$.
+Since the first two moments of these augmented variables are determined by the first two moments of $x_{k-1}$, $w_{k-1}$, and $v_{k}$, the same algorithmic machinery (e.g., linearization or sigma-point propagation) applies to $\chi_{k-1}$ and $\eta_{k}$ to approximate the first two moments of $x_{k}$ and $z_{k}$.
+Importantly, $w_{k-1}$ and $v_{k}$ are not treated as persistent states; they are i.i.d.
+zero-mean draws at each step and are included in the augmentation only to capture their nonlinear effect on the moments.
 2.2 The conventional framework and its problem
-While calculating $P_{y,k},S_{k}$, and $P_{xy,k}$ at the actual state is impossible, it is possible to approximate their values based on the predicted states.
-Different nonlinear filters (e.g.
-UKF, EKF2, CKF) take different approaches to approximating these covariance matrices.
-In the subsequent analysis, we denote these approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ since they are all estimated based on the prediction.
-Although the detailed process of calculating ${\color[rgb]{1,0,0}P_{y,k|k-1}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ differs, all existing types of nonlinear KFs directly replace $S_{k}$ and $P_{xy,k}$ in (7) and (8) with the approximated values.
-Namely, the Kalman gain is selected as:
-${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ (9)
-and the state covariance matrix is updated by:
-${\color[rgb]{1,0,0}P_{k|k,est}}$ $=P_{k|k-1}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k-1}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (10)
-$=P_{k|k-1}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
-On the other hand, the actual value of the updated state covariance matrix ${\color[rgb]{1,0,0}P_{k|k,ac}}$ given the selected Kalman gain in (9) can be calculated by substituting (9) into (6).
-Namely:
-${\color[rgb]{1,0,0}P_{k|k,ac}}=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}S_{k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$ (11)
-$\quad-P_{xy,k}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}P_{xy,k}^{T}$
-In other words, ${\color[rgb]{1,0,0}P_{k|k,ac}}$ plays the role of $P_{k|k}$ in (6), and ${\color[rgb]{1,0,0}P_{k|k,est}}$ plays the role of $P_{k|k,op}$ in (8).
-Intuitively, tr$({\color[rgb]{1,0,0}P_{k|k,ac}})$ tends to be larger than tr$({\color[rgb]{1,0,0}P_{k|k,est}})$ since tr$(P_{k|k,op})=\min$tr$(P_{k|k})\leq$ tr$(P_{k|k})$.
-However, this relationship of size is not always preserved after the approximations.
-For example, according to (10) and (11), when $P_{xy,k}=S_{k}=1$ and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}={\color[rgb]{1,0,0}S_{k|k-1}}=0.9$, we have ${\color[rgb]{1,0,0}P_{k|k,est}}-{\color[rgb]{1,0,0}P_{k|k,ac}}=0.1>0$.
-Nevertheless, in most cases in practice [17, 18], we do find that tr$({\color[rgb]{1,0,0}P_{k|k,ac}})>$tr$({\color[rgb]{1,0,0}P_{k|k,est}})$, meaning that using (10) will often (although not always) underestimate the actual covariance.
-In order to make this statement more rigorous in math, we consider the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$.
-Specifically, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k-1}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k-1}}$ is a random positive definite matrix.
-Note that we mark all the random covariance matrices in red in this section for clarity.
-Additionally, we assume that the “noises” that arise in the approximations are unbiased, so these random matrices should satisfy:
-$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k-1}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k-1}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}]=P_{xy,k}\end{cases}$ (12)
-These additional definitions and assumptions are made so that we can consider the “average” case and give a more mathematically rigorous definition of the overconfidence caused by (10).
-Namely, we have the following theorem.
+While calculating the actual values of $P_{k|k-1},P_{z,k}$, $S_{k}$, and $P_{xz,k}$ is generally impossible by just using the first two moments of the states, it is possible to make some approximations.
+Without loss of generality, this problem of approximation can be formulated as follows.
+The general approximation problem in nonlinear Kalman filters: Given an estimated mean $\hat{X}\approx\mathbb{E}[X]$ and an estimated covariance ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{x,\mathrm{est}}}\approx\mathrm{Var}(X)$ of the state $X$, and a nonlinear mapping $Y=g(X)$, we seek to approximate $\mathbb{E}[Y]$, $\mathrm{Var}(Y)$, and $\mathrm{Cov}(X,Y)$.
+The solution can be generally written as
+$\begin{cases}\hat{Y}:=\Phi_{e}\bigl(g;\hat{X},P_{x,\mathrm{est}}\bigr)\approx\mathbb{E}[Y],\\[4.0pt] {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{y,\mathrm{est}}}:=\Phi_{y}\bigl(g;\hat{X},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{x,\mathrm{est}}}\bigr)\approx\mathrm{Var}(Y),\\[4.0pt] {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xy,\mathrm{est}}}:=\Phi_{xy}\bigl(g;\hat{X},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{x,\mathrm{est}}}\bigr)\approx\mathrm{Cov}(X,Y),\end{cases}$ (10)
+where $\Phi_{e}$, $\Phi_{y}$, and $\Phi_{xy}$ denote the approximation mappings determined by the chosen nonlinear KF (e.g., EKF, UKF, CKF).
+The nonlinear mapping $g$ may represent either the state transition function $f$ (prediction step) or the measurement function $h$ (update step), so this approximation problem is typically solved at least twice in each iteration of a nonlinear KF.
+Specifically, $P_{k|k-1},P_{z,k}$, $S_{k}$, and $P_{xz,k}$ can be respectively approximated by:
+$\begin{cases}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}:=\Phi_{y}(f;\hat{x}_{k-1|k-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k-1|k-1,\text{est}}})+Q_{k-1},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}}:=\Phi_{y}(h;\hat{x}_{k|k-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}),\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}}+R_{k},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}:=\Phi_{xy}(h;\hat{x}_{k|k-1},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}).\end{cases}$ (11)
+Note that the subscript k|k1k|k-1 means that the covariance matrices are all estimated based on the predicted states $\hat{x}_{k|k-1}$.
+Although the corresponding $\Phi_{y},\Phi_{s}$, and $\Phi_{xy}$ differ, all existing types of nonlinear KFs directly replace $P_{k|k-1}$, $S_{k}$, and $P_{xz,k}$ in (8) and (9) with the approximated values in (11).
+Namely, the Kalman gain is selected as
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}},$ (12)
+and the state covariance matrix is updated by
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{est}}}$ $={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}^{T}}$ (13)
+$={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}.$
+On the other hand, the actual value of the updated state covariance matrix $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}$} given the selected Kalman gain in (12) can be calculated by substituting (12) into (7).
+Namely,
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}}=P_{k|k-1}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}S_{k}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}$ (14)
+$\quad-P_{xz,k}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}P_{xz,k}^{T}.$
+The covariance estimation error after the state update can therefore be calculated as
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err}}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{est}}}.$ (15)
+Noticing that the covariance estimation error before the state update is
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{err}}}:=P_{k|k-1}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}},$ (16)
+the change in the covariance estimation error is
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{err}}}$ (17)
+$\quad={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est}}}|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}=P_{k|k-1}}.$
+[…]
+To formalize this idea, we interpret the approximated covariance matrices $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}$}, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$}, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} produced by a nonlinear KF as (random) estimators of the actual quantities $P_{z,k}$, $S_{k}$, and $P_{xz,k}$.
+More precisely, we regard $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} as a random matrix, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}$} as a random positive semidefinite matrix, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} as a random positive definite matrix.
+For clarity, all such random covariance matrices are marked in red in this section.
+Additionally, we assume that these approximated covariance matrices are unbiased estimators and satisfy
+$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}}]=P_{z,k},\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}]=S_{k},\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}]=P_{xz,k}.\end{cases}$ (18)
+These additional definitions and assumptions are introduced to facilitate consideration of the “average” case and provide a more mathematically rigorous definition of the overconfidence caused by (13).
+Specifically, (18) should be understood as follows.
+If a systematic, nonzero-mean bias in these approximations were known a priori, it would be natural to modify the filter so as to subtract this bias, and to analyze only the residual approximation error.
+Thus, it is reasonable to model the remaining covariance errors as zero-mean random matrices.
+Note that this viewpoint is conceptually analogous to the standard treatment of deterministic biases in augmented-state or separate-bias Kalman filters, where unknown biases are initialized as zero-mean random variables.
+Namely, any known mean component is absorbed into the system model, and only the zero-mean uncertainty is left for the filter to estimate.
+With (18), we have the following theorem about the overconfidence of nonlinear Kalman filters.
 Theorem 1.
-Suppose that ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ and ${\color[rgb]{1,0,0}S_{k|k-1}}$ are random matrices that satisfy (12).
-${\color[rgb]{1,0,0}S_{k|k-1}}$ is symmetric and positive definite.
-Then, $P_{k|k,est}$ calculated from (10) satisfies:
-$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}]\succeq 0$ (13)
-where $\succeq$” represents positive semidefiniteness, and the equality only holds when ${\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ is constant.
+Suppose that $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} are random matrices that satisfy (18).
+$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} is symmetric and positive definite.
+Then, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}$} calculated from (17) satisfies:
+$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}}]\succeq 0,$ (19)
+where $\succeq$” represents positive semidefiniteness, and the equality only holds when ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}$ is constant almost surely.
+Note that this theorem does not require $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}$} to be independent.
 {pf}
-Let ${\color[rgb]{1,0,0}P}={\color[rgb]{1,0,0}P_{xy,k|k-1}}S_{k}^{-\frac{1}{2}}$, ${\color[rgb]{1,0,0}S}=S_{k}^{-\frac{1}{2}}{\color[rgb]{1,0,0}S_{k|k-1}}S_{k}^{-\frac{1}{2}}$. According to (12),
-$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P}]=\bar{P}=P_{xy,k}S_{k}^{-\frac{1}{2}}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S}]=I\end{cases}$ (14)
-According to (10, 11),
-${\color[rgb]{1,0,0}P_{k|k,ac}}-{\color[rgb]{1,0,0}P_{k|k,est}}$ (15)
-$=$ ${\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-2}P^{T}}-\bar{P}{\color[rgb]{1,0,0}S^{-1}P^{T}}-{\color[rgb]{1,0,0}PS^{-1}}\bar{P}^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}$
-$=$ $({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})({\color[rgb]{1,0,0}PS^{-1}}-\bar{P})^{T}+{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}$
-For any fixed unit vector $\bm{v}$, ${\color[rgb]{1,0,0}P^{T}}\bm{v}$ is a random vector with a mean of $\bar{P}^{T}\bm{v}$. According to Lemma 4 in Appendix A,
-$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}P}{\color[rgb]{1,0,0}S^{-1}}{\color[rgb]{1,0,0}P^{T}}\bm{v}]\geq\bm{v}^{T}\bar{P}\bar{P}^{T}\bm{v}.$ (16)
-Moving the right hand side to the left, the following must hold true:
-$\mathbb{E}[{\color[rgb]{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}]\succeq 0$ (17)
-Therefore, the expectation of (15) is always p.s.d.
-due to (17) and the fact that the first term in (15) is p.s.d.
-The equality only holds when ${\color[rgb]{1,0,0}PS^{-1}}$ is constant, which is equivalent to ${\color[rgb]{1,0,0}K_{est}}={\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}$ being constant (i.e., free from approximation errors).
-Theorem 1 suggests that using (10) to estimate the state covariance matrix tends to lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xy,k}$ are not the same as their actual values.
-Admittedly, in practice, the errors of the approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$ are not as simple as zero-mean random matrices.
+Let ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}S_{k}^{-\frac{1}{2}}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}=S_{k}^{-\frac{1}{2}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}S_{k}^{-\frac{1}{2}}$.
+According to (18),
+$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}]=\bar{P}=P_{xz,k}S_{k}^{-\frac{1}{2}},\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}]=I.\end{cases}$ (20)
+According to (13, 14, 18),
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err}}}$ (21)
+$=$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-2}P^{T}}-\bar{P}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}P^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}}\bar{P}^{T}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}P^{T}}$
+$=$ $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}}-\bar{P})({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}}-\bar{P})^{T}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}.$
+For any fixed unit vector $\bm{v}$, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{T}}\bm{v}$ is a random vector with a mean of $\bar{P}^{T}\bm{v}$.
+According to Lemma 4 in Appendix A,
+$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{T}}\bm{v}]\geq\bm{v}^{T}\bar{P}\bar{P}^{T}\bm{v}.$ (22)
+Moving the right hand side to the left, we get
+$\mathbb{E}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}P^{T}}-\bar{P}\bar{P}^{T}]\succeq 0.$ (23)
+Therefore, the expectation of (21) is always positive semi-definite.
+due to (23) and the fact that the first term in (21) is positive semi-definite.
+The equality only holds when $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}PS^{-1}$} is constant almost surely, which is equivalent to ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}$ being constant (i.e., free from approximation errors) almost surely.
+Theorem 1 suggests that using (13) to estimate the state covariance matrix tends to lead to overconfident covariance estimation when the estimated $S_{k}$ and $P_{xz,k}$ are not the same as their actual values.
+Admittedly, in practice, the errors of the approximated covariance matrices ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$} may not be as simple as zero-mean random matrices.
 However, Theorem 1 does give us a structured explanation as to why nonlinear KFs often underestimate the state covariance matrix.
 Note that an overconfident state covariance estimation can be very detrimental since it prevents the filter from correcting itself in subsequent iterations, thereby affecting the state estimation accuracy.
-Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xy,k}$ are related to the actual state, which is unknown.
+Again, this problem is unavoidable because the actual values of $S_{k}$ and $P_{xz,k}$ are related to the actual distribution of the states, which is unknown.
 Therefore, this problem exists for all types of nonlinear KFs as long as (13) is used to update the state covariance estimation.
 In plain language, the problem of the conventional framework originates from the assumption that the selected Kalman gain is optimal.
-However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the covariance matrix after the update.
-The reason is that the states will move to a new location after the update, and the previous approximations may be inaccurate.
+However, the Kalman gain, which was optimized based on the approximated measurement function near the predicted states, generally will not minimize the trace of the actual covariance.
+The reason is that the actual states can be far from the predicted states and the approximations made in the “predict” step may be inaccurate.
 An example of such a problem is shown in Fig.
-1, where we used an EKF to estimate a plane’s position.
+1, where we use an EKF to estimate a plane’s position.
 In the example, the state is the plane’s position along the x-axis, and the measurement function is the altitude of the ground beneath the plane, which is a nonlinear function of the state.
@@
 1”, in Fig.
-1), the EKF overestimates the position of the plane, and the absolute value of the state estimation error $|\hat{x}_{k|k}-x_{k}|$ becomes higher after the “update” step.
+1), the EKF overestimates the position of the plane, and the absolute value of the state estimation error $|x_{k}-\hat{x}_{k|k}|$ becomes higher after the “update” step.
 Meanwhile, the estimated standard deviation of the state, $\sigma_{k|k}$, becomes smaller after the “update” step.
@@
 As shown in Fig.
-1, with this additional step, the covariance matrix may not have the smallest possible trace after the update; instead, it will honestly reflect the effect of the Kalman gain on the state covariance matrix.
+1, with this additional step, the estimated covariance matrix may not have the smallest possible trace after the update; instead, it will honestly reflect the effect of the Kalman gain on the actual state covariance matrix.
 This is visualized by the wider yellow distribution in the “recalibrate” step, which now covers the true state.
-Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $\hat{x}_{k|k}$ and re-estimate the matrices $P_{y,k},S_{k}$ and $P_{xy}$.
-We denote these re-approximated covariance matrices as ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ since they are all estimated based on the updated states.
-In the new framework, the state covariance matrix is updated by replacing $S_{k}$ with ${\color[rgb]{1,0,0}S_{k|k}}$ and $P_{xy,k}$ with ${\color[rgb]{1,0,0}P_{xy,k|k}}$ in (11):
-${\color[rgb]{1,0,0}P_{k|k,new}}=P_{k|k-1}+{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}$ (18)
-$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}K_{est}^{T}}-{\color[rgb]{1,0,0}K_{est}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
-$=P_{k|k-1}+{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}$
-$\quad-{\color[rgb]{1,0,0}P_{xy,k|k}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}-{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k}^{T}}$
+Mathematically, “recalibrate” means to re-approximate the measurement functions around the updated states $\hat{x}_{k|k}$ and re-estimate the matrices $P_{z,k},S_{k}$ and $P_{xz}$.
+We denote these re-approximated covariance matrices as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} since they are all estimated based on the updated states.
+Similar to (11), these covariance matrices are calculated by
+$\begin{cases}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}}:=\Phi_{y}(h;\hat{x}_{k|k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}),\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}}+R_{k},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}:=\Phi_{xy}(h;\hat{x}_{k|k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}).\end{cases}$ (24)
+In the new framework, the state covariance matrix is updated by replacing $S_{k}$ with $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}$} and $P_{xz,k}$ with $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} in (14):
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}={{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{\text{est}}^{T}}$ (25)
+$\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{\text{est}}^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}K_{k,\text{est}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}^{T}}$
+$={{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}$
+$\quad-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}^{T}}.$
+In the new framework, the covariance estimation error after the state update can then be calculated as
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err,new}}}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\text{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}.$ (26)
+The change in the covariance estimation error after the update is
+${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{err}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{err,new}}}$ (27)
+$\quad={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}=P_{k|k-1}}.$
 Note that the proposed new framework selects the same Kalman gain and does the same state update as the conventional one; the only difference is that the state covariance update equation is now a direct approximation of (14), which is the actual state covariance regardless of the selection of the Kalman gain $K$.
 Intuitively, this can prevent the KF from overconfidence because the equation no longer relies on the problematic assumption that the Kalman gain is optimal.
-To explain this advantage more rigorously, we consider the newly approximated covariance matrices ${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$ as noisy versions of $P_{y,k},S_{k}$, and $P_{xy,k}$.
-Similar to our previous assumptions about ${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, we assume that ${\color[rgb]{1,0,0}P_{xy,k|k}}$ is a random matrix, ${\color[rgb]{1,0,0}P_{y,k|k}}$ is a random positive semidefinite matrix, and ${\color[rgb]{1,0,0}S_{k|k}}$ is a random positive definite matrix.
-Additionally, we assume that the “noises” that arise in these approximations are unbiased, so these new random matrices should also satisfy:
-$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{y,k|k}}]=P_{y,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}S_{k|k}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k}}]=P_{xy,k}\end{cases}$ (19)
+To explain this advantage more rigorously, we consider the newly approximated covariance matrices ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} produced by a nonlinear KF as (random) estimators of the actual quantities $P_{z,k}$, $S_{k}$, and $P_{xz,k}$.
+Similar to our previous assumptions about ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$}, we regard $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$} as a random matrix, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}$} as a random positive semidefinite matrix, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}$} as a random positive definite matrix.
+Additionally, we assume that these approximated covariance matrices are unbiased estimators and satisfy:
+$\begin{cases}\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}}]=P_{z,k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}]=S_{k}\\ \operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}]=P_{xz,k}\end{cases}$ (28)
 With these additional definitions and assumptions, we are able to give the following theorem about the advantage of the new framework:
 Theorem 2.
-Assume that $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{xy,k|k}})=P_{xy,k}$ and $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k-1}})=\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$.
-If ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$, ${\color[rgb]{1,0,0}P_{xy,k|k}}$, ${\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}S_{k|k}}$ are all independent from each other, ${\color[rgb]{1,0,0}P_{k|k,new}}$ calculated in (18) is an unbiased estimation of ${\color[rgb]{1,0,0}P_{k|k,ac}}$.
-Namely, we have:
-$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})=0$ (20)
+Assume that the conditions (18) and (28) hold.
+Further assume that the random pair $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}})$ is independent of the random pair $({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}})$.
+Then $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}$}, as defined in (27), satisfies
+$\operatorname{\mathbb{E}}\big[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}}\big]=0.$ (29)
+Consequently, if in addition $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\text{est}}}]=P_{k|k-1}$, then $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}$} calculated in (25) is an unbiased estimator of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}$}.
+Namely,
+$\operatorname{\mathbb{E}}\big[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{ac}}}\big]=0.$ (30)
 {pf}
 With the assumptions made in Theorem 2, we have:
-$\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}P_{k|k,new}}-{\color[rgb]{1,0,0}P_{k|k,ac}})$ (21)
-$=$ $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}S_{k|k}}-S_{k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
-$-\operatorname{\mathbb{E}}[({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k}){\color[rgb]{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}P_{xy,k|k-1}^{T}}]$
-$-\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}P_{xy,k|k-1}}{\color[rgb]{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}P_{xy,k|k}}-P_{xy,k})^{T}]$
-$=$ 00 0
-The last equality equals zero because the terms in parentheses are zero, due to the unbiased (e.g. $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}S_{k|k}})=S_{k}$) and independence assumptions.
-Theorem 2 suggests that the proposed new framework can generally prevent the filter from being overconfident if the approximation errors of the measurement function at the predicted and updated states are independent of each other.
+$\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta P_{k|k,\mathrm{err,new}}}]$ (31)
+$=$ $\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}-S_{k}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}]$
+$-\operatorname{\mathbb{E}}[({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}-P_{xz,k}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}^{T}}]$
+$-\operatorname{\mathbb{E}}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}^{-1}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}}-P_{xz,k})^{T}]$
+$=$  $0.$
+The last equality holds because the terms in parentheses are zero, due to the unbiasedness (i.e., $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}})=S_{k}$ and $\operatorname{\mathbb{E}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}})=P_{xz,k}$) and independence assumptions.
+Theorem 2 suggests that the proposed new framework can generally prevent the filter from being overconfident in the “update” step if the approximation errors of the measurement function at the predicted and updated states are independent of each other.
 Admittedly, this assumption about the independence of the approximation errors is very hard to verify and may not hold in practice.
-Nonetheless, the theorem does explain why recalibrating the covariance matrices $P_{y,k},S_{k}$, and $P_{xy,k}$ after the state update is advantageous for accurate state covariance estimation.
+Nonetheless, the theorem does explain why recalibrating the covariance matrices $P_{z,k},S_{k}$, and $P_{xz,k}$ after the state update is advantageous for accurate state covariance estimation.
 Specifically, the advantage comes from our replacement of (13) with (25), which highlights the fundamental idea of the new framework.
 That is, instead of reducing the updated state covariance matrix to the smallest possible value, we re-approximate the system after the state update to estimate the actual effect of the Kalman gain more accurately.
-In fact, in the new framework, the trace of the state covariance matrix may increase after the state update and recalibration if the system is strongly nonlinear and the predicted states have high variances, just like the example shown in Fig.
+In nonlinear KFs, the update can sometimes make the state estimate less accurate and increase the actual state covariance, as shown in Fig.
 1.
-This phenomenon means that the update fails to make the estimation more accurate and should be withdrawn.
-To address this situation, we also add a “back out” step to the framework, which allows the state estimator to back out of the update when the update is unhelpful.
+If this situation were detectable, a natural reaction would be to withdraw the update and use the more accurate predicted state (and the predicted state covariance) as the final output.
+Unfortunately, such detection is impossible based solely on the estimated state covariance matrix in the conventional KF framework, since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est}}$} calculated in (13) is guaranteed to be smaller than or equal to $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}.
+In the new framework, however, the additional “recalibrate” step makes such detection possible, which motivates us to introduce a “back out” step.
+In general, “back out” means to withdraw unhelpful updates and output the state estimate with a lower estimated mean-squared error.
+Specifically, by the end of the recalibrate step, the filter has two candidate estimates of the current state: the predicted state $\hat{x}_{k|k-1}$ with covariance $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}, and the updated state $\hat{x}_{k|k}$ with covariance $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}$}.
+Since the trace of the covariance matrix equals the expected squared Euclidean error of the corresponding state estimate, we choose as the final output the estimate whose covariance has the smaller trace.
+If the “back out” step is omitted, the filter may not only produce a suboptimal estimate, but can also become highly unstable: $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k,\mathrm{est,new}}$} computed in (25) is no longer guaranteed to remain bounded by $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}, and can be several orders of magnitude larger than $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{k|k-1,\mathrm{est}}$}, which is clearly undesirable.
 The pseudo-code of the new framework is shown in Algorithm 1.
 Note that the new framework degrades to the conventional framework when the “recalibrate” and “back out” steps are replaced with (13).
+Moreover, when all measurement functions are linear, the approximations of $P_{z,k}$, $S_{k}$, and $P_{xz,k}$ become identical before and after the recalibration step.
+In this case, the proposed framework is exactly equivalent to the conventional framework and preserves its optimality for linear systems.
 Algorithm 1 The proposed framework for nonlinear Kalman filters
@@
 2:Initialization:
-3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
-4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
+[…]
+[…]
 5:for every time step $k$ do
@@
 9:Update:
-10:     Estimate $\bm{\hat{y}}_{k|k-1},\bm{P}_{xy,k|k-1}$ and $\bm{P}_{y,k|k-1}$
+10:  Estimate $\bm{\hat{z}}_{k|k-1},\bm{P}_{xz,k|k-1}$ and $\bm{P}_{z,k|k-1}$
 11:  (based on $\bm{\hat{x}}_{k|k-1},\bm{P}_{k|k-1},$ and $\bm{h}$)
-12:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
-13:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
-14:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+12:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
+13:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
+14:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
 15:Recalibrate:
-16:     Estimate $\bm{P}_{xy,k|k}$ and $\bm{P}_{y,k|k}$
+16:  Estimate $\bm{P}_{xz,k|k}$ and $\bm{P}_{z,k|k}$
 17:  (based on $\bm{\hat{x}}_{k|k},\bm{P}_{k|k-1},$ and $\bm{h}$)
-18:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
-19:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+18:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
+19:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
 20:Back out:
@@
 25:end for
-26:where, $\bm{\hat{x}}$ is the estimated states, $\bm{P}$ is the states’ covariance matrix, $\bm{\hat{y}}$ is the estimated measurements, $\bm{P}_{xy}$ is the covariance between the states and the measurements, $\bm{P}_{y}$ is the covariance of the estimated measurements, $\bm{S}$ is the innovation (or residual) covariance, $\bm{K}$ is the Kalman gain.
-The subscript $k|k-1$ represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
-It is worth mentioning that when the measurement functions are all linear, they would require no approximation.
-Or, to be more precise, the approximation of the measurement functions would become independent of the states.
-In this case, the approximations of $P_{y,k},S_{k}$, and $P_{xy,k}$ will not change before and after the recalibration, and our framework will be equivalent to the conventional framework.
-To give readers an intuitive understanding of the effect of the “recalibrate” and “back out” steps, in Fig.
-2, we showed an example of using different types of nonlinear KFs to do the state update given the measurement $y=x^{3}/3-x^{2}/8-x+1.5383=0$.
-The measurement function is represented by the thick gray curve in the figure.
-The predicted state is $x_{1|0}=0$, its variance is $1.5^{2}$, and the standard deviation of the measurement is $\sigma_{y}=0.01$.
-This state estimation problem is challenging because the derivative of the measurement function at the predicted state is negative, while the actual state is to the left of the predicted state.
-The state and covariance estimations made by different algorithms shown in different rows in the figure.
-Generally, the conventional framework always underestimates the actual state estimation error, which can be visualized by the narrow standard deviation estimations in Fig.
-2 that cannot capture the true state.
-By contrast, the state covariance estimation made by the new framework is much more accurate.
-Additionally, the new framework is aware that the state estimation becomes less accurate after the update because the estimated variance increases.
-Therefore, after the “recalibrate” step, the “back out” step will be triggered, and the new framework will withdraw the update and use the predicted state and the predicted state variance as the final output.
-Figure 2: The comparison of the covariance estimation after the state update and recalibration using different frameworks and different types of nonlinear Kalman filters.
+26:where, $\bm{\hat{x}}$ is the estimated states, $\bm{P}$ is the states’ covariance matrix, $\bm{\hat{z}}$ is the estimated measurements, $\bm{P}_{xz}$ is the covariance between the states and the measurements, $\bm{P}_{z}$ is the covariance of the estimated measurements, $\bm{S}$ is the innovation (or residual) covariance, $\bm{K}_{k}$ is the Kalman gain.
+The subscript k|k1k|k-1 represents the estimated value before the state update, and $k|k$ represents the estimated value after the state update.
+To provide an intuitive illustration of the roles of the “recalibrate” and “back out” steps, Fig. 2 shows a single update for several nonlinear KFs applied to the scalar measurement model
+$z=h(x)+v,\qquad h(x)=x^{3}/3-x^{2}/8-x+1.5383,$ (32)
+with the observed measurement $z_{1}=0$.
+The measurement function $h(x)$ is plotted as the thick gray curve.
+The predicted state is $x_{1|0}=0$ with variance $1.5^{2}$, and the measurement noise standard deviation is $\sigma_{y}=0.01$.
+This example is particularly challenging because $h^{\prime}(x_{1|0})<0$ while the true state lies to the left of $x_{1|0}$, so a local linearization can easily mislead the filter.
+The state and covariance estimates produced by different algorithms are shown in separate rows in the figure.
+Under the conventional framework, all methods substantially underestimate the actual estimation error, as indicated by the overly narrow $\pm 1\sigma$ intervals that fail to cover the true state.
+In contrast, within the proposed framework, the covariance estimates are much more consistent with the true error.
+In this example, the recalibrated variance correctly indicates that the update is less reliable than the prediction, so the “back out” step is triggered and the filter discards the misleading update, outputting the predicted state and its variance as the final estimate.
+If the “back out” step is removed from the framework, the estimated state becomes less accurate for all filters.
+Moreover, in this example, the CKF’s estimated variance after the recalibration is approximately $144^{2}$, which is almost 10,000 times larger than the variance of the predicted state.
+Such a “covariance explosion” further underscores the necessity of the “back out” step.
+Figure 2: The comparison of the variance estimation after the state update and recalibration using different frameworks and different types of nonlinear Kalman filters.
+The recalibrated variance of CKF (new) is approximately $144^{2}$, almost 10,000 times larger than the predicted variance, which underscores the importance of the “back out” step.
+Finally, it is insightful to interpret the proposed framework through a machine learning perspective.
+In the conventional nonlinear KF framework, the Kalman gain is chosen based on approximations constructed around the predicted state.
+These local approximations can be viewed as an analogue of a “training set”: they are reasonably representative of the underlying nonlinear system but do not capture its full complexity.
+Computing the covariance using the same approximations is then analogous to evaluating an in-sample (training) loss, which tends to be systematically smaller than the true out-of-sample (testing) loss.
+This offers an intuitive explanation for why the conventional nonlinear KF framework often yields overconfident covariance estimates.
+In contrast, our new framework computes the covariance based on approximations constructed around the updated state, which can be viewed as an analogue of a “validation set.” In this interpretation, Theorem 2 states that, when the information used to design the gain and the information used to assess the covariance are independent, the resulting covariance (or loss) estimate is unbiased.
+Even though exact independence may not hold in practice, the machine learning literature provides strong empirical evidence that incorporating a validation-like mechanism is effective in mitigating overconfidence (overfitting) and improving estimation accuracy, which is precisely the role played by the “recalibrate” step introduced in this paper.
+As for the “back out” step, it can be interpreted as reverting to the previous checkpoint when the validation error increases during training, which is also necessary and important for finding the best estimates.
 2.4 Relationship to Iterated Kalman Filter
@@
 By contrast, the new framework requires no iteration, and the re-approximation is only done once.
-Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 2), which is only possible for EKF and EKF2.
+Secondly, IKF was initially only used on EKF, and the deduction of iterated EKF (IEKF) requires an explicit approximation of the measurement functions (rows 9 and 10 in Algorithm 3 in Appendix C), which is only possible for EKF and EKF2.
 Although a few attempts have been made to extend IKF to UKF and CKF [30, 31, 33, 34], a consensus on how to do this robustly has not been reached.
@@
 Namely, IKF always updates the states and covariance matrix simultaneously using (13), resulting in overconfident state estimations.
-When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{y,k},S_{k}$, and $P_{xy,k}$ can be less accurate, which can lead to a higher degree of overconfidence.
+When the algorithm cannot converge properly, the estimation can be even worse because the estimated $P_{z,k},S_{k}$, and $P_{xz,k}$ can be less accurate, which can lead to a higher degree of overconfidence.
 In general, in terms of making a more accurate estimation, finding the optimal Kalman gain is not as important as making the covariance estimation accurate, as we will show later in the simulation results.
-Algorithm 2 The iterated extended Kalman filter with the old framework
-1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
-2:Initialization:
-3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
-4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
-5:for every time step $k$ do
-6:    Predict:
-7:     $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
-8:     $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
-9:     $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
-10:    Update:
-11:     $\bm{x}_{0}=\bm{\hat{x}}_{k|k-1}$
-12:     for $i=1$ to 1000 do
-13:         $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{x}_{i-1}}$
-14:         $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{x}_{i-1})+\bm{H}_{k|k-1}(\bm{x}_{i-1}-\bm{\hat{x}}_{k|k-1})$
-15:         $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
-16:         $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
-17:         $\bm{S}_{k|k-1,i}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
-18:         $\bm{K}_{i}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1,i}^{-1}$
-19:         $\bm{x}_{i}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{i}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
-20:         if $i>1$ and $||\bm{x}_{i}-\bm{x}_{i-1}||>||\bm{x}_{i-1}-\bm{x}_{i-2}||$ then
-21:             $\bm{S}_{k|k-1,i}=\bm{S}_{k|k-1,i-1}$
-22:             $\bm{K}_{i}=\bm{K}_{i-1}$
-23:             $\bm{x}_{i}=\bm{x}_{i-1}$
-24:             break
-25:         end if
-26:         if $\max_{j}|1-\bm{x}_{i}(j)/\bm{x}_{i-1}(j)|<0.001$ then
-27:              break
-28:         end if
-29:     end for
-30:     $\bm{S}_{k|k-1}=\bm{S}_{k|k-1,i}$
-31:     $\bm{K}=\bm{K}_{i}$
-32:     $\bm{\hat{x}}_{k|k}=\bm{x}_{i}$
-33:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}-\bm{K}\bm{S}_{k|k-1}\bm{K}^{T}$
-34:end for
-35:Note that rows 16–21 are added to prevent the state estimations from diverging.
-Namely, when the sign of divergence is detected (row 16), the newest updates of the states, Kalman gain, and measurement residual covariance matrix are withdrawn (rows 17–19).
-If such a withdrawal is not made, we found in the simulation that the estimation accuracy can be much lower.
-This characteristic also suggests that the algorithm is not very robust and stable.
 Since IKF also uses (13) to update the state covariance matrix, it can also be considered as another framework for nonlinear KFs that is similar to the conventional framework.
 A natural question is: can we combine the new framework with the IKF to get a higher estimation accuracy?
-Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states (${\color[rgb]{1,0,0}P_{y,k|k-1}},{\color[rgb]{1,0,0}S_{k|k-1}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k-1}}$) are independent of the approximations made at the updated states (${\color[rgb]{1,0,0}P_{y,k|k}},{\color[rgb]{1,0,0}S_{k|k}}$, and ${\color[rgb]{1,0,0}P_{xy,k|k}}$).
+Unfortunately, the answer is no because a prerequisite for the new framework to give accurate covariance estimation is that the approximations made at the predicted states (${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k-1}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k-1}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k-1}$}) are independent of the approximations made at the updated states (${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{z,k|k}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{k|k}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{xz,k|k}$}).
 However, after IKF converges, these two sets of approximations become the same, which violates this prerequisite in Theorem 2.
@@
 3 Results
-To demonstrate the effectiveness of the proposed framework, we validated it in five different applications suitable for nonlinear KFs.
+To demonstrate the effectiveness of the proposed framework, we validate it in five different applications suitable for nonlinear KFs.
 A brief description of the five applications is presented in Table 1, and a detailed description of the system models and parameter setup can be found in Appendix B.
-For each scenario, we combined the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF.
-We also implemented the IEKF for comparison.
-The detailed EKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 35 in Appendix C.
+For each scenario, we combine the old and proposed framework with four different types of nonlinear KFs: EKF, EKF2, UKF, and CKF.
+We also implemente the IEKF for comparison.
+The detailed EKF, IEKF, EKF2, UKF, and CKF algorithms used in the paper can be found in Algorithms 25 in Appendix C.
 All these nine $(4+4+1)$ nonlinear KFs are used for state estimation under different measurement noise setups (the covariance of the measurement noises differs in different setups).
-To guarantee the fairness of the comparison and reduce the effect of randomness, each KF is simulated 10,000 times for each measurement noise setup, and all the KFs use the same random seed during the simulation.
-This means that the random initial states’ errors, random process noises, and random measurement noises are the same for the $i^{th}$ simulation of EKF and the $i^{th}$ simulation of UKF.
-In other words, all these random noises were generated for the 10,000 runs a priori and recalled for simulating each estimation method.
+To ensure a fair comparison and reduce the impact of Monte Carlo randomness, each KF is simulated 10,000 times for each measurement-noise level.
+All filters share the same realizations of the random initial state errors, process noises, and measurement noises.
+Concretely, these random sequences are generated from the same random seed so that the $i^{\text{th}}$ simulation of each method (e.g., EKF, UKF, CKF) is driven by identical noise realizations.
 Table 1: A brief description of the five applications of nonlinear Kalman filters investigated in this paper.
@@
 Number of iterations 30 100 100 100 180
+3.1 State estimation accuracy
 The root mean squared error (RMSE) of the state estimations given by different nonlinear KFs are shown in Figs.
@@
 The estimation error is calculated as the difference between the estimated state and the true state at the end of the final iteration of each simulation.
-Note that we only showed the RMSE of two representative states for systems with more than two states.
+Note that we only show the RMSE of two representative states for systems with more than two states.
 For example, in 3D target tracking, the six states are the object’s speed and position along the three axes, and Fig.
@@
 As can be seen from the five figures, the simulation results are very similar to each other.
-Therefore, only the results of the first task, 3D target tracking, are analyzed in depth in this section.
+Therefore, we mainly analyze the first application, 3D target tracking, in depth in this section.
 Figure 3: The root mean squared error of the state estimations under different measurement noise setups (3D target tracking).
 Figure 4: The root mean squared error of the state estimations under different measurement noise setups (terrain-referenced navigation).
-Figure 5: . The root mean squared error of the state estimations under different measurement noise setups (synchronous generator state estimation).
+Figure 5: The root mean squared error of the state estimations under different measurement noise setups (synchronous generator state estimation).
 Figure 6: The root mean squared error of the state estimations under different measurement noise setups (pendulum state estimation).
@@
 On the other hand, EKF is generally slightly worse than the other three types of nonlinear KFs because it only considers the first-order terms in the Taylor expansion.
-By comparing the performance of IEKF and EKF (with and without the new framework) in Figs.
-37, we can see that IEKF does perform better than EKF with the old framework, yet it is almost always worse than EKF with the new framework despite requiring longer computational time.
-As we mentioned in the Introduction, the main reason for such a difference is that IEKF still gives overconfident covariance estimation, especially when it fails to converge, which happens quite often in practice.
-As we mentioned in the Introduction, the fundamental idea of the proposed framework is to make a more accurate covariance estimation by re-approximating the measurement functions after the state update.
-To validate if the covariance estimation really becomes more accurate in the new framework, we compared the actual and estimated RMSE of the x-axis position (the first state in the first system) in the old and new framework in Fig.
-8.
-In the figure, for each type of nonlinear KF (represented by different line colors) and framework (the old and new frameworks are respectively shown in the top and bottom subfigures), the closeness between the actual RMSE (the solid line) and estimated RMSE (dotted line) represents the accuracy of the state covariance estimation.
-Apparently, when the standard deviations of the two measurements are lower than 0.1 meters, the conventional framework always significantly underestimates the covariance, sometimes by several orders of magnitude, regardless of the type of nonlinear KFs used.
-On the other hand, for the new framework, only the RMSE of EKF is underestimated, and the covariance estimation of other types of nonlinear KFs matches the actual value very well.
-This, again, can be explained by the fact that EKF neglects the second and higher-order information of the measurement functions, so its covariance estimation is less accurate than other methods.
-Nevertheless, the covariance estimation of EKF is still much more precise when combined with the new framework.
-Figure 8: The actual and estimated root mean squared error of the x-axis position estimation.
-Another interesting observation of Fig.
-8 is that the estimated RMSE of the x-axis position (the dashed lines) is very similar for different combinations of frameworks and KFs.
-This suggests that the accuracy of the estimation mainly relates to whether the covariance estimation is accurate or not, which leads to the conclusion mentioned at the end of the Introduction — to achieve a more accurate estimation, finding the optimal Kalman gain is not as important as accurately estimating the covariance.
+By comparing the performance of IEKF and EKF (with and without the proposed framework) in Figs.
+37, we observe that although the IEKF outperforms the EKF under the conventional framework, it is consistently inferior to the EKF implemented with the new framework, while also incurring higher computational cost.
+As discussed in the Introduction, this discrepancy arises because the IEKF continues to produce overconfident covariance estimates, particularly in cases where the iteration fails to converge — a situation that occurs frequently in practice.
+The necessity of the “back out” step is illustrated in Fig.
+8, which compares the RMSEs of pendulum state estimation (Application 3) with and without this step.
+The results show that, without the “back out” step, all four nonlinear KF variants become significantly less accurate when the measurement noise is low.
+This degradation can be attributed to the fact that the a posteriori covariance computed in (25) is not guaranteed to be bounded.
+As illustrated previously in Fig.
+2, without the “back out” step, the covariance may grow explosively, leading to severely inaccurate state estimates in subsequent iterations.
+It should be noted, however, that the impact of the “back out” step is application dependent.
+For example, as shown in Fig.
+9, removing this step in Application 1 noticeably degrades only the EKF performance, while EKF2, UKF, and CKF are almost unaffected.
+Nevertheless, it is still advisable to retain this step to guard against potential covariance explosion, since it incurs negligible computational overhead in the proposed framework.
+Appendix D presents the corresponding results for the other applications and reports, for all five applications, the percentage of update steps in which the “back out” mechanism is actually triggered.
+Figure 8: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (pendulum state estimation).
+Figure 9: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (3D target tracking).
+3.2 Covariance estimation accuracy
+As discussed previously, the proposed framework aims to improve covariance estimation by re-approximating the measurement functions after each state update.
+To evaluate whether the framework actually leads to more accurate covariance estimates, we introduce two metrics for assessing covariance estimation accuracy: the Average Normalized Estimation Error Squared (ANEES) and the Noncredibility Index (NCI).
+Specifically, the ANEES at time step $k$ is defined as [35]:
+$\mathrm{ANEES}_{k}:=\frac{1}{n_{x}N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}\bigl({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}\bigr)^{-1}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr),$ (33)
+where $n_{x}$ is the state dimension and $N_{\text{MC}}=10{,}000$ is the number of Monte Carlo simulations.
+The superscript $(i)$ denotes the $i^{\text{th}}$ Monte Carlo simulation.
+Here, $x_{k}^{(i)}$ and $\hat{x}_{k|k}^{(i)}$ are the true and estimated states at time step $k$, respectively, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}$} is the corresponding state estimation error covariance matrix.
+Note that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k|k-1,\mathrm{est}}}$ for the conventional framework, whereas ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k|k,\mathrm{est}}}$ for the proposed framework.
+When $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}$} equals the true error covariance and the estimation error is zero-mean Gaussian, we have $\operatorname{\mathbb{E}}[\mathrm{ANEES}_{k}]=1$.
+Therefore, the deviation of $\mathrm{ANEES}_{k}$ from 1 reflects the degree of overconfidence or conservatism in the KF.
+However, since the estimation error can be biased and non-Gaussian in practice, the value of ANEES alone may not be a sufficiently reliable indicator of overconfidence [35].
+Therefore, we also consider the metric NCI, which quantifies the mismatch between the estimated covariance and the empirical covariance of the estimation error.
+Specifically, the NCI at time step $k$ is defined as [36]:
+$\mathrm{NCI}_{k}:=\frac{10}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\log_{10}\frac{\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}\bigl({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}}\bigr)^{-1}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)}{\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}\bigl(P^{*}_{k}\bigr)^{-1}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)},$ (34)
+where $P^{*}_{k}$ is the sample error covariance at time step $k$, defined by
+$P^{*}_{k}:=\frac{1}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)\bigl(x_{k}^{(i)}-\hat{x}_{k|k}^{(i)}\bigr)^{T}.$ (35)
+By construction, $\mathrm{NCI}_{k}\approx 0$ when $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{k,\mathrm{est}}$} matches $P_{k}^{*}$, $\mathrm{NCI}_{k}>0$ indicates overconfidence, and $\mathrm{NCI}_{k}<0$ indicates conservatism.
+Given the definitions of ANEES and NCI for each step $k$, we define the global ANEES and NCI as the average across all time steps:
+$\begin{cases}\mathrm{ANEES}=\dfrac{1}{N_{\text{step}}}\displaystyle\sum_{k=1}^{N_{\text{step}}}\mathrm{ANEES}_{k},\\[8.00003pt] \mathrm{NCI}=\dfrac{1}{N_{\text{step}}}\displaystyle\sum_{k=1}^{N_{\text{step}}}\mathrm{NCI}_{k},\end{cases}$ (36)
+where $N_{\text{step}}$ is the total number of time steps in the simulation.
+Given these metrics, Fig.
+10 compares the global ANEES and NCI of various KF algorithms using the conventional and proposed frameworks in the target tracking application.
+Under the conventional framework, when the standard deviations of both measurements are below 0.1 m, all nonlinear KFs substantially underestimate the covariance, regardless of the specific algorithm.
+In contrast, under the proposed framework, only the EKF remains noticeably overconfident, while the ANEES and NCI of the other nonlinear KFs are very close to one and zero, respectively.
+This residual gap for the EKF is consistent with its neglect of second- and higher-order terms in the measurement functions; nevertheless, its covariance estimates are still significantly more accurate when combined with the proposed framework.
+Figure 10: The global Average Normalized Estimation Error Squared (ANEES) and Noncredibility Index (NCI) of various nonlinear Kalman filters under different measurement noise setups.
+The new framework significantly mitigates the problem of overconfidence.
+In Fig.
+11, we compare the estimated standard deviation of the first state $\hat{\sigma}_{1}$ using different KF variants and frameworks in the target tracking application.
+Specifically, $\hat{\sigma}_{1}$ is defined as the estimated posterior standard deviation of the $x$-axis position in the final iteration, averaged over all Monte Carlo simulations.
+Mathematically,
+$\hat{\sigma}_{1}:=\frac{1}{N_{\text{MC}}}\sum_{i=1}^{N_{\text{MC}}}\sqrt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P^{(i)}_{N_{\text{step}},\mathrm{est}}}(1,1)}.$ (37)
+Figure 11: Estimated standard deviation of the first state for various nonlinear Kalman filters under different measurement noise setups.
+Interestingly, Fig.
+11 suggests that the estimated standard deviation is almost the same for both frameworks, regardless of the KF algorithm.
+These results support the conclusion discussed in Section 2.4: the key to achieving more accurate state estimation is to reduce the gap between the estimated covariance and the actual covariance, rather than merely finding a better Kalman gain that minimizes the estimated covariance.
+3.3 Convergence characteristics
 Another important metric for evaluating the performance of different filters is the convergence characteristics.
-We visualized this metric by fixing the standard deviations of the two measurements to 0.01 m and plotted the RMSE of the state estimations after different numbers of iterations.
+We visualize this metric by fixing the standard deviations of the two measurements to 0.01 m and plotted the RMSE of the state estimations after different numbers of iterations.
 The results are shown in Fig.
-9.
+12.
 From the figure, we can see that all types of nonlinear KFs can converge faster after being combined with the new framework.
@@
 Figure 12: The root mean squared error of the state estimations after different numbers of iterations.
+3.4 Computational time
 Since the extra steps introduced in the new framework take time, it is also important to compare the computational time of the two frameworks.
 The results are summarized in Fig.
-10.
+13.
 Note that the average runtime shown in the figure is calculated from over 100,000 simulations run on the same computer.
-Specifically, for each combination of nonlinear KF and framework, we recorded the method’s runtime whenever it was used and calculated the average.
+Specifically, for each combination of nonlinear KF and framework, we record the method’s runtime whenever it was used and calculated the average.
 After the average runtime is calculated, the value is normalized with respect to the average runtime of the conventional EKF for the same application.
 In other words, the normalized average runtime of “EKF (old)” always equals one in Fig.
-10.
-Figure 10: Comparison of the computational time of different methods and frameworks.
+13.
+Figure 13: Comparison of the normalized computational time of different methods and frameworks.
 From the result, we can see that the new framework generally increases the computational time by 10–90%, varying for different nonlinear systems.
@@
 4 Conclusion
-In this paper, we proposed a new framework that can be applied to all types of nonlinear KFs.
+In this paper, we propose a new framework that can be applied to all types of nonlinear KFs.
 The new framework re-approximates the measurement functions after the state update, solving the problem of overconfident state covariance estimation, which is an inherent problem in the conventional framework for nonlinear KFs.
@@
 $x:\Omega\to\mathbb{R}^{n},\;y:\Omega\to\mathbb{R}^{n}$
-satisfy $\mathbb{E}[||x||^{2}]<\infty$ and $\mathbb{E}[||y||^{2}]<\infty$ (i.e., $x,y\in L^{2}(\Omega;\mathbb{R}^{n})$). Then, the operation on $L^{2}(\Omega;\mathbb{R}^{n})$ by
+[…]
 $f(u,v)=\mathbb{E}[u^{T}v]$
@@
 $\mathbb{E}[(ax+by)^{T}z]=a\mathbb{E}[x^{T}z]+b\mathbb{E}[y^{T}z].$
-Positive-definiteness is satisfied because for any $x\in L^{2}$, $\mathbb{E}[x^{T}x]\geq 0$, and the equality sign only holds when $x\equiv 0$.
+Positive-definiteness is satisfied because for any $x\in L^{2}$, $\mathbb{E}[x^{T}x]\geq 0$, and the equality sign only holds when $x=0$ almost surely.
 Lemma 4.
-Let ${\color[rgb]{1,0,0}S}$ be a random symmetric and positive definite matrix whose mean is the identity matrix $I$.
+Let $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S$} be a random symmetric and positive definite matrix whose mean is the identity matrix $I$.
 Let $\bm{v}$ be a random vector whose mean is $\bm{v}_{0}$.
 The following inequality holds:
-$\mathbb{E}(\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v})\geq\bm{v}_{0}^{T}\bm{v}_{0}$ (22)
-Note that the equality sign only holds when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
+$\mathbb{E}(\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v})\geq\bm{v}_{0}^{T}\bm{v}_{0}.$ (38)
+Note that the equality sign only holds when ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}$ is constant almost surely.
 {pf}
-Let $\bm{x}={\color[rgb]{1,0,0}S^{-\frac{1}{2}}}\bm{v},\bm{y}={\color[rgb]{1,0,0}S^{\frac{1}{2}}}\bm{v}_{0}$. From the definitions,
-$\begin{cases}\mathbb{E}[\bm{x}^{T}\bm{x}]=\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\\ \mathbb{E}[\bm{y}^{T}\bm{y}]=\mathbb{E}[\bm{v}_{0}^{T}{\color[rgb]{1,0,0}S}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\\ \mathbb{E}[\bm{x}^{T}\bm{y}]=\mathbb{E}[\bm{v}^{T}\bm{v}_{0}]=\bm{v}_{0}^{T}\bm{v}_{0}\end{cases}$ (23)
-According to Lemma 3, $\mathbb{E}[\bm{x}^{T}\bm{y}]$ is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$. Applying Cauchy–Schwarz inequality, we have:
-$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}S^{-1}}\bm{v}]\geq\frac{(\bm{v}_{0}^{T}\bm{v}_{0})^{2}}{\bm{v}_{0}^{T}\bm{v}_{0}}$ (24)
-which completes the proof. Note that the equality holds only when ${\color[rgb]{1,0,0}S^{-1}}\bm{v}$ is constant.
+Let $\bm{x}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-\frac{1}{2}}}\bm{v},\bm{y}:={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{\frac{1}{2}}}\bm{v}_{0}$.
+Since $\bm{v}$ is a random vector with a mean of $\bm{v}_{0}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S$} is a random symmetric matrix with a mean of the identity matrix,
+$\begin{cases}\mathbb{E}[\bm{x}^{T}\bm{x}]=\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}]\\ \mathbb{E}[\bm{y}^{T}\bm{y}]=\mathbb{E}[\bm{v}_{0}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}\bm{v}_{0}]=\bm{v}_{0}^{T}\mathbb{E}[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S}]\bm{v}_{0}=\bm{v}_{0}^{T}\bm{v}_{0}\\ \mathbb{E}[\bm{x}^{T}\bm{y}]=\mathbb{E}[\bm{v}^{T}\bm{v}_{0}]=\mathbb{E}[\bm{v}]^{T}\bm{v}_{0}=\bm{v}_{0}^{T}\bm{v}_{0}\end{cases}$ (39)
+According to Lemma 3, $\mathbb{E}[\bm{x}^{T}\bm{y}]$ is an inner product on $L^{2}(\Omega;\mathbb{R}^{n})$. Applying Cauchy–Schwarz inequality, we have
+$\mathbb{E}[\bm{v}^{T}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}]\geq\frac{(\bm{v}_{0}^{T}\bm{v}_{0})^{2}}{\bm{v}_{0}^{T}\bm{v}_{0}},$ (40)
+which completes the proof. Note that the equality holds only when ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{-1}}\bm{v}$ is constant almost surely.
 Appendix B The detailed simulation setup
@@
 As shown in Fig.
-11(a), a target moves in 3D space, and two sensors are used to track its location.
+14(a), a target moves in 3D space, and two sensors are used to track its location.
 The first sensor is at the origin, and the second sensor moves in a circle.
-The second sensor’s location $(x_{s1},x_{s2},x_{s3})$ at step $k=1,2,\ldots,30$ is represented by:
-$\begin{cases}x_{s1,k}=20+20\cos{\big{(}\frac{(k-1)\pi}{15}\big{)}}\\ x_{s2,k}=20+20\sin{\big{(}\frac{(k-1)\pi}{15}\big{)}}\\ x_{s3,k}=0\end{cases}$ (25)
+The second sensor’s location $(x_{s1},x_{s2},x_{s3})$ at step $k=1,2,\ldots,30$ is represented by
+$\begin{cases}x_{s1,k}=20+20\cos{\big(\frac{(k-1)\pi}{15}\big)},\\ x_{s2,k}=20+20\sin{\big(\frac{(k-1)\pi}{15}\big)},\\ x_{s3,k}=0.\end{cases}$ (41)
 The states are the target’s positions $(x_{1},x_{2},x_{3})$ and speeds $(v_{1},v_{2},v_{3})$ along the three axes.
-The measurements $y_{1},y_{2}$ are the distances between the target and the two sensors.
+The measurements $z_{1},z_{2}$ are the distances between the target and the two sensors.
 The inputs are the target’s accelerations along the three axes, which are assumed to be zero in this example.
-When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (26,27).
+When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (42) and (43).
 Note that the noises are omitted in the equations.
-$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ v_{1,k}\\ v_{2,k}\\ v_{3,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+v_{1,k-1}\\ x_{2,k-1}+v_{2,k-1}\\ x_{3,k-1}+v_{3,k-1}\\ v_{1,k-1}\\ v_{2,k-1}\\ v_{3,k-1}\end{bmatrix}$ (26)
-$\begin{bmatrix}y_{1,k}\\ y_{2,k}\end{bmatrix}=\begin{bmatrix}\sqrt{\sum_{i=1}^{3}x_{i,k}^{2}}\\ \sqrt{\sum_{i=1}^{3}(x_{i,k}-x_{si,k})^{2}}\end{bmatrix}$ (27)
+$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ v_{1,k}\\ v_{2,k}\\ v_{3,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+v_{1,k-1}\\ x_{2,k-1}+v_{2,k-1}\\ x_{3,k-1}+v_{3,k-1}\\ v_{1,k-1}\\ v_{2,k-1}\\ v_{3,k-1}\end{bmatrix}$, (42)
+$\begin{bmatrix}z_{1,k}\\ z_{2,k}\end{bmatrix}=\begin{bmatrix}\sqrt{\sum_{i=1}^{3}x_{i,k}^{2}}\\ \sqrt{\sum_{i=1}^{3}(x_{i,k}-x_{si,k})^{2}}\end{bmatrix}$. (43)
 The true initial states of the target are $x_{1,0}=10,x_{2,0}=-10,x_{3,0}=50,v_{1,0}=1,v_{2,0}=2,v_{3,0}=0$; the initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are 100, 100, 100, 0.01, 0.01, and 0.01.
@@
 As shown in Fig.
-11(b), a plane flies above an uneven landscape.
+14(b), a plane flies above an uneven landscape.
 The states are the planes’ positions along the two axes, denoted as $x_{1}$ and $x_{2}$.
-The measurement $y_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements.
+The measurement $z_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements.
 The plane knows the contour map in advance, so the terrain’s elevation can tell the aircraft its possible locations.
@@
 The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively.
-When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (28, 29).
+When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (44) and (45).
 Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.
-$\begin{bmatrix}x_{1,k}\\ x_{2,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+0.5\\ x_{2,k-1}\end{bmatrix}$ (28)
-$y_{k}=\sin{\sqrt{\big{(}\frac{x_{1,k}}{40}\big{)}^{2}+\big{(}\frac{x_{2,k}}{40}\big{)}^{2}}}$ (29)
+$\begin{bmatrix}x_{1,k}\\ x_{2,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+0.5\\ x_{2,k-1}\end{bmatrix}$, (44)
+$z_{k}=\sin{\sqrt{\big(\frac{x_{1,k}}{40}\big)^{2}+\big(\frac{x_{2,k}}{40}\big)^{2}}}.$ (45)
 The true initial states of the target are $x_{1,0}=10$ km, $x_{2,0}=10$ km; the initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are both 1 km2.
@@
 For simplicity, these states are denoted as $x_{1}$ to $x_{4}$.
-The measurement $y_{k}$ is electrical output power.
+The measurement $z_{k}$ is electrical output power.
 The system’s inputs include the mechanical input torque $T_{m}$, the steady-state internal voltage of the armature $E_{fd}$, and the terminal bus voltage $V_{t}$.
@@
 The input profiles are the same as in the paper, which are:
-$\begin{cases}u_{1,k}=0.8\\ u_{2,k}=2.11+0.0002k\\ u_{3,k}=1.002\end{cases}$ (30)
+$\begin{cases}u_{1,k}=0.8,\\ u_{2,k}=2.11+0.0002k,\\ u_{3,k}=1.002.\end{cases}$ (46)
 All the model parameters are selected to be the same as in [39].
-The system’s state transition and measurement functions can be written as (31, 32).
+The system’s state transition and measurement functions can be written as (47) and (48).
 Note that the noises are omitted from the equations.
-$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ x_{4,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}\\ x_{2,k-1}\\ x_{3,k-1}\\ x_{4,k-1}\end{bmatrix}$ (31)
-$+\begin{bmatrix}377x_{2,k-1}\Delta t\\ \frac{\Delta t}{13}\big{[}u_{1,k-1}-\frac{u_{3,k-1}x_{3,k-1}\sin{x_{1,k-1}}}{0.375}\big{]}\\ \frac{\Delta t}{0.131}[u_{2,k-1}-x_{3,k-1}]\\ \frac{-x_{4,k-1}\Delta t}{0.0131}\end{bmatrix}$
-$+\begin{bmatrix}0\\ \frac{\Delta t}{13}\big{[}0.9215u_{3,k-1}^{2}\sin(2x_{1,k-1})-0.05x_{2,k-1}\big{]}\\ \frac{-4.4933\Delta t}{0.131}(x_{3,k-1}-u_{3,k-1}\cos{x_{1,k-1}})\\ \frac{0.6911u_{3,k-1}\sin{x_{1,k-1}}\Delta t}{0.0131}\end{bmatrix}$
-$y_{k}=\frac{u_{3,k-1}x_{3,k}\sin{x_{1,k}}}{0.375}+0.9215u_{3,k-1}^{2}\sin(2x_{1,k})$ (32)
+$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ x_{4,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}\\ x_{2,k-1}\\ x_{3,k-1}\\ x_{4,k-1}\end{bmatrix}+\begin{bmatrix}377x_{2,k-1}\Delta t\\ \frac{\Delta t}{13}\big[u_{1,k-1}-\frac{u_{3,k-1}x_{3,k-1}\sin{x_{1,k-1}}}{0.375}\big]\\ \frac{\Delta t}{0.131}[u_{2,k-1}-x_{3,k-1}]\\ \frac{-x_{4,k-1}\Delta t}{0.0131}\end{bmatrix}$ (47)
+$+\begin{bmatrix}0\\ \frac{\Delta t}{13}\big[0.9215u_{3,k-1}^{2}\sin(2x_{1,k-1})-0.05x_{2,k-1}\big]\\ \frac{-4.4933\Delta t}{0.131}(x_{3,k-1}-u_{3,k-1}\cos{x_{1,k-1}})\\ \frac{0.6911u_{3,k-1}\sin{x_{1,k-1}}\Delta t}{0.0131}\end{bmatrix}$,
+$z_{k}=\frac{u_{3,k-1}x_{3,k}\sin{x_{1,k}}}{0.375}+0.9215u_{3,k-1}^{2}\sin(2x_{1,k}).$ (48)
 The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$.
-The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{-4}$, $10^{-10}$, $10^{-4}$, and $10^{-4}$.
+The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $^{-4}, 1010$$^{-10}$, $10^{-4}$, and $10^{-4}$.
 The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states.
-The process noise covariance matrix is also diagonal, and its diagonal values are $10^{-10}$, $10^{-16}$, $10^{-10}$, and $10^{-10}$.
+The process noise covariance matrix is also diagonal, and its diagonal values are $10^{-10}$, $^{-16}, 1010$$^{-10}$, and $10^{-10}$.
 B.4 Pendulum state estimation
 As shown in Fig.
-11(c), one end of the pendulum is fixed to the ceiling with a thin rope.
+14(c), one end of the pendulum is fixed to the ceiling with a thin rope.
 The system’s states are the pendulum’s angular position $\theta$ and angular speed $\omega$.
-The measurement $y_{k}$ is the horizontal component of tension in the rope.
+The measurement $z_{k}$ is the horizontal component of tension in the rope.
 The system input is the external torque exerted on the pendulum, which is zero in this example.
-The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is $g=9.8$m/s2.
+The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is g=9.8g=9.8m/s2.
 The time interval is set to $\Delta t=0.01$s.
-According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (33,34).
+According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (49) and (50).
 Note that the noises are omitted in the equations.
-$\begin{bmatrix}\omega_{k}\\ \theta_{k}\end{bmatrix}=\begin{bmatrix}\omega_{k-1}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \theta_{k-1}+\omega_{k-1}\Delta t\end{bmatrix}$ (33)
-$y_{k}=mg\cos{\theta_{k}}\sin{\theta_{k}}+ml\omega_{k}^{2}\sin{\theta_{k}}$ (34)
+$\begin{bmatrix}\omega_{k}\\ \theta_{k}\end{bmatrix}=\begin{bmatrix}\omega_{k-1}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \theta_{k-1}+\omega_{k-1}\Delta t\end{bmatrix}$, (49)
+$z_{k}=mg\cos{\theta_{k}}\sin{\theta_{k}}+ml\omega_{k}^{2}\sin{\theta_{k}}.$ (50)
 The true initial states of the system are $\omega_{0}=0,\theta_{0}=\pi/4$.
@@
 As shown in Fig.
-11(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$.
+14(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$.
 A battery’s two most important states are its state of charge (SOC) and state of health (SOH).
@@
 The states of this system are the SOC, SOH, and capacitor voltage $U_{c}$.
-The measurement $y_{k}$ is the battery’s terminal voltage.
+The measurement $z_{k}$ is the battery’s terminal voltage.
 The input is the charging current, which, in this example, is a three-level square wave shown below:
-$I_{k}=\begin{cases}-2,75\geq k>15\\ 2,165\geq k>105\\ 0,\text{else}\end{cases}$ (35)
+$I_{k}=\begin{cases}-2,&15<k\leq 75,\\ 2,&105<k\leq 165,\\ 0,&\text{else}.\end{cases}$ (51)
 The battery’s initial maximum capacity is $Q_{0}=1\,$Ah.
-The parameters in the equivalent circuit model are chosen to be $R_{1}=0.01,R_{2}=0.05,1/(R_{2}C_{1})=0.008$.
+The parameters in the equivalent circuit model are chosen to be $R_{1}=0.01\,\Omega,R_{2}=0.05\,\Omega,1/(R_{2}C_{1})=0.008\,\mathrm{s}$.
 The OCV is a function of SOC and SOH, and it can be written as:
-$OCV=\sum_{i=0}^{9}\big{(}\frac{SOH-0.8}{0.2}a_{100,i+1}+\frac{1-SOH}{0.2}a_{80,i+1}\big{)}$ (36)
-where $a_{100}=[1390.38,-6961.31,14760.31,-17230.92,$ 12055.71,5162.75,1330.60,196.37,15.60,2.96]12055.71,-5162.75,1330.60,-196.37,15.60,2.96]12055.71 , - 5162.75 , 1330.60 , - 196.37 , 15.60 , 2.96 ], $a_{80}=$ [813.94,4229.96,9345.49,11415.38,8396.15,3801.07,[813.94,-4229.96,9345.49,-11415.38,8396.15,-3801.07,[ 813.94 , - 4229.96 , 9345.49 , - 11415.38 , 8396.15 , - 3801.07 , 1043.09,165.29,14.28,2.96]1043.09,-165.29,14.28,2.96]1043.09 , - 165.29 , 14.28 , 2.96 ].
+$OCV=\sum_{i=0}^{9}\big(\frac{SOH-0.8}{0.2}a_{100,10-i}+\frac{1-SOH}{0.2}a_{80,10-i}\big)SOC^{i},$ (52)
+where $a_{100}=[1390.38,-6961.31,14760.31,-17230.92,$ 12055.71,5162.75,1330.60,196.37,15.60,2.96]12055.71,-5162.75,1330.60,-196.37,15.60,2.96], $a_{80}=$ [813.94,4229.96,9345.49,11415.38,8396.15,3801.07,[813.94,-4229.96,9345.49,-11415.38,8396.15,-3801.07, 1043.09,165.29,14.28,2.96]1043.09,-165.29,14.28,2.96].
 The time interval is set to $\Delta t=1$s, and the total simulation time is 180 seconds.
-The system’s state transition and measurement functions can be written as (37,38).
+The system’s state transition and measurement functions can be written as (53,54).
 Note that the noises are omitted in the equations.
 $\begin{bmatrix}SOC_{k}\\ U_{c,k}\\ SOH_{k}\end{bmatrix}$= $\begin{bmatrix}SOC_{k-1}\\ U_{c,k-1}\exp{\big(\frac{-\Delta t}{R_{2}C_{1}}\big)}\\ SOH_{k-1}\end{bmatrix}$ (53)
-$+\begin{bmatrix}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \big{[}1-\exp{\big{(}\frac{-\Delta t}{R_{2}C_{1}}\big{)}}\big{]}R_{2}I_{k-1}\\ 0\end{bmatrix}$
-$y_{k}=OCV(SOC_{k},SOH_{k})+U_{c,k}+R_{1}I_{k}$ (38)
+$+\begin{bmatrix}I_{k-1}\Delta t\\ \big[1-\exp{\big(\frac{-\Delta t}{R_{2}C_{1}}\big)}\big]R_{2}I_{k-1}\\ 0\end{bmatrix}$
+$z_{k}=OCV(SOC_{k},SOH_{k})+U_{c,k}+R_{1}I_{k}$ (54)
 The true initial states of the system are $SOC_{0}=60\%,U_{c,0}=0,SOH_{0}=90\%$.
@@
 Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
-The detailed algorithms of EKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 35.
+The detailed algorithms of EKF, IEKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 26.
 Note that all these algorithms degrade to the conventional framework when the “Recalibrate” and “Back out” steps are replaced by (13).
@@
 2:Initialization:
-3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
-4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
+[…]
+[…]
 5:for every time step $k$ do
@@
 11:  $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k-1}}$
-12:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})$
-13:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
-14:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
-15:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
-16:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
-17:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+12:  $\bm{\hat{z}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})$
+13:  $\bm{P}_{xz,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+14:  $\bm{P}_{z,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+15:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
+16:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
+17:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
 18:Recalibrate:
 19:  $\bm{H}_{k|k}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k|k}}$
-20:     $\bm{P}_{xy,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
-21:     $\bm{P}_{y,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
-22:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
-23:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+20:  $\bm{P}_{xz,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+21:  $\bm{P}_{z,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+22:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
+23:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
 24:Back out:
@@
 29:end for
+Algorithm 3 The iterated extended Kalman filter with the old framework
+1:Input: Process noise covariance matrix $\bm{Q}_{k}$, Measurement noise covariance matrix $\bm{R}_{k}$, state transition function $\bm{f}(\bm{x},\bm{u})$, measurement function $\bm{h}{(\bm{x})}$, system inputs $\bm{u}_{k}$, measurements $\bm{z}_{k}$
+2:Initialization:
+[…]
+[…]
+5:for every time step $k$ do
+6:Predict:
+7:  $\bm{F}_{k}=\frac{\partial\bm{f}}{\partial\bm{x}}|_{\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1}}$
+8:  $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
+9:  $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
+10:Update:
+11:  $\bm{x}_{0}=\bm{\hat{x}}_{k|k-1}$
+12:  for $i=1$ to 1000 do
+13:   $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{x}_{i-1}}$
+14:   $\bm{\hat{z}}_{k|k-1}=\bm{h}(\bm{x}_{i-1})+\bm{H}_{k|k-1}(\bm{x}_{i-1}-\bm{\hat{x}}_{k|k-1})$
+15:   $\bm{P}_{xz,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+16:   $\bm{P}_{z,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+17:   $\bm{S}_{k|k-1,i}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
+18:   $\bm{K}_{k,i}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1,i}^{-1}$
+19:   $\bm{x}_{i}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k,i}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
+[…]
+21:     $\bm{S}_{k|k-1,i}=\bm{S}_{k|k-1,i-1}$
+22:     $\bm{K}_{k,i}=\bm{K}_{k,i-1}$
+23:     $\bm{x}_{i}=\bm{x}_{i-1}$
+24:     break
+25:   end if
+26:   if $\max_{j}|1-\bm{x}_{i}(j)/\bm{x}_{i-1}(j)|<0.001$ then
+27:     break
+28:   end if
+29:  end for
+30:  $\bm{S}_{k|k-1}=\bm{S}_{k|k-1,i}$
+31:  $\bm{K}_{k}=\bm{K}_{i}$
+32:  $\bm{\hat{x}}_{k|k}=\bm{x}_{i}$
+33:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}-\bm{K}_{k}\bm{S}_{k|k-1}\bm{K}_{k}^{T}$
+34:end for
+35:Note that rows 16–21 are added to prevent the state estimations from diverging.
+Namely, when the sign of divergence is detected (row 16), the newest updates of the states, Kalman gain, and measurement residual covariance matrix are withdrawn (rows 17–19).
+If such a withdrawal is not made, we find in the simulation that the estimation accuracy can be much lower.
+This characteristic also suggests that the algorithm is not very robust and stable.
 Algorithm 4 The second-order extended Kalman filter with the new framework
@@
 2:Initialization:
-3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
-4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
+[…]
+[…]
 5:for every time step $k$ do
@@
 10:  $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}$
-11:      $+\frac{1}{2}\big{[}\text{tr}(\bm{F}_{i,k}^{*}\bm{P}_{k-1|k-1}\bm{F}_{j,k}^{*}\bm{P}_{k-1|k-1})\big{]}_{ij}+\bm{Q}_{k}$
+11:  $\frac{1}{2}\big[\text{tr}(\bm{F}_{i,k}^{*}\bm{P}_{k-1|k-1}\bm{F}_{j,k}^{*}\bm{P}_{k-1|k-1})\big]_{ij}+\bm{Q}_{k}$
 12:Update:
@@
 14:  $\bm{H}_{i,k|k-1}^{*}=\big[\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k|k-1}}\big]_{mn},i=1,\ldots,n_{m}$
-15:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{i}$
-16:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
-17:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
-18:      $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
-19:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
-20:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
-21:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+15:  $\bm{\hat{z}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})+\frac{1}{2}\big[\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1})\big]_{i}$
+16:  $\bm{P}_{xz,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+17:  $\bm{P}_{z,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
+18:  $\frac{1}{2}\big[\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k-1}^{*}\bm{P}_{k|k-1})\big]_{ij}$
+19:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
+20:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
+21:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
 22:Recalibrate:
@@
 24:  $\bm{H}_{i,k|k}^{*}=\big[\frac{\partial^{2}h_{i}}{\partial x_{m}\partial x_{n}}|_{\bm{\hat{x}}_{k|k}}\big]_{mn},i=1,\ldots,n_{m}$
-25:     $\bm{P}_{xy,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
-26:     $\bm{P}_{y,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
-27:       $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
-28:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
-29:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+25:  $\bm{P}_{xz,k|k}=\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+26:  $\bm{P}_{z,k|k}=\bm{H}_{k|k}\bm{P}_{k|k-1}\bm{H}_{k|k}^{T}$
+27:   $\frac{1}{2}\big[\text{tr}(\bm{H}_{i,k|k}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k}^{*}\bm{P}_{k|k-1})\big]_{ij}$
+28:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
+29:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
 30:Back out:
@@
 2:Initialization:
-3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
-4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
+[…]
+[…]
 5:for every time step $k$ do
 6:Predict:
-7:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$,
+7:  $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=\text{chol}(\bm{P}_{k-1|k-1})$,
 8:    where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
@@
 13:  $\bm{P}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}^{*}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{X}}_{i,k|k-1}^{*}$
-14:        $-\bm{\hat{x}}_{k|k-1}]^{T}+\bm{Q}_{k}$
+14:    $\bm{\hat{x}}_{k|k-1}]^{T}+\bm{Q}_{k}$
 15:Update:
-16:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k|k-1})$
+16:  $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=\text{chol}(\bm{P}_{k|k-1})$
 17:  $\bm{\mathcal{X}}_{i,k|k-1}=\bm{\hat{x}}_{k|k-1}+\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
 18:  $\bm{\mathcal{X}}_{i+n_{x},k|k-1}=\bm{\hat{x}}_{k|k-1}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
-19:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
-20:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k-1}$
-21:     $\bm{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
-22:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
-23:     $\bm{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
-24:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
-25:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
-26:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
-27:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+19:  $\bm{\mathcal{Z}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
+20:  $\bm{\hat{z}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Z}}_{i,k|k-1}$
+21:  $\bm{P}_{xz,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}$
+22:   $\bm{\hat{z}}_{k|k-1}]^{T}$
+23:  $\bm{P}_{z,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}$
+24:   $\bm{\hat{z}}_{k|k-1}]^{T}$
+25:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
+26:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
+27:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
 28:Recalibrate:
@@
 30:  $\bm{\mathcal{X}}_{i+n_{x},k|k}=\bm{\hat{x}}_{k|k}-\sqrt{n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
-31:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
-32:     $\bm{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k}$
-33:     $\bm{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
-34:     $\bm{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
-35:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
-36:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+31:  $\bm{\mathcal{Z}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
+32:  $\bm{\hat{z}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Z}}_{i,k|k}$
+33:  $\bm{P}_{xz,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
+34:  $\bm{P}_{z,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
+35:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
+36:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
 37:Back out:
@@
 2:Initialization:
-3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
-4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
-5:$\lambda=\alpha^{2}(n_{x}+\kappa)-n_{x}$
-6:$W_{i}^{(c)}=W_{i}^{(m)}=1/(2n_{x}+2\lambda),i=1,\ldots,n_{x}$
-7:$W_{0}^{(c)}=\lambda/(n_{x}+\lambda)+1-\alpha^{2}+\beta$
-8:$W_{0}^{(m)}=\lambda/(n_{x}+\lambda)$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 9:for every time step $k$ do
 10:Predict:
-11:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$, where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
-12:     $[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=i^{th}\text{ column of }\bm{L},i=1,...,n_{x}$
+11:  $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=\text{chol}(\bm{P}_{k-1|k-1})$, where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
+12:  $\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=i^{th}\text{ column of }\bm{L},i=1,...,n_{x}$
 13:  $\bm{\mathcal{X}}_{0,k-1}=\bm{\hat{x}}_{k-1|k-1}$
@@
 19:Update:
-20:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k})$
+20:  $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=\text{chol}(\bm{P}_{k-1|k})$
 21:  $\bm{\mathcal{X}}_{0,k|k-1}=\bm{\hat{x}}_{k|k-1}$
@@
 23:  $\bm{\mathcal{X}}_{i+n_{x},k|k-1}=\bm{\hat{x}}_{k|k-1}-\sqrt{\lambda+n_{x}}\bm{L}_{i},i=1,\ldots,n_{x}$
-24:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
-25:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k-1}$
-26:     $\bm{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
-27:     $\bm{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
-28:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
-29:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
-30:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
+24:  $\bm{\mathcal{Z}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
+25:  $\bm{\hat{z}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Z}}_{i,k|k-1}$
+26:  $\bm{P}_{xz,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}]^{T}$
+27:  $\bm{P}_{z,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}]^{T}$
+28:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
+29:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
+30:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
 31:Recalibrate:
-32:     $\bm{\mathcal{X}}_{i,k|k}=\bm{\mathcal{X}}_{i,k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1}),i=0,\ldots,2n_{x}$
-33:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
-34:     $\bm{\hat{y}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k}$
-35:     $\bm{P}_{xy,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
-36:     $\bm{P}_{y,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
-37:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
-38:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
+32:  $\bm{\mathcal{X}}_{i,k|k}=\bm{\mathcal{X}}_{i,k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1}),i=0,\ldots,2n_{x}$
+33:  $\bm{\mathcal{Z}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=0,\ldots,2n_{x}$
+34:  $\bm{\hat{z}}_{k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Z}}_{i,k|k}$
+35:  $\bm{P}_{xz,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
+36:  $\bm{P}_{z,k|k}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
+37:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
+38:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
 39:Back out:
@@
 44:end for
+Appendix D Supplementary results related to the “back out” step
+Fig.
+8, 9, 1517 compare the state estimation accuracy with and without the “back out” step.
+Additionally, Fig.
+1822 quantify how often the proposed “back out” step is activated in practice by reporting the percentage of state updates that trigger this safeguard.
+Figure 15: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (terrain-referenced navigation).
+Figure 16: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (synchronous generator state estimation).
+Figure 17: The root mean squared error of the state estimations with and without the “back out” step under different measurement noise setups (battery state estimation).
+Figure 18: Percentage of state updates that trigger “back out” (3D target tracking).
+Figure 19: Percentage of state updates that trigger “back out” (terrain-referenced navigation).
+Figure 20: Percentage of state updates that trigger “back out” (synchronous generator state estimation).
+Figure 21: Percentage of state updates that trigger “back out” (pendulum state estimation).
+Figure 22: Percentage of state updates that trigger “back out” (battery state estimation).